### v1

# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

Abstract
We study the problem of gradient descent learning of a single-index target function $f_{*}(\boldsymbol{x})=\textstyle\sigma_{*}\mathopen{}\mathclose{{}\left(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle}\right)$ under isotropic Gaussian data in $\mathbb{R}^{d}$, where the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ is an unknown degree $q$ polynomial with information exponent $p$ (defined as the lowest degree in the Hermite expansion). Prior works showed that gradient-based training of neural networks can learn this target with $n\gtrsim d^{\Theta(p)}$ samples, and such statistical complexity is predicted to be necessary by the correlational statistical query lower bound. Surprisingly, we prove that a two-layer neural network optimized by an SGD-based algorithm learns $f_{*}$ of arbitrary polynomial link function with a sample and runtime complexity of $n\asymp T\asymp C(q)\cdot d\mathrm{polylog}d$, where constant $C(q)$ only depends on the degree of $\sigma_{*}$, regardless of information exponent; this dimension dependence matches the information theoretic limit up to polylogarithmic factors. Core to our analysis is the reuse of minibatch in the gradient computation, which gives rise to higher-order information beyond correlational queries.
1 Introduction
Single-index models are a classical class of functions that capture low-dimensional structure in the learning problem. To efficiently estimate such functions, the learning algorithm should extract the relevant (one-dimensional) subspace from high-dimensional observations; hence this problem setting has been extensively studied in deep learning theory [BL20, BES+22, BBSS22, MHPG+23, WWF24], to examine the adaptivity to low-dimensional targets and benefit of representation learning in neural networks (NNs) optimized by gradient descent (GD). In this work we study the learning of a single-index target function with polynomial link function under isotropic Gaussian data:
$y_{i}=f_{*}(\boldsymbol{x}_{i})+\varsigma_{i},\quad f_{*}(\boldsymbol{x}_{i})=\sigma_{*}(\langle\boldsymbol{x}_{i},\boldsymbol{\theta}\rangle),\quad\boldsymbol{x}_{i}\stackrel{{\scriptstyle\mathrm{\tiny{i.i.d.}}}}{{\sim}}\mathcal{N}(0,\boldsymbol{I}_{d}),$ (1.1)
where $\varsigma_{i}$ is i.i.d. label noise, $\boldsymbol{\theta}\in\mathbb{R}^{d}$ is the direction of index features, and we assume the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ is a degree-$q$ polynomial with information exponent $p$ defined as the index of the first non-zero coefficient in the Hermite expansion (see Definition 1).
Equation (1.1) requires the estimation of the one-dimensional link function $\sigma_{*}$ and the relevant direction $\boldsymbol{\theta}$; it is known that learning is information theoretically possible with $n\gtrsim d$ training examples [Bac17, DPVLB24]. Indeed, when $\sigma_{*}$ is polynomial, such statistical complexity can be achieved up to logarithmic factors by a tailored algorithm that exploit the specific structure of the low-dimensional target [CM20]. On the other hand, for gradient-based training of two-layer NNs, existing works established a sample complexity of $n\gtrsim d^{\Theta(p)}$ [BAGJ21, BBSS22, DNGL23], which presents a gap between the information theoretic limit and what is computationally achievable by (S)GD. Such a gap is also predicted by the correlational statistical query (SQ) lower bound [DLS22, AAM23], which roughly states that for a CSQ algorithm to learn (isotropic) Gaussian single-index models using less than exponential compute, a sample size of $n\gtrsim d^{p/2}$ is necessary.
Although CSQ lower bounds are frequently cited to imply a fundamental barrier of learning via SGD (with the squared loss), strictly speaking, the CSQ model does not include empirical risk minimization with gradient descent, due to the non-adversarial noise and existence of non-correlational terms in the gradient computation. Very recently, [DTA+24] exploited higher-order terms in the gradient update arising from the reuse of the same training data, and showed that for certain link functions with high information exponent ($p>2$), two-layer NNs may still achieve weak recovery (i.e., nontrivial overlap with $\boldsymbol{\theta}$) after two GD steps with $O(d)$ batch size. While this presents evidence that GD-trained NNs can learn $f_{*}$ beyond the sample complexity suggested by the CSQ lower bound, the weak recovery statement in [DTA+24] may not translate to statistical guarantees; moreover, the class of functions where SGD can achieve vanishing generalization error is not fully characterized, as only a few specific examples of link functions are discussed.
(a) Online SGD (weak recovery).
(b) Same-batch GD (generalization error).
Figure 1: We train a two-layer ReLU NN (3.1) with $N=1024$ neurons using SGD with step size $\eta=1/d$ to learn a single-index target $f_{*}(\boldsymbol{x})=\mathsf{He}_{3}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)$; heatmaps are values averaged over 10 runs. $(a)$ online SGD with batch size $B=8$; $(b)$ GD on the same batch of size $n$ for $T=2^{14}$ steps. We only report weak recovery (i.e., overlap between parameters $\boldsymbol{w}$ and target $\boldsymbol{\theta}$, averaged across neurons) for online SGD since the generalization error does not drop.
Given the existence of (non-NN) algorithms that learn any single-index polynomials with $n=\tilde{O}(d)$ samples [CM20, DPVLB24], it is natural to ask if gradient-based training of NNs can achieve similar statistical efficiency for the same function class. Motivated by observations in [DTA+24] that SGD with reused batch may break the “curse of information exponent”, we aim to address the following question:
Can a two-layer NN optimized by SGD with reused batch learn arbitrary polynomial single-index models near the information-theoretic limit $n\asymp d$, regardless of the information exponent?
Empirically, the separation between one-pass (online) and multi-pass SGD is clearly observed in Figure 1, where we trained the same two-layer ReLU neural network to learn a single-index polynomial with information exponent $p=3$. We see that SGD with reused batch (Figure 1(b)) reaches low generalization error using roughly $n\asymp d$ samples, whereas online SGD fails to achieve even weak recovery with much larger sample size $n=\Omega(d^{2})$. Our main contribution is to establish this near-optimal sample complexity for two-layer NNs trained by a variant of SGD with reused training data.
1.1 Our Contributions
We answer the above question in the affirmative by showing that for (1.1) with arbitrary polynomial link function, SGD training on a natural class of shallow NNs can achieve small generalization error using polynomial compute and $n=\tilde{O}(d)$ training examples, if we employ a layer-wise optimization procedure (analogous to that in [BES+22, DLS22, AAM23]) and reuse of the same minibatch:
Theorem (informal).
A shallow NN with $N=\tilde{O}_{d}(1)$ neurons can learn arbitrary single-index polynomials up to small population loss: $\mathbb{E}_{\boldsymbol{x}}[(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]=o_{d,\mathbb{P}}(1)$, using $n=\tilde{O}_{d}(d)$ samples, and an SGD-based algorithm (with reused training data) minimizing the squared loss objective in $T=\tilde{O}_{d}(d)$ gradient steps.
We make the following remarks on our main result.
  • The theorem suggests that NN + SGD with reused batch can match the statistical efficiency of SQ algorithms tailored for low-dimensional polynomial regression [CM20]. Furthermore, the sample complexity is information theoretically optimal up to polylogarithmic factors, and surpasses the CSQ lower bound for $p>2$ (see Figure 2); this disproves a conjecture in [AAM23] stating that $n\asymp d^{p/2}$ is the optimal sample complexity for empirical risk minimization with SGD.
  • A key observation in our analysis is that with suitable activation function, SGD with reused batch can go beyond correlational queries and implement (a subclass of) SQ algorithms. This enables polynomial transformations to the labels that reduce the information exponent to (at most) 2, and hence optimization can escape the high-entropy “equator” at initialization in polylogarithmic time.
Upon completion of this work, we became aware of the preprint [ADK+24] showing weak recovery with similar sample complexity, also by exploiting the reuse of training data. Our work was conducted independently and simultaneously, and their contribution is not reflected in the present manuscript (beyond this paragraph).
Information theoretic limitSGD + batch reuse [This work] SQ algorithm [CM20]Smoothed SGD [DNGL23] CSQ lower bound [DLS22]One-pass SGD [BAGJ21]Kernel methods [GMMM21]$d$$\tilde{\Theta}(d)$$\tilde{\Theta}(d^{p/2})$$\tilde{\Theta}(d^{p-1})$$\Theta(d^{q})$
Figure 2: Sample complexity of learning single-index model where the link function $\sigma_{*}$ has degree $q$ and information exponent $p$. For the CSQ lower bound, we translate the tolerance to sample complexity using the i.i.d. concentration heuristic $\tau\approx n^{-1/2}$. We restrict ourselves to algorithms using polynomial compute; this excludes the sphere-covering procedure in [DPVLB24] or exponential-width neural network in [Bac17, TS24].
2 Problem Setting and Prior Works
Notations.
[…]
2.1 Complexity of Learning Single-index Models
We aim to learn a single-index model (1.1) where the link function $\sigma_{*}$ is a degree-$q$ polynomial with information exponent $p$ defined as follows.
Definition 1 (Information exponent [BAGJ21]).
Let $\{\mathsf{He}_{j}\}_{j=0}^{\infty}$ denote the normalized Hermite polynomials. The information exponent of $g\in L^{2}(\gamma)$, denoted by $\mathrm{IE}(g):=p\in\mathbb{N}_{+}$, is the index of the first non-zero Hermite coefficient of $g$, that is, given the Hermite expansion $g(z)=\sum_{i=0}^{\infty}\alpha_{i}\mathsf{He}_{i}(z)$, $p:=\min\{i>0:\alpha_{i}\neq 0\}$.
By definition we always have $p\leq q$. Note that $f_{*}$ contains $\Theta(d)$ parameters to be estimated, and hence information theoretically $n\gtrsim d$ samples are both sufficient and necessary for learning [MM18, BKM+19, DPVLB24]; however the sample complexity achieved by different (polynomial time) learning algorithms depends on structure of the link function.
  • Kernel Methods. Rotationally invariant kernels cannot adapt to the low-dimensional structure of single-index $f_{*}$ and hence suffer from the curse of dimensionality [YS19, GMMM21, DWY21, BES+22]. By a standard dimension argument [KMS20, HSSVG21, AAM22], we know that in the isotropic data setting, kernel methods (including neural networks in the lazy regime [JGH18, COB19]) require $n\gtrsim d^{q}$ samples to learn degree-$q$ polynomials in $\mathbb{R}^{d}$.
  • Gradient-based Training of NNs. While NNs can easily approximate a single-index model [Bac17], the statistical complexity of gradient-based learning established in prior works scales as $n\gtrsim d^{\Theta(p)}$: in the well-specified setting, [BAGJ21] proved a sample complexity of $n=\tilde{\Theta}(d^{p-1})$ for online SGD, which is later improved to $\tilde{\Theta}(d^{p/2})$ by a smoothed objective [DNGL23]; as for the misspecified setting, [BBSS22, DKL+23] showed that $n\gtrsim d^{p}$ samples suffice, and in some cases a $\tilde{\Theta}(d^{p-1})$ complexity is achievable [AAM23, OSSW24]. Consequently, at the information-theoretic limit ($n\asymp d$), existing results can only cover the learning of low information exponent targets [AAM22, BMZ23, BES+23]. It is worth noting that this exponential dependence on $p$ also appears in the CSQ lower bounds [DLS22, AAM22], which is often considered to be indicative of the performance of SGD learning with the squared loss (see Section 2.2).
  • Statistical Query Learners. If we do not restrict ourselves to correlational queries, then (1.1) can be efficiently solved near the information-theoretic limit. Specifically, [CM20] proposed an SQ algorithm that learns any single-index polynomials using $n=\tilde{O}(d)$ samples; the key ingredient is to construct nonlinear transformations to the labels that lowers the information exponent to $2$ (similar preprocessing also appeared in context of phase retrieval [MM18, BKM+19]). This is consistent with the recently established SQ lower bound [DPVLB24], which suggests a statistical complexity independent of the information exponent $p$ when $\sigma_{*}$ is polynomial.
2.2 Can Gradient Descent Go Beyond Correlational Queries?
Correlational statistical query.
A statistical query (SQ) learner [Kea98, Rey20] accesses the target $f_{*}$ through noisy queries $\tilde{\phi}$ with error tolerance $\tau$: $|\tilde{\phi}-\mathbb{E}_{\boldsymbol{x},y}[\phi(\boldsymbol{x},y)]|\leq\tau$. Lower bound on the performance of SQ algorithm is a classical measure of computational hardness. In the context of gradient-based optimization, an often-studied subclass of SQ is the correlational statistical query (CSQ) [BF02] where the query is restricted to (noisy version of) $\mathbb{E}_{\boldsymbol{x},y}[\phi(\boldsymbol{x})y]$. To see the connection between CSQ and SGD, consider the gradient of expected squared loss for one neuron $f_{\boldsymbol{w}}(\boldsymbol{x})$:
$\nabla_{\boldsymbol{w}}\mathbb{E}_{\boldsymbol{x},y}(f_{\boldsymbol{w}}(\boldsymbol{x})-y)^{2}\propto-\mathbb{E}_{\boldsymbol{x},y}[\underbrace{y\cdot\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})}_{\text{correlational query}}]\,+\,\mathbb{E}_{\boldsymbol{x}}[\underbrace{f_{\boldsymbol{w}}(\boldsymbol{x})\cdot\nabla_{\boldsymbol{w}}f_{\boldsymbol{w}}(\boldsymbol{x})}_{\text{\!\!\!\!can be evaluated without $y$\!\!\!\!}}].$ (2.1)
One can see that information of the target function is encoded in the correlation term in the gradient. To infer the statistical efficiency of GD in the empirical risk minimization setting, we replace the population gradient with the empirical average $\nabla_{\boldsymbol{w}}(\frac{1}{n}\sum_{i=1}^{n}(f_{\boldsymbol{w}}(\boldsymbol{x}_{i})-y_{i})^{2})$, and heuristically equate the CSQ tolerance $\tau$ with the scale of i.i.d. concentration error $n^{-1/2}$.
For the Gaussian single-index model class with information exponent $p$, [DLS22] proved a lower bound stating that a CSQ learner either has access to queries with tolerance $\tau\lesssim d^{-p/4}$, or exponentially many queries are needed to learn $f_{*}$ with small population loss. Using the heuristic $\tau\approx n^{-1/2}$, this suggests a sample complexity lower bound $n\gtrsim d^{p/2}$ for polynomial time CSQ algorithm. This lower bound can be achieved by a landscape smoothing procedure [DNGL23] (in the well-specified setting), and is conjectured to be optimal for empirical risk minimization with SGD [AAM23].
SGD with reused data.
As discussed in Section 2.1, the gap between SQ and CSQ algorithms primarily stems from the existence of label transformations that decrease the information exponent. While such transformation cannot be utilized by a CSQ learner, [DTA+24] argued that they may arise from two consecutive gradient updates using the same minibatch. For illustrative purposes, consider an example where one neuron $f_{\boldsymbol{w}}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}\rangle)$ is updated by two GD steps using the same training example $(\boldsymbol{x},y)$, starting from zero initialization $\boldsymbol{w}^{0}=\mathbf{0}$ (we focus on the correlational term in the loss for simplicity):
$\!\!\!\boldsymbol{w}^{2}=$ $~{}\boldsymbol{w}^{1}+\eta\cdot y\sigma^{\prime}(\langle\boldsymbol{x},\boldsymbol{w}^{1}\rangle)\boldsymbol{x}=\eta\sigma^{\prime}(0)\underbrace{y\cdot\boldsymbol{x}}_{\text{\!\!\!CSQ term\!\!\!}}+~{}\eta\underbrace{y\sigma^{\prime}(\eta\sigma^{\prime}(0)\mathopen{}\mathclose{{}\left\|{\boldsymbol{x}}}\right\|^{2}\cdot y)\boldsymbol{x}}_{\text{non-CSQ term}}.$ (2.2)
Under appropriate learning rate scaling $\eta\cdot\mathopen{}\mathclose{{}\left\|{\boldsymbol{x}}}\right\|^{2}=\Theta(1)$, one can see that in the second gradient step, the label $y$ is transformed by the nonlinearity $\sigma^{\prime}$, even though the loss function itself is not modified. Based on this observation, [DTA+24] showed that if the non-CSQ term in (2.2) reduces the information exponent to $1$, then weak recovery (i.e., nontrivial overlap between the first-layer parameters $\boldsymbol{w}$ and index features $\boldsymbol{\theta}$) can be achieved after two GD steps with $n=\Theta(d)$ samples.
2.3 Challenges in Establishing Statistical Guarantees
Importantly, the analysis in [DTA+24] does not lead to concrete learnability guarantees for the class of single-index polynomials for the following reasons: $(i)$ it is not clear if an appropriate nonlinear transformation that lowers the information exponent can always be extracted from SGD with reused batch, and $(ii)$ the weak recovery guarantee may not translate to a sample complexity for the trained NN to achieve small generalization error. We elaborate these technical challenges below.
SGD decreases information exponent.
To show weak recovery, [DTA+24, Definition 3.1] assumed that the student activation $\sigma$ can reduce the information exponent of the labels to $1$; while a few examples are given (e.g., odd-degree Hermite polynomials), the existence of such transformations in SGD is not guaranteed:
  • The label transformation employed in prior SQ algorithms [CM20] is based on thresholding, which reduces the information exponent to $2$ for any polynomial $\sigma_{*}$; however, isolating such function from SGD updates on the squared loss is challenging. Instead, we show in Proposition 5 that monomial transformation suffices, and it can be extracted from SGD via Taylor expansion.
  • If the link function $\sigma_{*}$ is even, then its information exponent after arbitrary nonlinear transformation is at least $2$; such functions are predicted not be not learnable by SGD in the $n\asymp d$ regime [DTA+24]. To handle this setting, we analyze the SGD update up to $\mathrm{polylog}(d)$ time, at which a nontrivial overlap can be established by a Grönwall-type argument similar to [BAGJ21]. This recovers results on phase retrieval when $\sigma_{*}(z)=z^{2}$ which requires $n=\Omega(d\log d)$.
From weak recovery to sample complexity.
Note that weak recovery (i.e., $|\langle\boldsymbol{w},\boldsymbol{\theta}\rangle|>\varepsilon$ for some small constant $\varepsilon>0$) is generally insufficient to establish low generalization error of the trained NN. Therefore, we need to show that starting from a nontrivial overlap, subsequent gradient steps can achieve strong recovery of the index features (i.e., $|\langle\boldsymbol{w},\boldsymbol{\theta}\rangle|>1-\varepsilon$), despite the link misspecification. After the first-layer parameters align with the target function, we can train the second-layer parameters with SGD to learn the link function $\sigma_{*}$ with the aid of random bias units.
3 SGD Achieves Almost-linear Sample Complexity
We consider the learning of single-index polynomials with degree $q$ and information exponent $p$:
Assumption 1.
The target function is given as $f_{*}(\boldsymbol{x})=\sigma_{*}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)$, where the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ admits the Hermite decomposition $\sigma_{*}(z)=\sum_{i=p}^{q}\alpha_{i}\mathsf{He}_{i}(z)$.
3.1 Training Algorithm
We train the following two-layer neural network with $N$ neurons using SGD to minimize the squared loss:
$f_{\boldsymbol{\Theta}}(\boldsymbol{x})=\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}(\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle+b_{j}),$ (3.1)
where $\boldsymbol{\Theta}=(\boldsymbol{w}_{j},a_{j},b_{j})_{j=1}^{N}$ are trainable parameters, and $\sigma_{j}:\mathbb{R}\to\mathbb{R}$ is the activation function defined as the sum of Hermite polynomials up to degree $C_{\sigma}$: $\sigma_{j}(z):=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, where $C_{\sigma}$ only depends on the degree of link function $\sigma_{*}$. Note that we allow each neuron to have a different nonlinearity as indicated by the subscript in $\sigma_{j}$; this subscript is omitted when we focus on the dynamics of one single neuron. Our SGD training procedure is described in Algorithm 1, and below we outline the key ingredients of the algorithm.
  • Algorithm 1 employs a layer-wise training strategy common in the recent feature learning theory literature [DLS22, BES+22, BBSS22, AAM23, MHWSE23], where in the first stage, we optimize the first-layer parameters $\{\boldsymbol{w}_{j}\}_{j=1}^{N}$ with normalized SGD to learn the low-dimensional latent representation (index features $\boldsymbol{\theta}$), and in the second phase, we train the second-layer $\{a_{j}\}_{j=1}^{N}$ to fit the unknown link function $\sigma_{*}$.
  • The most important part in Phase I of Algorithm 1 is the reuse of the same minibatch in the gradient computation. Specifically, we sample a fresh batch of training examples in every two GD steps; this enables us to extract non-CSQ terms from two consecutive gradient updates outlined in (2.2).
  • We introduce an interpolation step between the current and previous iterates with hyperparameter $\xi$ to stabilize the training dynamics; this resembles a negative momentum term often seen in optimization algorithms [AZ18, ZLBH19]; the role of this interpolation step is discussed in Section 4.2. We also use a projected gradient update ~$\tilde{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})=(\boldsymbol{I}_{d}-\boldsymbol{w}^{2t}{\boldsymbol{w}^{2t}}^{\top})\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})$ for steps $2t$ and $12t+$ 1, where $\nabla_{\boldsymbol{w}}$ is the Euclidean gradient. Similar use of projection also appeared in [DNGL23, AAM23].
1
Input : Learning rates $\eta^{t}$, momentum parameters $\{\xi_{j}^{t}\}_{j=1}^{N}$, number of steps $T_{1},T_{2}$,
$\ell_{2}$ regularization $\lambda$.
2
3Initialize $\boldsymbol{w}^{0}_{j}\sim\mathbb{S}^{d-1}(1)$, $a_{j}\sim\mathrm{Unif}\{\pm r_{a}\}$.
4Phase I: normalized SGD on first-layer parameters
5       for $t=0$ to $T_{1}$ do
6             if $t>0$ is even then
7                   Draw i.i.d. sample $(\boldsymbol{x},y)$.
8                   Interpolate ${\boldsymbol{w}}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\xi_{j}^{t}(\boldsymbol{w}^{t}_{j}-\boldsymbol{w}^{t-2}_{j})$.
9                   Normalize $\boldsymbol{w}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}/\|\boldsymbol{w}^{t}_{j}\|$.
10            $\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2},\quad(j=1,\dots,N)$.
11
12
13Initialize $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$.
14Phase II: SGD on second-layer parameters
15       $\hat{\boldsymbol{a}}\leftarrow\mathrm{argmin}_{\boldsymbol{a}\in\mathbb{R}^{N}}\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\mathopen{}\mathclose{{}\left(f_{\boldsymbol{\Theta}}(\boldsymbol{x}_{i})-y_{i}}\right)^{2}+\lambda\|\boldsymbol{a}\|^{2}.$
16
Output : Prediction function $\boldsymbol{x}\mapsto f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})$ with $\hat{\boldsymbol{\Theta}}=(\hat{a}_{j},\boldsymbol{w}^{T_{1}}_{j},b_{j})_{j=1}^{N}$.
Algorithm 1 Gradient-based training of two-layer neural network
3.2 Convergence and Statistical Complexity
Weak Recovery Guarantee.
We first consider the “search phase” of SGD, and show that after running Phase I of Algorithm 1 for $T=\mathrm{polylog}(d)$ steps, a subset of parameters $\boldsymbol{w}$ achieve nontrivial overlap with the target direction $\boldsymbol{\theta}$. We denote $H(g;j)$ as the $j$-th Hermite coefficient of some $g\in L^{2}(\gamma)$. Our main theorems handle polynomial activations satisfying the following condition.
Assumption 2.
We require the activation function to be a polynomial $\sigma(z)=\sum_{i=0}^{C_{\sigma}}\beta_{i}\mathsf{He}_{i}(z)$ and its degree $C_{\sigma}$ to be sufficiently large so that $C_{\sigma}\geq C_{q}$ holds ($C_{q}$ is defined in Proposition 5). For all $2\leq\ell\leq C_{\sigma}$ and $k=0,1$, we assume that $H\big{(}\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k\big{)}\neq 0$.
More precisely, a given $\sigma_{*}$, the above condition only needs to be met for one pair of $(k,\ell)$, as specified in Appendix B.1. We show that this condition is satisfied for a wide range of polynomials with degree $C_{\sigma}$.
Lemma 2.
Given any $\ell\geq 2$ and $k\geq 0$. For $C_{\sigma}\geq\frac{2\ell+k-1}{\ell}$, if we choose $\{\beta_{i}\}_{i=0}^{C_{\sigma}}$ where each $\beta_{i}$ is randomly drawn from some interval $[a_{i},b_{i}]$ with $a_{i}<b_{i}$, then $H(\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k)\neq 0$ with probability one.
The following theorem states that $n=\tilde{\Theta}(d)$ samples are sufficient to achieve weak recovery.
Theorem 1.
Under Assumptions 1 and 2, for suitable choices of hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $\xi_{j}^{t}=O(1)$, there exists constant $C(q)$ such that after Phase I of Algorithm 1 is run for $2T_{1,1}=C(q)\cdot d\mathrm{polylog}(d)$ steps, with high probability, there exists a subset of neurons $\boldsymbol{w}_{j}^{2T_{1}}\in\mathcal{W}$ with $|\mathcal{W}|=\tilde{\Theta}(N)$ such that $\big{|}\langle\boldsymbol{w}_{j}^{2T_{1}},\boldsymbol{\theta}\rangle\big{|}>c$ for some $c\gtrsim 1/\mathrm{polylog}(d)$.
Recall that at random initialization we have $\langle\boldsymbol{w},\boldsymbol{\theta}\rangle\approx d^{-1/2}$ with high probability. The theorem hence implies that SGD “escapes from mediocrity” after seeing $n=\tilde{O}(d)$ samples, analogous to the information exponent $p=2$ setting studied in [BAGJ21]. We remark that due to the small second-layer initialization, the squared loss is dominated by the correlation loss, which allows us to track the evolution of each neuron independently; similar use of vanishing initialization also appeared in [BES+22, AAM23].
Strong recovery and sample complexity.
After weak recovery is achieved, we continue Phase I to amplify the alignment. Due to the nontrivial overlap between $\boldsymbol{w}$ and $\boldsymbol{\theta}$, the objective is no longer dominated by the lowest degree in the Hermite expansion. Therefore, to establish strong recovery ($\langle\boldsymbol{w},\boldsymbol{\theta}\rangle>1-\varepsilon$), we place an additional assumption on the activation function.
Assumption 3.
Recall the Hermite expansions $\sigma_{*}(z)=\sum_{i=p}^{q}\alpha_{i}\mathsf{He}_{i}(z)$, $\sigma_{j}(z)=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, we assume the coefficients satisfy $\alpha_{i}\beta_{j,i}\geq 0$ for $p\leq i\leq q$.
This assumption is easily verified in the well-specified setting $\sigma_{*}=\sigma$ [BAGJ21] since $\alpha_{i}=\beta_{i}$, and under link misspecification, it has been directly assumed in prior work [MHWSE23]. We follow [OSSW24] and show that by randomizing the Hermite coefficients of the activation function, a subset of neurons satisfy the above assumption for any degree-$q$ polynomial link function $\sigma_{*}$.
Lemma 3.
If we set $\sigma_{j}(z)=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, where for each neuron we sample $\beta_{j,i}\stackrel{{\scriptstyle\mathrm{\tiny{i.i.d.}}}}{{\sim}}\mathrm{Unif}(\{\pm r_{i}\})$ with appropriate constant $r_{i}$, then Assumption 2 and 3 are satisfied in $\operatorname{exp}\mathopen{}\mathclose{{}\left({-\Theta(q)}}\right)$-fraction of neurons.
The proof is found in Appendix B.1. Note that in our construction of activation functions for both assumptions, we do not exploit knowledge of the link function $\sigma_{*}$ other than its degree $q$ which decides the constant $C_{\sigma}$. The next theorem shows that by running Phase I for $\tilde{\Theta}(d)$ more steps, a subset of neurons can achieve almost perfect overlap with the index features.
Theorem 2.
For student neurons with activation satisfying Assumptions 2 and 3 and parameter $\boldsymbol{w}_{j}$ starting from a nontrivial overlap $c$ specified in Theorem 1, for suitable hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $\xi_{j}^{t}=1-\tilde{O}(\varepsilon)$, if we continue to run Phase I of Algorithm 1 for $2T_{1,2}=\tilde{\Theta}_{d}(d\varepsilon^{-2})$ steps, then we achieve $\big{\langle}\boldsymbol{w}_{j}^{2(T_{1,1}+T_{1,2})},\boldsymbol{\theta}\big{\rangle}>1-\varepsilon$ with high probability.
Combining Theorem 1,2 and Lemma 3, we know that after $T_{1}=2(T_{1,1}+T_{1,2})$ steps, there exists some constant $C(q)$ depending only on the degree $q$ such that $1/(C(q)\cdot\mathrm{polylog}(d))$ fraction of the neurons become $\varepsilon$-close to the target direction $\boldsymbol{\theta}$. The following proposition shows that after strong recovery, training the second-layer parameters in Phase II is sufficient for the NN model (3.1) to achieve small generalization error.
Proposition 4.
After Phase I terminates, for suitable $\lambda>0$, the output of Phase II satisfies
$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon.$
with probability 1 as $d\to\infty$, if we set $T_{2}=C(q)N^{4}\mathrm{polylog}(d)\varepsilon^{-2}$, $N=C(q)\mathrm{polylog}(d)\varepsilon^{-1}$ for some constant $C(q)$ depending on the target degree $q$.
Putting things together.
Combining the above theorems, we conclude that in order for two-layer NN (3.1) trained by Algorithm 1 to achieve $\varepsilon$ population squared loss, it is sufficient to set
$n=T_{1}+T_{2}\asymp C(q)\cdot(d\varepsilon^{-2}\vee\varepsilon^{-6})\mathrm{polylog}(d),\quad N\asymp C(q)\cdot\varepsilon^{-1}\mathrm{polylog}(d),$ (3.2)
where constant $C(q)$ only depends on the target degree $q$ (although exponentially). Hence we may set $\varepsilon^{-1}\asymp\mathrm{polylog}d$ to conclude an almost-linear sample and computational complexity for learning arbitrary single-index polynomials up to $o_{d}(1)$ population error. This establishes the informal theorem in Section 1.
4 Proof Sketch
In this section we outline the high-level ideas and key steps in our derivation.
4.1 Monomial Transformation Reduces Information Exponent
To prove the main theorem, we first establish the existence of nonlinear label transformation that $(i)$ reduces the information exponent, and $(ii)$ can be easily extracted from SGD updates. If we ignore desideratum $(ii)$, then for polynomial link functions, transformations that decrease the information exponent to at most $2$ have been constructed in [CM20, Section 2.1]. However, prior results are based on the thresholding function with specific offset, and it is not clear if such function naturally arises from SGD with batch reuse. The following proposition shows that the effect of thresholding can also be achieved by a simple monomial transformation.
Proposition 5.
Let $f:\mathbb{R}\to\mathbb{R}$ be any polynomial with degree up to $p$ and $\mathopen{}\mathclose{{}\left\|{f}}\right\|_{L^{2}(\gamma)}^{2}=1$, then
  • (i)
    There exists some $i\leq C_{q}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(f^{i})\leq 2$, where constant $C_{q}$ only depends on $q$.
  • (ii)
    Let $f^{\mathrm{odd}}:\mathbb{R}\to\mathbb{R}$ be the odd part of $f$ with $\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{\mathrm{odd}}(t)^{2}]\geq\rho>0$. Then there exists some $i\leq C_{q,\rho}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(f^{i})=1$, where constant $C_{q,\rho}$ only depends on $q$ and $\rho$.
We make the following remarks.
  • The proposition implies that for any polynomial link function that is not even, there exists some power $i\in\mathbb{N}_{+}$ only depending on the degree of $\sigma_{*}$ such that raising the function to the $i$-the power reduces the information exponent to $1$. For even link functions, the information exponent after arbitrary transformation is at least $2$, and this lowest value can also be achieved by monomial transformation. Furthermore, we provide a uniform upper-bound on the required degree of transformation $i$ via a compactness argument.
  • The advantage of working with monomial transformations is that they can be obtained from two GD steps on the same training example, by Taylor expanding the activation $\sigma^{\prime}$. In Section 4.2, we build upon this observation to show that Phase I of Algorithm 1 achieves weak recovery using $n\gtrsim d\,\mathrm{polylog}(d)$ samples.
Intuition behind the analysis.
Our proof is inspired by [CM20] which introduced a (non-polynomial) label transformation that reduces the information exponent of any degree-$q$ polynomial to at most $2$. To prove the existence of monomial transformation for the same purpose, we first show that for a fixed link function $\sigma_{*}$, there exists some $i$ such that the $i$-th power of the link function has information exponent $2$, which mirrors the transformation used in [CM20]. Then, we make use of the compactness of the space of link functions to define a test function and obtain a uniform bound on $i$. As for the polynomial transformation for non-even functions, we exploit the asymmetry of $\sigma_{*}$ to further reduce the information exponent to 1.
4.2 SGD with Batch Reuse Implements Polynomial Transformation
Now we present a more formal discussion of (2.2) to illustrate how polynomial transformation can be utilized in batch reuse SGD. We let $\eta^{t}\equiv\eta$. When one neuron $f_{\boldsymbol{w}}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}\rangle)$ is updated by two GD steps using the same sample $(\boldsymbol{x},y)$, starting from $\boldsymbol{w}^{0}:=\boldsymbol{\omega}$, the alignment with $\boldsymbol{\theta}$ becomes
$\langle\boldsymbol{\theta},\boldsymbol{w}^{2}\rangle=\mathopen{}\mathclose{{}\left\langle\boldsymbol{\theta},\big{[}\boldsymbol{w}^{1}+\eta\cdot y\sigma^{\prime}(\langle\boldsymbol{x},\boldsymbol{w}^{1}\rangle)\boldsymbol{x}\big{]}}\right\rangle=\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle+$ (4.1)
$\eta\bigg{[}y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle+\sum_{i=0}^{C_{\sigma}-1}\underbrace{(\eta\|\boldsymbol{x}\|^{2})^{i}y^{i+1}(i!)^{-1}(\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{i}\sigma^{(i+1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle}_{=:\psi_{i}}\bigg{]}.$ (4.2)
We take $\eta\leq c_{\eta}d^{-1}$ with a small constant $c_{\eta}$ so that $\eta\|\boldsymbol{x}\|^{2}\ll 1$. Crucially, the strength of each term in (4.2) can vary depending on properties of the link function $\sigma_{*}$, which is unknown. Hence a careful analysis is required to ensure that the suitable monomial transformation is always singled out from the gradient update. We therefore divide our analysis into four cases (for simplicity we present the noiseless setting below).
  • (I)
    If $\mathrm{IE}(\sigma_{*})=1$. All terms in the summation in (4.2) with $i\geq 2$ decay as fast as $\eta\|\boldsymbol{x}\|^{2}\ll 1$. On the other hand, the expectation of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ is roughly $\alpha_{1}\beta_{1}\gtrsim 1$. Therefore we may isolate the informative term $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$. This case is discussed in Section B.3.3.
  • (II)
    Else if $\mathrm{IE}((\sigma_{*})^{I})=1$ for some $2\leq I\leq C_{\sigma}$. Let $I$ be the lowest degree of monomial transformation such that $\mathrm{IE}((\sigma_{*})^{I})=1$. Since $\sigma_{*},\cdots,\sigma_{*}^{I-1}$ have information exponent larger than $1$, expectations of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ and $\psi_{i}\ (i=2,\cdots,I-2)$ scales as $\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle\asymp d^{-1/2}$. For $i=I-1$, because $\sigma_{*}^{I}$ has information exponent $1$,
    $\mathbb{E}[\psi_{I-1}]=\,$ $\mathbb{E}\big{[}(\eta\|\boldsymbol{x}\|^{2})^{I-1}y^{I}((I-1)!)^{-1}(\sigma^{(1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{I-1}\sigma^{(I)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle\big{]}$ (4.3)
[…]
    For $i\geq I$, $\psi_{i}$ decays as $c_{\eta}^{I}$, which is smaller than the scale of $\psi_{I-1}\asymp c_{\eta}^{I-1}$. Hence the term $\psi_{I-1}\gtrsim c_{\eta}^{I-1}$ is singled out. This case is discussed in Section B.3.2.
  • (III)
    Else if $\mathrm{IE}(\sigma_{*})=2$. We have $\mathbb{E}[y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle]\approx 2\alpha_{2}\beta_{2}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$. Also, for $i\geq 2$, since $\sigma_{*}^{2},\cdots,\sigma_{*}^{C_{\sigma}}$ have information exponent at least $2$, expectation of $\psi_{i}$ is roughly of order $(\eta\|\boldsymbol{x}\|^{2})^{i}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$. Therefore, the term $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ is singled out, and the expectation scales as $\alpha_{2}\beta_{2}d^{-1/2}$ at initialization. This case is discussed in Section B.3.3.
  • (IV)
    Else $\mathrm{IE}((\sigma_{*})^{I})=2$ for some $2\leq I\leq C_{\sigma}$. In this case, since $\sigma_{*},\cdots,\sigma_{*}^{I-1}$ have information exponent larger than $2$, expectations of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ and $\psi_{i}\ (i=2,\cdots,I-2)$ are at most $(\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle)^{2}\asymp d^{-1}$. And at $i=I-1$, because $\sigma_{*}^{I}$ has information exponent $2$,
    $\mathbb{E}[\psi_{I-1}]=\,$ $\mathbb{E}\big{[}(\eta\|\boldsymbol{x}\|^{2})^{I-1}y^{I}((I-1)!)^{-1}(\sigma^{(1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{I-1}\sigma^{(I)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle\big{]}$ (4.5)
[…]
    As for $i\geq I$, because $\sigma_{*}^{i}$ has information exponent larger than $1$ for $I+1\leq i\leq C_{\sigma}$, $\psi_{i}$ decays as $c_{\eta}^{I}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$, which is smaller than $\psi_{I-1}\asymp c_{\eta}^{I-1}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$. Thus, the term $\psi_{I-1}$ is dominating, whose scale is roughly $c_{\eta}^{I-1}d^{-1/2}$ at initialization. This case is discussed in Section B.3.1.
Why interpolation is required.
[…]
Combining the four cases yields the following lemma on the evolution of alignment.
Lemma 6.
Under the assumptions per Theorem 1, one of the following holds:
  • (i)
    $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\gamma+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\nu^{2t}_{j}.$
  • (ii)
    $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\gamma\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\nu^{2t}_{j}.$
Here $\gamma\gtrsim c_{\eta}^{I-1}$ is a constant that depends on $\sigma_{*}$ and $\nu^{2t}_{j}$ is a mean-zero noise.
For (i), taking expectation immediately yields that weak recovery is achieved within $(\eta(1-\xi)\gamma)^{-1}=\tilde{O}(d)$ steps. For (ii), $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle=:\kappa^{t}$ can be approximated by a differential equation $\frac{\mathrm{d}\kappa^{t}}{\mathrm{d}t}=\eta(1-\xi)\gamma\kappa^{t}$. Solving this yields $\kappa^{t}=\kappa^{0}\exp(\eta(1-\xi)\gamma t)\approx d^{-\frac{1}{2}}\exp(\eta(1-\xi)\gamma t)$, and weak recovery is obtained within $t\lesssim(\eta(1-\xi)\gamma)^{-1}\cdot\log d=\tilde{O}(d)$ steps, similar to the analysis in [BAGJ21].
4.3 Analysis of Phase II and Statistical Guarantees
Once strong recovery is achieved for the first-layer parameters, we turn to Phase II and optimize the second-layer with $\ell_{2}$ regularization. Since the objective is strongly convex, gradient-based optimization can efficiently minimize the empirical loss. The learnability guarantee follows from standard analysis analogous to that in [AAM22, DLS22, BES+22], where we construct a “certificate” second-layer $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ that achieves small loss and small norm:
$\mathbb{E}_{\boldsymbol{x}}\mathopen{}\mathclose{{}\left(f_{*}(\boldsymbol{x})-\frac{1}{N}\sum_{j=1}^{N}a^{*}_{j}\sigma_{j}\big{(}\langle{\boldsymbol{w}_{j}^{T_{1}}},\boldsymbol{x}\rangle+b_{j}\big{)}}\right)^{2}\leq\varepsilon^{*},\quad\|\boldsymbol{a}^{*}\|\lesssim r^{*},$
from which the population loss of the regularized empirical risk minimizer can be bounded via standard Rademacher complexity argument. To construct such a certificate, we make use of the random bias units $\{b_{j}\}_{j=1}^{N}$ to approximate the link function $\sigma_{*}$ as done in [DLS22, BBSS22, OSSW24].
5 Conclusion and Future Directions
In this work we showed that a two-layer neural network (3.1) trained by SGD with reused batch can learn arbitrary single-index polynomials up to $\varepsilon$ population error using $n=\tilde{O}(d\varepsilon^{-2})$ samples and compute. Our analysis is based on the observation that by reusing the same minibatch twice in the gradient computation, a non-correlational term arises in the SGD update that transforms the labels (despite the loss function is not modified). Specifically, following the definition in [DPVLB24], we know that polynomial $\sigma_{*}$ has generative exponent at most $2$, which implies the existence of nonlinear transformation $\mathcal{T}:\mathbb{R}\to\mathbb{R}$ such that the information exponent $p_{*}$ becomes at most $2$, i.e.,
$\mathbb{E}[\mathcal{T}(y)\mathsf{He}_{i}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)]\neq 0,\quad\text{for $i=1$ or $2$.}$
We show that restricting $\mathcal{T}$ to be polynomial is sufficient, and such transformation can be extracted by Taylor-expanding the SGD update. Then we show via careful analysis of the trajectory that strong recovery and low population error can be achieved under suitable activation function.
Future directions.
First, our analysis only handles link functions with generative exponent $p_{*}\leq 2$; while this covers arbitrary polynomial $\sigma_{*}$ analogous to [CM20], it is interesting to examine whether SGD with reused batch can learn targets with $p_{*}\geq 3$ with a sample complexity matching the SQ lower bound. It is also possible that ERM algorithms on i.i.d. data can achieve a statistical complexity beyond the SQ lower bound due to non-adversarial noise [DH21, DH24]; such mechanism is not exploited in our current analysis. Additional interesting directions include extension to multi-index [BAGJ22, BBPV23, CWPPS23, Gla23], hierarchical polynomials [NDL23], and additive models [OSSW24]. Lastly, the SGD algorithm that we employ requires a layer-wise training procedure and a specific batch reuse schedule; one may therefore ask if standard multi-pass SGD training of all parameters simultaneously (as reported in Figure 1) also achieves the same statistical efficiency.
Acknowledgements
The authors thank Gerard Ben Arous, Joan Bruna, Alex Damian, Marco Mondelli, and Eshaan Nichani for the discussions and feedback on the manuscript. JDL acknowledges support of the ARO under MURI Award W911NF-11-1-0304, NSF CCF 2002272, NSF IIS 2107304, NSF CIF 2212262, ONR Young Investigator Award, and NSF CAREER Award 2144994. KO was partially supported by JST ACT-X (JPMJAX23C4). TS was partially supported by JSPS KAKENHI (24K02905) and JST CREST (JPMJCR2015).
Appendix A Polynomial Transformation
Proof of Proposition 5. We use a thresholding and compactness argument inspired by [CM20].
A.1 Proof for Even Functions $(i)$
We divide the analysis into the following steps.
(i-1): Monomials reducing the information exponent. Define $\tau(f)=\max_{-2\leq t\leq 2}|f(t)|$. This entails that if $|f(t)|\geq\tau(f)$, then we have $|t|>2$.
Consider the following expectation:
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}.$ (A.1)
We evaluate the case when $i$ is even. (A.1) can be lower bounded as
[…]
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.3)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.4)
$\geq$ $~{}\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|\geq 2\tau(f)]\bigg{(}\frac{2\tau(f)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.5)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.6)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{\tau(f)}{2\tau(f)}\bigg{)}^{i}(0^{2}-1)\bigg{]}$ (A.7)
$\geq$ $~{}3\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[|f(t)|\geq 2\tau(f)]-2^{-i}.$ (A.8)
Note that $\mathbb{P}[|f(t)|\geq 2\tau(f)]$ is positive (since $f$ is polynomial) and independent of $i$, while $2^{-i}$ decays to 00 as $i$ increases. Therefore, for sufficiently large $i\in\mathbb{N}$, (A.1) is positive and hence ${\mathrm{IE}}(f^{i})\leq 2$. The subsequent analysis aims to provide an upper bound on $i$.
(i-2): Construction of test function. We introduce the notation $H(\cdot;j)$ which takes any function (in $L^{1}$) and returns its $j$-th Hermite coefficient. We consider the following test function:
$\mathscr{H}(f):=\sum_{i=2}^{\infty}\bigg{(}\frac{H(f^{i};2)}{2^{\frac{i}{2}}(2i-1)^{\frac{iq}{2}}}\bigg{)}^{2}.$ (A.9)
(i-3): Lower bound of test function via compactness. Let $\mathcal{F}_{q}$ be a set of polynomials with degree up to $q$ with unit $L^{2}$ norm. Because $\mathscr{H}(f)$ is positive for any $f\in\mathcal{F}_{q}$, $H(f^{i};2)$ is continuous with respect to $f$, and $\mathcal{F}_{q}$ is a compact set, $\inf_{f\in\mathcal{F}_{q}}\mathscr{H}(f)$ admits a minimum value $\mathscr{H}_{0}$ which is positive.
(i-4): Conclusion via hypercontractivity. Because $f$ is a polynomial with degree at most $q$, Gaussian hypercontractivity [O’D14] yields that
$2H(f^{i};2)^{2}\leq\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}(f(t))^{2i}\big{]}\leq(2i-1)^{iq}\big{(}\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}f(t)^{2}\big{]}\big{)}^{i}=(2i-1)^{iq}.$ (A.10)
Therefore, for all polynomials in $\mathcal{F}_{q}$, a partial sum of (A.9) is uniformly bounded by
$\bigg{|}\sum_{i=j}^{\infty}\bigg{(}\frac{H(f^{i};2)}{2^{\frac{i}{2}}(2i-1)^{\frac{iq}{2}}}\bigg{)}^{2}\bigg{|}\leq\sum_{i=j}^{\infty}2^{-i-1}=2^{-j}\to 0\quad(j\to\infty).$ (A.11)
Combining this with the fact that $\mathscr{H}(f)\geq\mathscr{H}_{0}>0$, we know that there exists some $C_{q}\leq 1+\log_{2}(\mathscr{H}_{0}^{-1})$ such that
$\sum_{i=2}^{C_{q}}\bigg{(}\frac{H(f^{i};2)}{2^{\frac{i}{2}}(2i-1)^{\frac{iq}{2}}}\bigg{)}^{2}>\frac{1}{2}\mathscr{H}_{0}>0,$ (A.12)
for all polynomials in $\mathcal{F}_{q}$. This means that there is at least one $i\leq C_{q}$ such that $H(f^{i};2)\neq 0$.
A.2 Proof for Non-even Functions $(ii)$
(ii-1): Monomials reducing the information exponent. We prove that some exponentiation of $g:=f^{2}$ has non-zero first Hermite coefficient. Denote $g^{\mathrm{odd}}$ as the odd part of $g$, and similarly $g^{\mathrm{even}}$. Let $\upsilon(g)\in\mathbb{R}_{+}$ be the value at which the followings hold:
  • (a)
    $g^{\mathrm{odd}}(t)>0$ for all $t\geq\upsilon(g)$ and $g^{\mathrm{odd}}(t)<0$ for all $t\leq-\upsilon(g)$.
  • (b)
    $g^{\mathrm{even}}(t)>|g^{\mathrm{odd}}(t)|$ for all $t\geq\upsilon(g)$ and $t\leq-\upsilon(g)$.
  • (c)
    For for all $t\geq\upsilon(g)$ and $t\leq-\upsilon(g)$, $g(s)=g(t)$ (as an equation of $s$) only has two real-valued solutions with opposing signs.
Such threshold $\upsilon(g)$ exists because the tail of $g=f^{2}$ is dominated by the highest degree which is even. Then, we let $\tau(g)=\max_{-\upsilon(g)\leq t\leq\upsilon(g)}|g(t)|$.
Consider the following expectation:
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\bigg{(}\frac{g(t)}{2\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.13)
(A.13) is decomposed as
[…]
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[2\tau(g)\leq|g(t)|<3\tau(g)]\bigg{(}\frac{g(t)}{3\tau(f)}\bigg{)}^{i}t\bigg{]}$ (A.15)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.16)
We first evaluate the first term. Because of (c), $g(t)=3\tau(f)$ has two real-valued solutions $\alpha<0<\beta$. Because of (a) and (b), $g(\beta)=g^{\mathrm{even}}(\beta)+g^{\mathrm{odd}}(\beta)=3\tau(f)>g^{\mathrm{even}}(-\beta)+g^{\mathrm{odd}}(-\beta)=g^{\mathrm{odd}}(-\beta)$. Because $\lim_{t\to-\infty}g^{\mathrm{odd}}(t)=+\infty$, and $\alpha$ is the only solution in $t<0$, we have $\alpha<-\beta$. Moreover, for all $t>\beta$, we have $g(t)=g^{\mathrm{even}}(t)+g^{\mathrm{odd}}(t)>g^{\mathrm{even}}(-t)+g^{\mathrm{odd}}(-t)=g^{\mathrm{odd}}(-t)$. Combining the above, the first term of (A.16) is bounded as
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|g(t)|\geq 3\tau(f)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.17)
$=\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\beta\leq t\leq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.18)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\leq\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.19)
$=\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[\beta\leq t\leq-\alpha]t\big{]}+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}-\bigg{(}\frac{g(-t)}{3\tau(g)}\bigg{)}^{i}\bigg{)}t\bigg{]}$ (A.20)
$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\beta\leq t\leq-\alpha\big{]}.$ (A.21)
Following the exact same reasoning, we know that the second term of (A.16) is positive. Finally, the third term which is bounded by
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\big{]}\geq-\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.22)
Putting things together,
(A.16)$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\beta\leq t\leq-\alpha\big{]}-\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.23)
The first term is independent of $i$ and positive, while the second term goes to zero as $i$ grows. Therefore, there exists some $i$ such that $\mathrm{IE}(g^{i};1)=1$.
(ii-2): Construction of test function. This time we consider the following function:
$\mathscr{H}(f):=\sum_{i=2}^{\infty}\bigg{(}\frac{H(f^{i};1)}{2^{\frac{i}{2}}(2i-1)^{\frac{iq}{2}}}\bigg{)}^{2}.$ (A.24)
(ii-3): Lower bound of test function via compactness. Let $\mathcal{F}_{q}$ be a set of unit $L^{2}$-norm polynomials with degree up to $q$ and $\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{\mathrm{odd}}(t)^{2}]\geq c$. Since $\mathscr{H}(f)$ is always positive for $\mathcal{F}_{q}$, $\mathscr{H}(f)$ is continuous with respect to $f$, and $\mathcal{F}_{q}$ is a compact set, $\inf_{f\in\mathcal{F}_{q}}\mathscr{H}(f)$ has the minimum value $\mathscr{H}_{0}$ that is positive. Note that $\mathscr{H}(f)$ might depends on $c$.
(ii-4): Conclusion via hypercontractivity. Using the same argument as in (i), we conclude that there exists some $C_{q,c}$ such that
$\sum_{i=2}^{C_{q}}\bigg{(}\frac{H(f^{i};1)}{2^{i}(2i-1)^{\frac{iq}{2}}}\bigg{)}^{2}>\frac{1}{2}\mathscr{H}_{0}>0.$ (A.25)
Because $\mathscr{H}_{0}$ depends on $c$, $C_{q,c}$ depends on $c$ as well as $q$. ∎
Appendix B SGD with Reused Batch
In this section we establish the statistical and computational complexity of Algorithm 1. Recall that the algorithm first trains the first-layer parameters with $T_{1}$ steps of SGD update, where we reuse the same sample for two consecutive steps. The analysis of first-layer training is divided into two phases: $(i)$ weak recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\geq\varepsilon$), and $(ii)$ strong recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\geq 1-\varepsilon$). We then train the second-layer parameters after strong recovery is achieved.
The section is organized as follows.
  • Section B.1 verifies the conditions on the activation function $\sigma$ to guarantee weak and strong recovery.
  • Section B.2 isolates a (nearly) constant fraction of neurons at initialization with an alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ above a certain threshold. We focus on such neurons in Phase I of first-layer training.
  • Section B.3 lower bounds the expected update of alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ of two gradient steps, and Section B.4 shows that neurons yield weak recovery within $2T_{1,1}=\tilde{O}(d)$ steps.
  • Section B.5 discusses how to convert weak recovery to strong recovery using $2T_{1,2}=\tilde{O}(d\varepsilon^{-2})$ SGD steps.
  • Finally, Section B.6 analyzes the second-layer training and concludes the proof.
In the following proofs, we introduce constants $c_{i}$ and $C_{i}$, which depends on $d$ at most polylogarithmically. Specifically, the asymptotic strength of the constants is ordered as follows.
$1\simeq C_{1}\lesssim c_{1}^{-1}\lesssim C_{2}\lesssim C_{3}\lesssim c_{2}^{-1}$ (B.1)
$c_{\eta}$ in the main text can be taken as $c_{\eta}=c_{1}$, where $c_{1}$ should satisfy $\lim_{d\to\infty}c_{1}=0$, but the convergence can be arbitrarily slow. This requirement comes from the fact that we do not know the exact value of $H(\sigma_{*}^{I};k_{*})$, which might be very small. To ensure that the signal is isolated, taking $\eta\asymp c_{1}d^{-1}$ with arbitrarily slow $c_{1}$ suffices. $C_{2}$ can also be arbitrarily slow, as long as it satisfies $C_{2}=\mathrm{poly}(c_{1}^{-1})$. $C_{3}=\mathrm{polylog}(d)$ will be used to represent polylogarithmic factor that comes from high probability bounds.
B.1 Conditions on the Activation Function
B.1.1 Verifying Assumption 2
In the following, we focus on the activation function of a single neuron and omit the subscript that distinguishes different neurons. Recall that we consider polynomial activation functions written as
$\sigma(z):=\sum_{j=0}^{C_{\sigma}}\beta_{j}\mathsf{He}_{j}(z).$ (B.2)
For weak recovery, we can use any polynomial that has degree $C_{\sigma}\geq C_{q}$ as long as the following condition holds: If $\mathrm{IE}(\sigma_{*})\geq 2$ and there exists some $i\leq C_{\sigma}$ such that $\mathrm{IE}(\sigma_{*}^{i})=1$, $\sigma$ should satisfy
$H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}\neq 0.$ (B.3)
If $\mathrm{IE}(\sigma_{*})\geq 3$ and there does not exist any $i\leq C_{\sigma}$ such that $\mathrm{IE}(\sigma_{*}^{i})=1$ (in this case there exists some $i\leq C_{q}$ such that $\mathrm{IE}(\sigma_{*}^{i})=2$), $\sigma$ should satisfy
$H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};1\bigg{)}\neq 0.$ (B.4)
Below we prove Lemma 2 which shows that the above conditions are met with probability $1$ for randomly drawn the Hermite coefficients.
Proof of Lemma 2. We note that $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)=\mathbb{E}[\sigma^{(i)}(\sigma^{(1)})^{i-1}\mathsf{He}_{k}]$ is a polynomial of $\{\beta_{j}\}_{j=0}^{C_{\sigma}}$. This polynomial is not identically equal to zero. To confirm this, consider $\sigma=x^{C_{\sigma}}+x^{C_{\sigma}-1}$. Because $\sigma^{(i)}(\sigma^{(1)})^{i-1}$ is expanded as a sum of $x^{l}(i(C_{\sigma}-3)\leq l\leq i(C_{\sigma}-2)+1$ with positive coefficients and each $x^{l}$ is a sum of $\mathsf{He}_{l},\mathsf{He}_{l-2}\cdots$ with positive coefficients, $\sigma^{(i)}(\sigma^{(1)})^{i-1}$ has all positive Hermite coefficients for degree $10,1,\cdots,i(C_{\sigma}-2)+$ 1. If $k\leq i(C_{\sigma}-2)+1$, this choice of $\sigma$ yields $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)>0$, which confirms that $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)$ as a polynomial of $\{\beta_{j}\}_{j=0}^{C_{\sigma}}$ is not identically equal to zero.
Now, the assertion follows from the Schwartz–Zippel Lemma [Sch80], or the fact that zeros of a non-zero polynomial form a measure-zero set. ∎
B.1.2 Verifying Assumption 3
On the other hand, for the strong recovery we require an additional condition on the activation function due to link misspecification, which is also introduced in [BAGJ21, MHWSE23]:
$\sum_{j=p}^{p}j!\alpha_{j}\beta_{j}s^{j}>0\quad\text{for all $s>0$}.$ (B.5)
In order to meet Assumption 2 and (3) simultaneously, we follow [OSSW24] and randomize the activation function. Specifically, the activation function should satisfy
  • (I)
    If $\mathrm{IE}(\sigma_{*})=1$. We require $\beta_{1}>0$ and $\sum_{j=1}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
  • (II)
    Else if $\mathrm{IE}((\sigma_{*})^{I})=1$ for some $2\leq I\leq C_{\sigma}$. We require $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};1)$. Also, $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
  • (III)
    Else if $\mathrm{IE}(\sigma_{*})=2$. We require $\beta_{2}>0$ and $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
  • (IV)
    Else $\mathrm{IE}((\sigma_{*})^{I})=2$ for some $2\leq I\leq C_{q}$. We require $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};2)$. Also, $\sum_{j=3}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
Now we prove Lemma 3 which verifies the existence of an activation function that satisfies the assumptions above with non-zero probability. The construction does not depend on the link function itself, but only its degree $q$.
Proof of Lemma 3. Let $c$ be a sufficiently small constant, and $C_{\sigma}$ be the minimum odd integer with $C_{\sigma}\geq\max\{C_{q}+1,q+2,3\}$. With probability $1/2$, we choose the coefficients as $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$, and $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $2\leq j\leq C_{\sigma}$. Then, it is easy to see (I) and (III) are met with probability at least $2^{-q}$, because they hold when $\mathrm{sign}(\alpha_{j})=\mathrm{sign}(\beta_{j})$ holds in the summation.
By taking $c$ sufficiently small, we have
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)=\underbrace{i!\beta_{I}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c).$ (B.6)
When $I$ is even, by adjusting the sign of $\beta_{1}$, $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};1)$ with probability $\frac{1}{2}$. Note that the sign of $\beta_{1}$ is independent from whether $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds. This holds with probability at least $2^{-q+1}$. Thus we verified (II) for even $I$.
In the same vein, we can verify (IV) for even $I\leq C_{q}\leq C_{\sigma}-1$. We have
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)=\underbrace{(I+1)!\beta_{I+1}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}),$ (B.7)
and (IV) can be verified using the same argument.
Otherwise (also with probability $1/2$), we choose the coefficients as $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}=\pm 1$ to verify (II) and (IV) for odd $I$. It is easy to see that $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds for (I) and $\sum_{j=3}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ for (III). In addition, $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};0)>0$ and $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};1)>0$. Therefore, by taking $c$ sufficiently small, flipping the sign of $\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1}$ can change the sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};k)$ for both (II) and (IV) with odd $I$. Combining all cases yields the desired claim. ∎
B.2 Random Initialization
In Section B.3.1 we focus on the neurons with slightly larger initial alignment that satisfy $\kappa^{0}_{j}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}$ at initialization, where constant $C_{2}$ grows at most polylogarithmically in $d$. The following lemma states that roughly a constant portion of the neurons satisfies this initial alignment condition.
Lemma 7.
At the time of initialization, $\kappa^{0}_{j}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{0}$ satisfies the following:
$\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]=\mathbb{P}[\kappa^{0}_{j}\leq-2C_{2}d^{-\frac{1}{2}}]\gtrsim e^{-16C_{2}^{2}}=\tilde{\Omega}(1).$ (B.8)
We make use of the following lemma.
Lemma 8 (Theorem 2 of [CCM11]).
For any $\beta>1$ and $s\in\mathbb{R}$, we have
$\frac{\sqrt{2e(\beta-1)}}{2\beta\sqrt{\pi}}e^{-\frac{\beta s^{2}}{2}}\leq\int_{s}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{t^{2}}{2}}\mathrm{d}t$ (B.9)
Proof of Lemma 7. Because $\kappa^{0}=v^{\top}w\overset{\mathrm{d}}{=}\frac{{\sf e}_{1}^{\top}g}{g}$, where $g\sim\mathcal{N}(0,\boldsymbol{I}_{d})$,
$\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$ $=\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\sf e}_{1}^{\top}g\geq 4C_{2}\land\|g\|\leq 2d^{\frac{1}{2}}\bigg{]}$ (B.10)
$\geq\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\sf e}_{1}^{\top}g\geq 4C_{2}\bigg{]}-\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\|g\|\geq 2d^{\frac{1}{2}}\bigg{]}$ (B.11)
$\gtrsim\frac{\sqrt{2e(\beta-1)}}{2\beta\sqrt{\pi}}e^{-8\beta C_{2}^{2}}-e^{-\Omega(d)}.$ (B.12)
By letting $\beta=2$, we have that $\mathbb{P}[\kappa^{0}_{j}\geq C_{2}d^{-\frac{1}{2}}]\gtrsim e^{-16C_{2}^{2}}.$ Due to symmetry, $\mathbb{P}[\kappa^{0}_{j}\leq 2C_{2}d^{-\frac{1}{2}}]=\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$. ∎
B.3 Population Update
We first analyze the training of first-layer parameters by evaluating the expected (population) update of two gradient steps with the same training example. At each step, the parameters are updated as
$\boldsymbol{w}^{t+1}_{j}$ $\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\big{(}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2}\big{)}$ (B.13)
$=\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}^{2}+2\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}.$ (B.14)
While the second term scales with $\eta^{t}a_{j}^{2}=\eta^{t}c_{a}^{2}$, the third term scales with $\eta^{t}a_{j}=\eta^{t}c_{a}$. Thus, by setting the second-layer scale $c_{a}$ sufficiently small, we can ignore the interaction of neurons; similar mechanism also appeared in [BES+22, AAM23]. Specifically, in the following, we show that the strength of the signal is $\boldsymbol{\theta}^{\top}\boldsymbol{w}_{j}^{t}\gtrsim d^{-\frac{1}{2}}$. Thus, by simply letting $c_{a}\lesssim C_{3}^{-1}d^{-\frac{1}{2}}$, we can ignore the effect of the squared term. Thus we may focus on the following correlational update:
$\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}+\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}.$ (B.15)
Due to the absence of interaction between neurons, we omit the subscript $j$ for the index of neurons and ignore the prefactor of $N$ (which can be absorbed into the learning rate); multiplying $N$ to $\eta^{t}$ specified below recovers the scaling of $\eta^{t}$ presented in the main text.
B.3.1 When $\mathrm{IE}(\sigma_{*}^{I})=2$ with $I\geq 2$
First we consider the most technically difficult case, when $\mathrm{IE}[\sigma_{*}]\geq 3$ and the information exponent cannot be lowered to 1 for $i\leq C_{\sigma}$; in this case, from Proposition 5 we know that there exists some $2\leq i\leq C_{\sigma}$ such that $\mathrm{IE}[\sigma_{*}^{i}]=2$ and we let $I$ be the first such $i$.
[…]
Lemma 9.
Starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $C_{2}d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot\eta\tilde{\eta}c_{1}^{I-1}H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};1\bigg{)}H\big{(}(\sigma_{*})^{I};2\big{)}\kappa+\eta\tilde{\eta}\nu,$ (B.16)
where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
Proof.  We first compute one gradient step from $\boldsymbol{w}=\boldsymbol{\omega}$ with a fresh sample $(\boldsymbol{x},y)$.
~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{x}.$ (B.17)
Then, with a projection matrix $\boldsymbol{P}_{\boldsymbol{\omega}}=I-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}$, the updated parameter becomes
$\boldsymbol{w}\leftarrow\boldsymbol{w}+\boldsymbol{P}_{\boldsymbol{\omega}}\eta y\sigma^{\prime}(\boldsymbol{w}^{\top}\boldsymbol{x})\boldsymbol{x}=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x},$ (B.18)
and the next gradient step with the same sample is computed as
~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})$ $=\eta y\sigma^{\prime}(\boldsymbol{w}^{\top}\boldsymbol{x})\boldsymbol{x}$ (B.19)
$=y\sigma^{\prime}\big{(}(\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})^{\top}\boldsymbol{x}\big{)}\boldsymbol{x}$ (B.20)
$=y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{x},$ (B.21)
here we used the notation $\|\boldsymbol{v}\|_{\boldsymbol{A}}^{2}=\boldsymbol{v}^{\top}\boldsymbol{A}\boldsymbol{v}$ for a vector $\boldsymbol{v}\in\mathbb{R}^{d}$ and a positive symmetric matrix $\boldsymbol{A}\in\mathbb{R}^{d\times d}$. From (B.18) and (B.21), the parameter after the two steps is obtained as
$\boldsymbol{w}$ $\leftarrow\boldsymbol{w}+\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})$ (B.22)
$=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+\eta y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}$ (B.23)
$=\boldsymbol{\omega}+\eta\boldsymbol{g},$ (B.24)
where we defined
$\boldsymbol{g}:=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.25)
Finally, normalization yields
$\boldsymbol{w}$ $\leftarrow\frac{\boldsymbol{w}-\xi(\boldsymbol{w}-\boldsymbol{\omega})}{\|\boldsymbol{w}-\xi(\boldsymbol{w}-\boldsymbol{\omega})\|}=\frac{\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}.$ (B.26)
Therefore, the update of the alignment is
$\boldsymbol{\theta}^{\top}\boldsymbol{w}$ $=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{(1+\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2})^{\frac{1}{2}}}$ (B.27)
$\geq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}.$ (B.28)
On the other hand, we have
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}.$ (B.29)
We evaluate the expectation of (B.28). For the $j$-th Hermite polynomial $\mathsf{He}_{j}$ and $\boldsymbol{u}\in\mathbb{S}^{d-1}$, we have that
$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{i}({\sf e}_{1}^{\top}\boldsymbol{x})f(\boldsymbol{u}^{\top}\boldsymbol{x}){\sf e}_{1}^{\top}\boldsymbol{x}]$ (B.30)
$=j({\sf e}_{1}^{\top}x)^{j-1}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j-1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]+({\sf e}_{1}^{\top}\boldsymbol{x})^{j+1}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})],$ (B.31)
$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{j}(x_{1})f(\boldsymbol{u}^{\top}\boldsymbol{x}){\sf e}_{2}^{\top}\boldsymbol{x}]=({\sf e}_{1}^{\top}\boldsymbol{x})^{j}({\sf e}_{2}^{\top}\boldsymbol{x})\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})].$ (B.32)
Therefore,
$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{j}({\sf e}_{1}^{\top}\boldsymbol{x})f(\boldsymbol{u}^{\top}\boldsymbol{x})\boldsymbol{x}]$ (B.33)
$=\begin{pmatrix}j({\sf e}_{1}^{\top}\boldsymbol{x})^{j-1}\\ 0\\ \vdots\\ 0\end{pmatrix}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j-1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]{\sf e}_{1}+({\sf e}_{1}^{\top}\boldsymbol{x})^{j}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]\boldsymbol{u}.$ (B.34)
Hence the first term of $\boldsymbol{g}$ (B.25) can be exapanded as
$\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ $=\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}\bigg{(}\sum_{j=p}^{q}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}j\sum_{j=0}^{C_{q}}\beta_{j}\mathsf{He}_{j}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\bigg{)}\boldsymbol{x}\bigg{]}$ (B.35)
$=\boldsymbol{P}_{\boldsymbol{\omega}}\sum_{j=p}^{q}\bigg{[}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j}\boldsymbol{\omega}\bigg{]}$ (B.36)
$=\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}.$ (B.37)
The coefficient is evaluated as
$\bigg{|}\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\bigg{|}\lesssim\kappa^{p-1}\leq\kappa^{2}.$ (B.38)
For the second term of $\boldsymbol{g}$ (B.25), we first bound the difference in replacing $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ with $d$,
$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}$ (B.39)
$=2\big{|}\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)h(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2},\boldsymbol{\omega}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})]\big{|},$ (B.40)
where $h$ is a polynomial with degree at most $(C_{q}+q)^{C_{q}-1}+q+1$ and coefficients are all $O(1)$. (B.40) is further upper bounded as
(B.40)$\leq 2\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)^{2}]^{\frac{1}{2}}\mathbb{E}[(h(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2},\boldsymbol{\omega}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}))^{2}]^{\frac{1}{2}},$ (B.41)
by Cauchy-Schwarz inequality. $\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)^{2}]^{\frac{1}{2}}=\eta(2d-1)^{\frac{1}{2}}$, and the expectation of $\mathbb{E}[h^{2}]$ is $O(1)$ when $\eta\leq d^{-1}$. Therefore, (B.40) is bounded by $C_{1}\eta d^{\frac{1}{2}}$.
Now, we consider $\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$. The following decomposition can be made.
$\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.42)
$=\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.43)
$=\sum_{k=0}^{j-1}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose k}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.44)
We evaluate each term of (B.44) except for $k=I-1$. Each term of (B.44) is a constant multiple of $\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}$. and we can evaluate the constant by
$\bigg{|}\boldsymbol{\theta}^{\top}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.45)
$=(\eta d)^{k}\bigg{|}\mathbb{E}\bigg{[}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{k}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}^{k+1}\mathsf{He}_{j-k-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}.$ (B.46)
When $k\leq I-2$, $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{k+1}$ has information exponent larger than $2$. Therefore, we have
[…]
$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{k}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{k-l+1}\mathsf{He}_{j-k-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.48)
$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\sum_{m=3}^{\infty}m!\kappa^{q(k-l+1)}H\big{(}(\sigma_{*})^{k-l+1};m\big{)}H\big{(}(\sigma^{\prime})^{k}\mathsf{He}_{j-k-1};m-1\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}\bigg{|}$ (B.49)
$\lesssim(\eta d)^{k}\kappa^{2}.$ (B.50)
When $k\geq I$, we know that $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{k+1},\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ have information exponent larger than $1$, and $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1},\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))$ have information exponent larger than $2$. Thus, the expectation in (B.46) is bounded by
[…]
$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\sum_{m=2}^{\infty}m!\kappa^{q(k-l+1)}H\big{(}(\sigma_{*})^{k-l+1};m-1\big{)}H\big{(}(\sigma^{\prime})^{k}\mathsf{He}_{j-k-1};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}\bigg{|}$ (B.52)
$\lesssim(\eta d)^{k}\kappa.$ (B.53)
Now, consider the case when $k=I-1$:
$\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}\big{)}^{I}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.54)
$=\sum_{l=0}^{I}\sum_{j=1}^{C_{\sigma}}(\eta d)^{k}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.55)
Note that $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1}$ has information exponent larger than $2$. Therefore, for $l\geq 1$, we have
$\bigg{|}\boldsymbol{\theta}^{\top}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.56)
$\lesssim(\eta d)^{I-1}\bigg{|}\mathbb{E}\bigg{[}\sum_{m=3}^{\infty}m!\kappa^{m-1}H\big{(}\mathsf{He}_{j-I}(\sigma^{\prime})^{I-1};m-1\big{)}H\big{(}\big{(}\sigma_{*}\big{)}^{I-l};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.57)
$\lesssim(\eta d)^{I-1}\kappa^{2}.$ (B.58)
And for $l=0$, $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ has information exponent $2$ and we have
$\sum_{j=1}^{C_{\sigma}}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}\mathsf{He}_{j-I}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.59)
$=(\eta d)^{I-1}\sum_{m=2}^{\infty}m!\kappa^{m-1}H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};m-1\bigg{)}H\big{(}(\sigma_{*})^{I};m\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$ (B.60)
If $H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};1\bigg{)}\neq 0$, we have
(B.60)$=(1+O(\kappa))\cdot(\eta d)^{I-1}\underbrace{2H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};1\bigg{)}H\big{(}(\sigma_{*})^{I};2\big{)}}_{=:\gamma}\kappa\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}.$ (B.61)
We have that
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.62)
$=(1+O(\kappa))\cdot(\eta d)^{I-1}\gamma\kappa\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+O(\kappa^{2}+C_{1}\eta d^{\frac{1}{2}}+(\eta d)^{I-2}\kappa^{2}+(\eta d)^{I}\kappa+(\eta d)^{I-1}\kappa^{2})$ (B.63)
When $d^{-\frac{1}{2}}C_{3}\leq\kappa\leq c_{2}$ and $\eta=c_{1}d^{-1}$, we have
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa$ (B.64)
Together with the bound on the first term ((B.37) and (B.38)), the expectation of $\boldsymbol{g}$ (B.25) is evaluated as
$\boldsymbol{\theta}^{\top}\mathbb{E}[\boldsymbol{g}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa.$ (B.65)
By using this, the expected update of the alignment becomes
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\mathbb{E}[\|\boldsymbol{g}\|^{2}]-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}].$ (B.66)
Note that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\kappa\lesssim c_{1}$. Thus,
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa-C_{1}\eta\tilde{\eta}^{2}(\kappa\eta d+\eta^{2}\tilde{\eta}d).$ (B.67)
When $\tilde{\eta}\leq c_{2}$, $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ is evaluated as
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa.$ (B.68)
In the same way, using (B.29), we also have the opposite bound:
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\leq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa$ (B.69)
Regarding the noise, recall that
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}.$ (B.70)
$\eta^{-1}\tilde{\eta}^{-1}\big{(}\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}[\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}]\big{)}=\ \boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$ is a mean-zero polynomial of Gaussian inputs, with all coefficients and variances of inputs bounded by $O(1)$. Notice that normalization does not increase the absolute value of the noise. Thus, regarding $\nu=\eta^{-1}\tilde{\eta}^{-1}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{w}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\big{)}$, we have that
$\mathbb{P}\big{[}\big{|}\eta^{-1}\tilde{\eta}^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{w}-\mathbb{E}[\eta^{-1}\tilde{\eta}^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{w}]\big{|}>t\big{]}\leq\mathbb{P}\bigg{[}\bigg{|}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}\bigg{[}\boldsymbol{\theta}^{\top}\boldsymbol{g}\bigg{]}\bigg{|}>t\bigg{]}\leq\exp(-t^{1/C_{1}}/C_{1}).$ (B.71)
This completes the proof. ∎
B.3.2 When $\mathrm{IE}(\sigma_{*}^{I})=1$ with $I\geq 2$
[…]
Lemma 10.
For the case of $\mathrm{IE}(\sigma_{*}^{I})=1$, starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot\eta\tilde{\eta}c_{1}^{I-1}H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}H\big{(}(\sigma_{*})^{I};1\big{)}+\eta\tilde{\eta}\nu,$ (B.72)
where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
Proof.  Similarly to Lemma 9, update of the alignment is evaluated as
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\geq\kappa+\eta c_{2}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}$ (B.73)
and
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta c_{2}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2},$ (B.74)
where
$\boldsymbol{g}=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.75)
From (B.37) and (B.38) we know that the first term of $\boldsymbol{g}$ (B.75) is evaluated as
$\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta},$ (B.76)
where the sum of coefficients are bounded by $\lesssim\kappa$.
We then consider the second term of $\boldsymbol{g}$ (B.75). We can replace $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ by $d$ with the following bound similarly to (B.40):
$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}\leq C_{1}\eta d^{\frac{1}{2}}.$ (B.77)
For the term $\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$. the following decomposition can be made.
$\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.78)
$=\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.79)
$=\sum_{k=0}^{j-1}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose k}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.80)
We evaluate each term of (B.80) except for $k=I-1$. Similarly to the bounds on (B.46), each term is a constant multiple of $\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$, and we want to bound
$\bigg{|}\boldsymbol{\theta}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}.$ (B.81)
When $k\leq I-2$, $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{k+1}$ has information exponent larger than $1$. Therefore, we have
(B.81)$\lesssim(\eta d)^{k}\kappa.$ (B.82)
When $k\geq I$, we have
(B.81)$\lesssim(\eta d)^{k}\leq c_{1}^{I}.$ (B.83)
Now we consider the case when $k=I-1$:
$\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}\big{)}^{I}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.84)
$=\sum_{l=0}^{I}\sum_{j=1}^{C_{\sigma}}(\eta d)^{k}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.85)
Note that $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1}$ has information exponent larger than $1$. Therefore, for $l\geq 1$, we have
$\bigg{|}\boldsymbol{\theta}^{\top}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.86)
$\lesssim(\eta d)^{I-1}\bigg{|}\mathbb{E}\bigg{[}\sum_{m=2}^{\infty}m!\kappa^{m-1}H\big{(}\mathsf{He}_{j-I}(\sigma^{\prime})^{I-1};m-1\big{)}H\big{(}\big{(}\sigma_{*}\big{)}^{I-l};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.87)
$\lesssim(\eta d)^{I-1}\kappa.$ (B.88)
And for $l=0$, $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ has information exponent $1$ and we have
$\sum_{j=1}^{C_{\sigma}}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}\mathsf{He}_{j-I}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.89)
$=(\eta d)^{I-1}\sum_{m=1}^{\infty}m!\kappa^{m-1}H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};m-1\bigg{)}H\big{(}(\sigma_{*})^{I};m\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$ (B.90)
If $H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}\neq 0$, we have
(B.90)$=(1+O(\kappa))\cdot(\eta d)^{I-1}\underbrace{H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}H\big{(}(\sigma_{*})^{I};1\big{)}}_{=:\gamma}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}.$ (B.91)
Note that
$H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}=H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}.$ (B.92)
Now we have that
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.93)
$=(1+O(\kappa))\cdot(\eta d)^{I-1}\gamma\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+O(\kappa+C_{1}\eta d^{\frac{1}{2}}+(\eta d)^{I-2}\kappa+(\eta d)^{I}+(\eta d)^{I-1}\kappa).$ (B.94)
When $d^{-\frac{1}{2}}C_{3}\leq\kappa\leq c_{2}$ and $\eta=c_{1}d^{-1}$, we have
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma$ (B.95)
Together with the bound on the first term ((B.76)), the expectation of $\boldsymbol{g}$ is evaluated as
$\boldsymbol{\theta}^{\top}\mathbb{E}[\boldsymbol{g}]=(1+O(c_{1}))\cdot c_{1}^{I-1}.$ (B.96)
By using this and (B.73), the expected update of the alignment becomes
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\mathbb{E}[\|\boldsymbol{g}\|^{2}]-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}].$ (B.97)
Note that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\kappa\lesssim c_{1}$. Taking $\tilde{\eta}\leq c_{2}$ yields that
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma.$ (B.98)
The upper bound can be obtained in similar fashion,
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\leq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma.$ (B.99)
Finally, the noise can be handled in the exact same way as that of Lemma 9, the details of which we omit. ∎
B.3.3 When $\mathrm{IE}(\sigma_{*}^{I})=1,2$ with $I=1$
We finally consider the case when $\mathrm{IE}[\sigma_{*}]=1$ or when $\mathrm{IE}[\sigma_{*}]=2$ and $\mathrm{IE}[\sigma_{*}^{i}]\geq 2$ for $i=2,\cdots,C_{\sigma}$.
Lemma 11.
Starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then for $\mathrm{IE}[\sigma_{*}]=1$, the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot 2\eta\tilde{\eta}\alpha_{1}\beta_{1}+\eta\tilde{\eta}\nu,$ (B.100)
and when $\mathrm{IE}[\sigma_{*}]=2$ and $\mathrm{IE}[\sigma_{*}^{i}]\geq 2$ for all $i=2,\cdots,C_{q}$,
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot 4\eta\tilde{\eta}\alpha_{2}\beta_{2}\kappa+\eta\tilde{\eta}\nu,$ (B.101)
where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
Proof.  Similar to Lemma 9, the update of the alignment is
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\geq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}$ (B.102)
and
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}$ (B.103)
where
$\boldsymbol{g}=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.104)
We first consider the case when $\mathrm{IE}[\sigma_{*}]=1$. We have
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=\alpha_{1}\beta_{1}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}=\alpha_{1}\beta_{1}(1-\kappa^{2}).$ (B.105)
Because we take $\eta=c_{1}d^{-1}$ with a vanishing constant $c_{1}$ and assume that $\kappa\leq c_{2}$, we have
$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}\lesssim\eta d\beta_{1}^{2}\lesssim c_{1}.$ (B.106)
and
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]=2(1+O(c_{1}))\cdot\alpha_{1}\beta_{1}.$ (B.107)
The first claim follows from the fact that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\eta=c_{1}d^{-1}$.
Next we consider the case when $\mathrm{IE}[\sigma_{*}^{i}]=2$ for $i=1,2,\cdots,C_{q}$. We have
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=2\alpha_{2}\beta_{2}\kappa\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}=2\alpha_{2}\beta_{2}(1-\kappa^{2})\kappa.$ (B.108)
On the other hand, because $y,\cdots,y^{C_{\sigma}}$ has information exponent larger than $1$, we have
$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}\lesssim\eta d\kappa=c_{1}\kappa.$ (B.109)
Thus, when $\kappa\leq c_{1}$,
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]=(1+O(c_{1}))\cdot 4\alpha_{2}\beta_{2}\kappa,$ (B.110)
which establishes the second claim. ∎
B.4 Stochastic Update
As a result of the previous subsection, by choosing either of $\xi_{j}=1-\tilde{\eta}_{j}$ or $\xi_{j}=1+\tilde{\eta}_{j}$, we obtained that
  • (i)
    When $\mathrm{IE}[\sigma_{*}^{i}]=1$ for some $i\leq C_{\sigma}$, the update can be written as
    $\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\gamma+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}\leq\kappa^{2(t+1)}_{j}\leq\kappa^{2t}_{j}+2\eta^{2t}_{j}\tilde{\eta}_{j}\gamma+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}.$ (B.111)
  • (ii)
    Otherwise, the update can be written as
    $\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\gamma\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}\leq\kappa^{2(t+1)}_{j}\leq\kappa^{2t}_{j}+2\eta^{2t}_{j}\tilde{\eta}_{j}\gamma\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}.$ (B.112)
Here $\eta_{j}^{2t}=\eta_{j}^{2t+1}=c_{1}d^{-1}$, $\tilde{\eta}_{j}\leq c_{2}$, and $\gamma\geq c_{2}$ that depends on $\sigma_{*}$. $\nu^{2t}_{j}$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu^{2t}_{j}|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$. We assumed that $d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for the former and $C_{2}d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for the latter. For each neuron $j$, we sample $T_{1,1,j}\sim\mathrm{Unif}(\{1,\cdots,T_{1,1}\})$ and let $\eta^{t}=c_{1}d^{-1}$. For $t=0,\cdots,2(T_{1,1,j}-1)$, we let $\xi_{j}^{t}\equiv 1-\tilde{\eta}_{j}$ or $\xi_{j}^{t}\equiv 1+\tilde{\eta}_{j}$ with equiprobability, where $\tilde{\eta}_{j}=c_{2}$, and $\xi_{j}^{t}=1$ for $t=2T_{1,1,j},\cdots,2(T_{1,1}-1)$.
The goal of this subsection is to prove the following lemma.
Lemma 12.
Let $T_{1}=\tilde{\Theta}(d)$. If the initial alignment satisfies $\kappa_{j}^{0}\geq C_{2}d^{-\frac{1}{2}}$ (for (i)) or $\kappa_{j}^{0}\geq 2C_{2}d^{-\frac{1}{2}}$ (for (ii)), then we have $\frac{1}{4}c_{2}\leq\kappa^{2T_{1,1}}_{j}\leq c_{2}$ for at least $1/\mathrm{polylog}(d)$ fraction of neurons, with high probability.
Proof.
(i) the case of $\mathrm{IE}=1$.
If $d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for all $t=0,1,\cdots,\tau\ (0\leq\tau\leq T_{1,j}-1)$, we have
$\kappa^{2(\tau+1)}_{j}$ $\geq\kappa^{2\tau}_{j}+\eta^{\tau}\tilde{\eta}_{j}\gamma+\eta^{\tau}\tilde{\eta}_{j}\nu^{2\tau}_{j}$ (B.113)
$=\kappa^{0}_{j}+\sum_{s=0}^{\tau}\eta^{s}\tilde{\eta}_{j}\gamma+\sum_{s=0}^{\tau}\eta^{s}\tilde{\eta}_{j}\nu^{2s}_{j}$ (B.114)
$\geq 2d^{-\frac{1}{2}}+c_{1}c_{2}d^{-1}\tilde{\varepsilon}\gamma(s+1)-c_{1}c_{2}^{2}d^{-1}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}_{j}\bigg{|}.$ (B.115)
With high probability, $\sum_{s=0}^{\tau}\nu^{2s}_{j}$ is bounded by $C_{3}\sqrt{\tau+1}$. If $\tau+1\leq 4\gamma^{-2}C_{3}^{2}$, by letting $c_{2}^{2}\leq\frac{1}{2}\gamma c_{1}^{-1}C_{3}^{-2}d^{\frac{1}{2}}$, we have $c_{1}c_{2}d^{-1}C_{3}\sqrt{\tau+1}\leq d^{-\frac{1}{2}}$. If $\tau+1\geq 4\gamma^{-2}C_{3}^{2}$, we have $C_{3}\sqrt{\tau+1}\leq\frac{1}{2}\gamma(\tau+1)$. Thus, in either case,
(B.115)$\geq d^{-\frac{1}{2}}+\frac{1}{2}c_{1}c_{2}^{2}d^{-1}\gamma(\tau+1).$ ) . (B.116)
This verifies that $\kappa^{2t}_{j}\geq C_{1}d^{-\frac{1}{2}}$ holds for $t=\tau+1$. By induction,
$\kappa^{2t}_{j}$ $\geq d^{-\frac{1}{2}}+\frac{1}{2}c_{1}c_{2}^{2}d^{-1}\gamma(\tau+1)$ (B.117)
holds for $t$ until $\kappa^{2t}$ gets larger than $c_{2}$. By letting $T_{1,1}\geq c_{1}^{-1}c_{2}^{-2}\gamma^{-1}d$, we have $\kappa^{2t}\geq\frac{1}{2}c_{2}$ for some $t\leq T_{1,1}$, with high probability.
Now together with (B.113), we have
$\kappa^{2(t+1)}_{j}$ $\leq\kappa^{2t}_{j}+\eta^{t}\tilde{\eta}_{j}\cdot 2\gamma+\eta^{t}\tilde{\eta}_{j}\nu^{2t}_{j}.$ (B.118)
Hence we obtain the following bound with high probability,
$|\kappa^{2(t+1)}_{j}-\kappa^{2t}_{j}|\leq\eta^{t}_{j}\tilde{\eta}_{j}(2\gamma+C_{3})\lesssim c_{1}c_{2}^{2}C_{3}d^{-1}=:\Delta_{1}.$ (B.119)
When $\kappa^{2\tau}\geq\frac{1}{2}c_{1}$ holds for some $\tau$, with high probability, we have
$\frac{1}{4}c_{2}\leq\kappa^{2\tau+s}\leq c_{2}$ (B.120)
for all $0\leq s\leq\Delta_{1}/4c_{2}$. Because $1/4\Delta_{1}c_{2}=\tilde{\Theta}(d)$ and $T_{1,1}$ is also $\tilde{\Theta}(d)$, with probability $\tilde{\Theta}(1)$, $2T_{1,1,j}$ satisfies $2\tau\leq 2T_{1,1,j}\leq 2\tau+1/4\Delta_{1}c_{2}$. This establishes the first assertion.
(ii) the case of $\mathrm{IE}=2$:
If $C_{2}d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for all $t=0,1,\cdots,\tau\ (0\leq\tau\leq T_{1,j}-1)$, we have
$\kappa^{2(\tau+1)}_{j}$ $\geq\kappa^{2\tau}_{j}+\eta^{\tau}\tilde{\eta}_{j}\kappa_{j}^{2\tau}\gamma+\eta^{\tau}\tilde{\eta}_{j}\nu^{2s}_{j}$ (B.121)
$=\kappa^{0}_{j}+c_{1}c_{4}d^{-1}\gamma\sum_{s=0}^{\tau}\kappa_{j}^{2s}+c_{1}c_{2}d^{-1}\sum_{s=0}^{\tau}\nu^{2s}_{j}.$ (B.122)
With high probability, $\sum_{s=0}^{\tau}\nu^{2s}_{j}$ is bounded by $C_{3}\sqrt{\tau+1}$. If $\tau+1\leq 4\gamma^{-2}C_{2}^{-2}C_{3}^{2}d$, by letting $c_{2}\leq\frac{1}{2}\gamma c_{1}^{-1}C_{2}^{2}C_{3}^{-2}$, we have $c_{1}c_{2}d^{-1}C_{3}\sqrt{\tau+1}\leq C_{2}d^{-\frac{1}{2}}$. If $\tau+1\geq 4\gamma^{-2}C_{2}^{-2}C_{3}^{2}d$, we have $C_{3}\sqrt{\tau+1}\leq\frac{1}{2}\gamma(\tau+1)C_{2}d^{-\frac{1}{2}}$. Thus, in either case,
(B.122)$\geq\kappa_{j}^{0}/2+\frac{1}{2}c_{1}c_{4}d^{-1}\tilde{\varepsilon}\gamma\sum_{s=0}^{\tau}\kappa_{j}^{2s}.$ (B.123)
This verifies that $\kappa^{2t}_{j}\geq C_{1}d^{-\frac{1}{2}}$ holds for $t=\tau+1$. By induction,
$\kappa^{2t}$ $\geq\kappa_{j}^{0}/2+\frac{1}{2}c_{1}c_{2}d^{-1}\gamma\sum_{s=0}^{\tau}\kappa_{j}^{2s}$ (B.124)
holds for $t$ until $\kappa^{2t}$ gets larger than $c_{2}$. By letting $T_{1,1}\geq\log_{(1+\frac{1}{2}c_{1}c_{2}d^{-1})}\frac{c_{2}}{C_{1}d^{-\frac{1}{2}}}$, we have $\kappa^{2t}\geq\frac{1}{2}c_{2}$ for some $t\leq T_{1,1}$, with high probability. Similarly to the case of (i), we can verify that $\frac{1}{4}c_{2}\leq\kappa^{2t}\leq c_{2}$ for $\tilde{\Theta}(d)$ steps. Therefore, we obtain the second assertion. ∎
B.5 From Weak Recovery to Strong Recovery
In the previous subsection, we proved that after $t=2T_{1,1}=\tilde{\Theta}(d)$ steps, with probability $\tilde{\Omega}(1)$ over the randomness of initialization, $T_{1,1,j}$, and $\tilde{\eta}_{j}$, neurons achieve small alignment with the target direction $\frac{1}{4}c_{2}\leq\kappa^{2T_{1,1}}_{j}\leq c_{2}$. This subsection discusses how to convert the weak recovery into the strong recovery. We focus on the neurons that satisfy $\alpha_{j}\beta_{j}\geq 0$ for all $j$ as specified in Assumption 3.
For each neuron $j$, we let $c_{a}\eta^{t}=\eta=c_{1}d^{-1}$ if $t$ is even and $\eta^{t}=0$ if $t$ is odd, for $t=2T_{1,1},\cdots,2(T_{1,1}+T_{1,2}-1)$. The momentum is defined as $\xi_{j}^{t}=1-\tilde{\eta}$, where $\tilde{\eta}=c_{2}\varepsilon$.
In the following, we show that the strength of the signal is greater than some constant $\varepsilon$. Thus, for second-layer initialization $c_{a}\lesssim\varepsilon$, the effect of the interaction term $\nabla_{\boldsymbol{w}}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}^{2}$ can be ignored, and we drop the subscript that distinguishes $N$ neurons.
Lemma 13.
Consider the neuron that satisfies $\frac{1}{4}c_{2}\leq\kappa^{2T_{1,1}}_{j}\leq c_{2}$. We have
$\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(T_{1,1}+T_{1,2})}\geq 1-\varepsilon,$ (B.125)
with high probability, where $T_{1,2}=\Theta{O}_{d}(d\varepsilon^{-2})$.
Proof.  The expected gradient (of the correlation term) can be computed as
$\mathbb{E}\big{[}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}$ $=\mathbb{E}\bigg{[}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}\sum_{j=p}^{q}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}\sum_{j=0}^{C_{q}}\beta_{j}\mathsf{He}_{j}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\bigg{)}\bigg{]}$ (B.126)
$=\sum_{j=p}^{q}\big{[}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j}\boldsymbol{w}^{2t}\big{]}.$ (B.127)
Applying $P_{\boldsymbol{w}^{2t}}$, we have
$\mathbb{E}\big{[}P_{\boldsymbol{w}^{2t}}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}=(\boldsymbol{\theta}-({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{\theta})\boldsymbol{w}^{2t})\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}.$ (B.128)
The update of the alignment is
$\kappa^{2(t+1)}\geq\kappa^{2t}+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2},$ (B.129)
where
$\boldsymbol{g}=P_{\boldsymbol{w}^{2t}}yc_{a}\sigma^{\prime}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\boldsymbol{x}.$ (B.130)
From (B.128), the expectation of (B.130) is bounded by
$\mathbb{E}[\kappa^{2(t+1)}]$ $\geq\kappa^{2t}+\eta\tilde{\eta}(1-(\kappa^{2t})^{2})\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}-C_{3}\eta^{2}\tilde{\eta}^{2}d(\kappa^{2t}+\eta\tilde{\eta}).$ (B.131)
$\geq\kappa^{2t}+\eta\tilde{\eta}\tilde{\varepsilon}p!\alpha_{p}\beta_{p}(\kappa^{2t})^{p-1}-C_{3}\eta^{2}\tilde{\eta}^{2}d(\kappa^{2t}+\eta\tilde{\eta}).$ (B.132)
By letting $\eta\leq c_{1}d^{-1}$ and $\tilde{\eta}\leq c_{2}\varepsilon$, we have
$\mathbb{E}[\kappa^{2(t+1)}]\geq\kappa^{2t}+\frac{1}{2}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(\kappa^{2t})^{p-1}.$ (B.133)
Because the noise $\nu^{2t}:=\eta^{-1}\tilde{\eta}^{-1}(\kappa^{2(t+1)}-\mathbb{E}[\kappa^{2(t+1)}])$ satisfies $\mathbb{P}[|\nu^{2t}|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$,
$\kappa^{2(T{1,1}+t)}\geq\kappa^{2T_{1,1}}+\frac{1}{2}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}\sum_{s=T_{1,1}}^{T_{1,1}+t-1}(\kappa^{2s})^{p-1}+\eta\tilde{\eta}\sum_{s=T_{1,1}}^{T_{1,1}+t-1}\nu^{2s},$ (B.134)
with high probability. The third term is bounded by $\eta\tilde{\eta}C_{3}t\leq\frac{1}{8}c_{2}$ when $t\leq\frac{1}{8}\eta^{-1}\tilde{\eta}^{-1}C_{3}^{-1}c_{2}$ and by $\eta\tilde{\eta}C_{3}\sqrt{t}\leq 4\eta^{\frac{3}{2}}\tilde{\eta}^{\frac{3}{2}}c_{2}^{-\frac{1}{2}}C_{3}^{\frac{3}{2}}t\leq\frac{1}{2}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(c_{2}/8)^{p-1}$ when $t\leq\frac{1}{8}\eta^{-1}\tilde{\eta}^{-1}C_{3}^{-1}c_{2}$ and $\varepsilon=\tilde{O}(d^{-1})$.
Therefore, if $\kappa^{2s}\geq\frac{1}{8}c_{2}$ holds for all $s=T_{1,1},\cdots,T_{1,1}+t-1$, we have
$\kappa^{2(T_{1,1}+t)}\geq\frac{1}{8}c_{2}+\frac{1}{4}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(c_{2}/8)^{p-1}t,$ (B.135)
with high probability. Thus, by induction, $\kappa^{2(T{1,1}+t)}\geq\frac{1}{8}c_{2}$ holds and (B.135) holds for all $t$, until we get $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$. Because of (B.135), we have $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ until $t\leq T_{1,2}$, where $T_{1,2}\geq\big{(}\frac{1}{4}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(c_{2}/8)^{p-1}\big{)}^{-1}=\tilde{O}\big{(}(\eta\tilde{\eta}\varepsilon)^{-1}\big{)}=\tilde{O}(d\varepsilon^{-2})$.
After we achieve the strong recovery $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ for some $t$, $\kappa^{2(T{1,1}+s)}$ may get smaller than $1-\varepsilon$. However, by letting $s^{\prime}$ be the first such step, because at each step the alignment only moves at most $\tilde{O}(\eta\tilde{\eta})=\tilde{O}(d^{-1}\varepsilon)$, $s^{\prime}$ should still satisfies $\kappa^{2(T{1,1}+s^{\prime})}\geq 1-2\varepsilon\geq c_{2}$. Thus, (B.135) holds again until $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$. Therefore, we can guarantee $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ after $t\geq T_{1,2}$, with high probability. ∎
B.6 Second Layer Training
This subsection proves the generalization error after training the second layer. Let $f_{\boldsymbol{a}}(\boldsymbol{x})=f_{\boldsymbol{\Theta}}(\boldsymbol{x})$ for $\boldsymbol{\Theta}=(\hat{\boldsymbol{w}}_{j},a_{j},\hat{b}_{j})_{j=1}^{N}$ where $\boldsymbol{a}\in\mathbb{R}^{N}$ and $(\hat{\boldsymbol{w}}_{j},\hat{b}_{j})_{j=1}^{N}$ are the parameters trained in the first stage. Here we let $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ be the “certificate” such that $\|\boldsymbol{a}^{*}\|=\tilde{O}(N)$ that is shown to exist in Lemma 15 (here we suppress dependence on constants $p,q$). The following lemma is a complete version of Proposition 4.
Lemma 14.
There exists a $4q$-th order polynomial $Q(R_{w},b,q^{\prime})$ of $R_{\boldsymbol{w}}=\max_{j}\mathopen{}\mathclose{{}\left\|{\boldsymbol{w}_{j}}}\right\|$ and $b=(b_{j})_{j=1}^{N}$ such that, if we set $\lambda=\Theta\mathopen{}\mathclose{{}\left(\sqrt{\frac{2}{T_{2}\delta_{0}}N^{2}Q(R_{w},b,q^{\prime})}}\right)$ for some $\delta_{0}>0$, the ridge estimator $\hat{\boldsymbol{a}}$ satisfies
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.136)
with probability $1-\delta_{0}$. Therefore, by taking $T_{2}=\tilde{\Theta}((N^{4}Q(R_{w},b,q)+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}])\varepsilon^{-2})$, and $N=\tilde{\Theta}(\varepsilon^{-1})$, we have
$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon.$
Proof.  Let $P_{T_{2}}$ be the empirical distribution of the second stage: $P_{T_{2}}:=\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\delta_{\boldsymbol{x}_{i}}$. Let $\psi(\boldsymbol{x})=(\sigma(\langle\boldsymbol{x},\widehat{\boldsymbol{w}}_{j})\rangle+b_{j}))_{j=1}^{N}$ so that $f_{\boldsymbol{a}}(\boldsymbol{x})=\langle\boldsymbol{a},\psi(\boldsymbol{x})\rangle$.
Part (1).
Here, we first bound the second term $\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}$. Since ^$\hat{\mathcal{L}}(f_{\hat{\boldsymbol{a}}})+\lambda\|\hat{\boldsymbol{a}}\|^{2}\leq\hat{\mathcal{L}}(f_{\boldsymbol{a}^{*}})+\lambda\|\boldsymbol{a}^{*}\|^{2}$, we have that
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}+\lambda\|\hat{\boldsymbol{a}}\|^{2}$ (B.137)
$\leq\|f_{\boldsymbol{a}^{*}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}+\frac{2}{T_{2}}\sum_{i=1}^{T_{2}}(f_{\boldsymbol{a}^{*}}(\boldsymbol{x}_{i})-f_{\hat{\boldsymbol{a}}}(\boldsymbol{x}_{i}))\varepsilon_{i}+\lambda\|\boldsymbol{a}^{*}\|^{2}.$ (B.138)
Now, by the Cauchy-Schwarz inequality, we have
$\frac{2}{T_{2}}\sum_{i=1}^{T_{2}}(f_{\boldsymbol{a}^{*}}(x_{i})-f_{\hat{\boldsymbol{a}}}(x_{i}))\varepsilon_{i}$ $=(\boldsymbol{a}^{*}-\hat{\boldsymbol{a}})^{\top}\frac{2}{T_{2}}\sum_{i=1}^{T_{2}}\psi(\boldsymbol{x}_{i})\varepsilon_{i}$
$\leq 2\|\boldsymbol{a}^{*}-\hat{\boldsymbol{a}}\|\sqrt{\frac{\sum_{i,j}\varepsilon_{i}\varepsilon_{j}\psi(\boldsymbol{x}_{i})^{\top}\psi(\boldsymbol{x}_{j})}{T_{2}^{2}}}.$
By applying Markov’s inequality to the right hand side, it can be further bounded by
$\|\boldsymbol{a}^{*}-\hat{\boldsymbol{a}}\|\sqrt{\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{T_{2}\delta_{1}}}\leq\frac{\lambda}{2}\|\hat{\boldsymbol{a}}\|^{2}+\frac{\lambda}{2}\|\boldsymbol{a}^{*}\|^{2}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{T_{2}\delta_{1}\lambda},$
with probability $1-\delta_{1}$. Thus, by combining with (B.137), we arrive at
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}+\frac{\lambda}{2}\|\hat{\boldsymbol{a}}\|^{2}\leq\|f_{\boldsymbol{a}^{*}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{T_{2}\delta_{1}\lambda}+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2}.$
Here, by using the evaluation $\|f_{\boldsymbol{a}^{*}}-f_{*}\|_{L^{2}(P_{T_{2}})}=\tilde{O}(N^{-1}+\varepsilon)$ in Lemma 15, the right hand side can be further bounded by
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}+\frac{\lambda}{2}\|\hat{\boldsymbol{a}}\|^{2}\leq\tilde{O}(N^{-2}+\varepsilon^{2})+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{T_{2}\delta_{1}\lambda}+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2}.$
Part (2).
Next we lower bound $\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}$ by noticing that
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}$ (B.139)
$=\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}+\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ (B.140)
$=\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{x})}^{2}-2\mathopen{}\mathclose{{}\left(\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}f_{\hat{\boldsymbol{a}}}(\boldsymbol{x}_{i})f_{*}(\boldsymbol{x}_{i})-\mathbb{E}[f_{\hat{\boldsymbol{a}}}(\boldsymbol{x}_{i})f_{*}(\boldsymbol{x}_{i})]}\right)$ (B.141)
$~{}~{}~{}~{}~{}+\|f_{*}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{*}\|_{L^{2}(P_{x})}^{2}+\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}.$ (B.142)
The first two terms of Eq. (B.142) can be bounded by
$\mathopen{}\mathclose{{}\left|\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{x})}^{2}}\right|$ $=\mathopen{}\mathclose{{}\left|\hat{\boldsymbol{a}}^{\top}\mathopen{}\mathclose{{}\left(\frac{\sum_{i=1}^{T_{2}}\psi(\boldsymbol{x}_{i})\psi(\boldsymbol{x}_{i})^{\top}}{T_{2}}-\mathbb{E}_{\boldsymbol{x}}[\psi(\boldsymbol{x})\psi(\boldsymbol{x})^{\top}]}\right)\hat{\boldsymbol{a}}}\right|$
$\leq\|\hat{\boldsymbol{a}}\|^{2}\sup_{\boldsymbol{a}:\|\boldsymbol{a}\|\leq 1}\mathopen{}\mathclose{{}\left|\|f_{\boldsymbol{a}}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{\boldsymbol{a}}\|_{L^{2}(P_{x})}^{2}}\right|.$
The standard Rademacher complexity bound yields that
$\mathbb{E}_{(x_{i})_{i=1}^{T_{2}}}\mathopen{}\mathclose{{}\left[\sup_{\boldsymbol{a}\in\mathbb{R}^{N}:\|\boldsymbol{a}\|\leq 1}\mathopen{}\mathclose{{}\left|\|f_{\boldsymbol{a}}\|_{L^{2}(P_{x})}^{2}-\|f_{\boldsymbol{a}}\|_{L^{2}(P_{T_{2}})}^{2}}\right|}\right]$ (B.143)
$\leq$ $2\mathbb{E}_{(x_{i},\sigma_{t})_{t=1}^{T_{2}}}\mathopen{}\mathclose{{}\left[\sup_{\boldsymbol{a}\in\mathbb{R}^{N}:\|\boldsymbol{a}\|\leq 1}\mathopen{}\mathclose{{}\left|\frac{1}{T_{2}}\sum_{t=1}^{T_{2}}\sigma_{t}f_{\boldsymbol{a}}(x_{i})^{2}}\right|}\right]$ (B.144)
$\leq$ $2\sqrt{\mathbb{E}_{(\boldsymbol{x}_{i})_{i=1}^{T_{2}}}\mathopen{}\mathclose{{}\left[\sup_{\boldsymbol{a}\in\mathbb{R}^{N}:\|\boldsymbol{a}\|\leq 1}\frac{1}{T_{2}^{2}}\sum_{i=1}^{T_{2}}(\boldsymbol{a}^{\top}\psi(\boldsymbol{x}_{i}))^{4}}\right]}$ (B.145)
$\leq$ $2\sqrt{\mathbb{E}_{(\boldsymbol{x}_{i})_{i=1}^{T_{2}}}\mathopen{}\mathclose{{}\left[\frac{1}{T_{2}^{2}}\sum_{i=1}^{T_{2}}\|\psi(\boldsymbol{x}_{i})\|^{4}}\right]}$ (B.146)
$=$ $2\sqrt{\frac{1}{T_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]},$ (B.147)
where $(\sigma_{i})_{i=1}^{T_{2}}$ is the i.i.d. Rademacher sequence which is independent of $(\boldsymbol{x}_{i})_{i=1}^{T_{2}}$. Hence, Markov’s inequality yields that
$\mathopen{}\mathclose{{}\left|\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{x})}^{2}}\right|$ $=2\|\hat{\boldsymbol{a}}\|^{2}\sqrt{\frac{1}{T_{2}\delta_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]},$
with probabilty $1-\delta_{2}$.
The third term in Eq. (B.142) can be evaluated as
$2\mathopen{}\mathclose{{}\left(\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}f_{\hat{\boldsymbol{a}}}(\boldsymbol{x}_{i})f_{*}(\boldsymbol{x}_{i})-\mathbb{E}_{\boldsymbol{x}}[f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})f_{*}(\boldsymbol{x})]}\right)$
$=\hat{\boldsymbol{a}}^{\top}\mathopen{}\mathclose{{}\left(\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}(\psi(\boldsymbol{x}_{i})f_{*}(\boldsymbol{x}_{i})-\mathbb{E}_{\boldsymbol{x}}[\psi(\boldsymbol{x})f_{*}(\boldsymbol{x})])}\right)$
$\leq\|\hat{\boldsymbol{a}}\|\sqrt{\frac{1}{T_{2}^{2}}\sum_{i=1}^{T_{2}}\sum_{j=1}^{T_{2}}(\psi(\boldsymbol{x}_{i})f_{*}(\boldsymbol{x}_{i})-\mathbb{E}_{\boldsymbol{x}}[\psi(\boldsymbol{x})f_{*}(\boldsymbol{x})])^{\top}(\psi(\boldsymbol{x}_{j})f_{*}(\boldsymbol{x}_{j})-\mathbb{E}_{\boldsymbol{x}}[\psi(\boldsymbol{x})f_{*}(\boldsymbol{x})])}$
$\leq\|\hat{\boldsymbol{a}}\|\sqrt{\frac{1}{T_{2}\delta_{3}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})f_{*}(\boldsymbol{x})-\mathbb{E}_{\boldsymbol{x}}[\psi(\boldsymbol{x})f_{*}(\boldsymbol{x})]\|^{2}]}$
$\leq\|\hat{\boldsymbol{a}}\|\sqrt{\frac{1}{T_{2}\delta_{3}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}+\|f_{*}(\boldsymbol{x})\|^{4}]}$
$\leq\frac{\lambda}{4}\|\hat{\boldsymbol{a}}\|^{2}+\frac{1}{\lambda T_{2}\delta_{3}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}+\|f_{*}(\boldsymbol{x})\|^{4}],$
with probability $1-\delta_{3}$ where we used Markov’s inequality again in the second inequality.
Finally, the fourth and fifth term in Eq. (B.142) can be bounded as
$\mathopen{}\mathclose{{}\left|\|f_{*}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{*}\|_{L^{2}(P_{x})}^{2}}\right|$ $=\sqrt{\mathopen{}\mathclose{{}\left(\|f_{*}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{*}\|_{L^{2}(P_{x})}^{2}}\right)^{2}}$
$\leq\sqrt{\frac{1}{T_{2}\delta_{4}}\mathbb{E}_{\boldsymbol{x}}[(f^{*}(\boldsymbol{x})^{4}-\|f_{*}\|_{L^{2}(P_{x})}^{2})^{2}]}$
$\leq\sqrt{\frac{1}{T_{2}\delta_{4}}\mathbb{E}_{\boldsymbol{x}}[(f^{*}(\boldsymbol{x}))^{4}]},$
with probability $1-\delta_{4}$ where we used Markov’s inequality in the last inequality.
Combining these inequalities, we finally arrive at
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}+\mathopen{}\mathclose{{}\left(\frac{\lambda}{4}-\sqrt{\frac{2}{T_{2}\delta_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}}\right)\|\hat{\boldsymbol{a}}\|^{2}$
$\leq\tilde{O}(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda}\mathopen{}\mathclose{{}\left(\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{1}}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{3}}+\frac{\mathbb{E}_{\boldsymbol{x}}[(f^{*})^{4}]}{\delta_{3}}}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$
with probability $1-\sum_{j=1}^{4}\delta_{j}$. Hence, by setting $\lambda\geq 8\sqrt{\frac{2}{T_{2}\delta_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}$, we have that
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$
$\leq\tilde{O}(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda}\mathopen{}\mathclose{{}\left(\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{1}}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}{\delta_{3}}+\frac{\mathbb{E}_{\boldsymbol{x}}[(f^{*})^{4}]}{\delta_{3}}}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2}.$
When the activation function $\sigma$ is a polynomial, then each $\psi_{j}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle+b_{j})$ is an order $q$-polynomial of a Gaussian random variable $\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle$ which has mean 0 and variance $\mathbb{E}[\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle^{2}]=\|\boldsymbol{w}_{j}\|^{2}=\tilde{O}(1)$. Then, if we let $R_{w}:=\max_{j}\|\boldsymbol{w}_{j}\|=\tilde{O}(1)$, the term $\max_{j}\max\{\mathbb{E}_{\boldsymbol{x}}[\psi(\boldsymbol{x})_{j}^{2}],\mathbb{E}_{\boldsymbol{x}}[\psi(\boldsymbol{x})_{j}^{4}]\}$ can be bounded by a $4q$-th order polynomial of $R_{w}$ and $b$, which can be denoted by $Q(R_{w},b,4q)$.
Part (3).
By combining evaluations of (1) and (2) together, if we let $\lambda=8\sqrt{\frac{2}{T_{2}\delta_{0}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}$ for some $\delta_{0}>0$, (by ignoring polylogarithmic factors) we obtain that
$!\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.148)
with probability $1-4\delta_{0}$. Thus, since $\|\boldsymbol{a}^{*}\|^{2}=\tilde{O}(N)$, by setting $T_{2}=\tilde{\Theta}((N^{4}Q(R_{w},b,q^{\prime})+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}])\varepsilon^{-2})$, and $N=\tilde{\Theta}(\varepsilon^{-1})$, we obtain that (B.136)$\lesssim\varepsilon$. ∎
Finally, we provide the approximation guarantee: If $\sigma$ is a degree-$q$ polynomial, we have the following result, which follows Lemmas 29 and 30 of [OSSW24].
Lemma 15.
Suppose that there exist at least $N^{\prime}=\tilde{\Theta}(N)$ neurons that satisfy $\|\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta}\|\leq\varepsilon$ and $\sigma$ is a polynomial link function with degree at least $q$. Let $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$ with $C_{b}=\tilde{O}(1)$ , and consider approximation of a ridge function $h(\boldsymbol{\theta}^{\top}\boldsymbol{x})$ with its degree at most $q$. Then, there exists $a_{1},\dots,a_{N}$ such that
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1}+\varepsilon)$ (B.149)
with high probability, where $(\boldsymbol{x},y)$ is a random sample, and we omit dependence on the degree $q$ in the big-$O$ notation. Moreover, we have $\sum_{j=1}^{N}a_{j}^{2}=\tilde{O}(N)$.
We rely on the following result.
Lemma 16.
Suppose that $C_{b}\geq q$. For any polynomial $h(s)$ with its degree at most $q$, there exists $\bar{v}(b;h)$ with $|\bar{v}(b;h)|\lesssim C_{b}$ such that for all $s$,
$\mathbb{E}[\bar{v}(b;h)\sigma(\delta s+b)]=h(s).$ (B.150)
Proof.  When $g_{q}(s)=\sigma(s)$ is a degree-$q$ polynomial,
$g_{q}(s)=\int_{b=-q}^{0}\sigma(s+b)\mathrm{d}b$ (B.151)
is also a degree-$q$ polynomial. Let us repeatedly define
$g_{q-i}(s):=g_{q-(i-1)}(s+1)-g_{q-(i-1)}(s)\quad(i=1,2,\cdots,q),$ (B.152)
and let $(c_{i,j})$ be coefficients so that $(s-1)^{i}=\sum_{j=0}^{i}c_{i,j}s^{j}$ holds for all $z$. Then, by induction, $g_{i}(s)$ is a degree-$i$ polynomial. Moreover, we have
$g_{q-i}(s)$ $=\sum_{j=0}^{i}c_{i,j}\int_{b=-q}^{0}\sigma(s+b+j)\mathrm{d}b$ (B.153)
$=2C_{b}\mathbb{E}_{b\sim\mathrm{Unif}([-C_{b},C_{b}])}\bigg{[}\bigg{(}\sum_{j=0}^{i}c_{i,j}\mathbbm{1}[j-q\leq b\leq j]\bigg{)}\sigma(s+b)\bigg{]},$ (B.154)
when $C_{b}\geq q$. Therefore, for any polynomial $h(s)$ with its degree at most $q$, there exists $\bar{v}(b;h)$ with $|\bar{v}(b;h)|\lesssim C_{b}$ such that for all $s$,
$\mathbb{E}[\bar{v}(b;h)\sigma(\delta s+b)]=h(s).$ (B.155)
Proof of Lemma 15. We now discretize Lemma 16. We focus on $N^{\prime}$ neurons that satisfy $\|\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta}\|\leq\varepsilon$ (by letting $a_{j}=0$ otherwise).
For $A=\tilde{\Theta}(N^{\prime})=\tilde{\Theta}(N)$ (with a small hidden constant), we consider $2A$ intervals $[-C_{b},C_{b}(-1+\frac{1}{A})),[C_{b}(-1+\frac{1}{A}),C_{b}(-1+\frac{2}{A})),\cdots,[C_{b}(1-\frac{1}{A}),C_{b}]$. By taking the hidden constant sufficiently small, for each interval there exists at least one $b_{j}$. Then, for $b_{j}$ corresponding to $[C_{b}(-1+\frac{i}{A}),C_{b}(-1+\frac{i+1}{A}))$, we set $a_{j}=\frac{N}{2}\int_{C_{b}(-1+\frac{i}{A})}^{C_{b}(-1+\frac{i+1}{A}))}\bar{v}(b;h)\mathrm{d}b$. Here we note that $|a_{j}|=\tilde{O}(1)$ holds for all $j$. If each interval contains more than one $b_{j}$, we ignore all but one by letting $a_{j}=0$ except for the one. By doing so, because of Lipschitzness of $\sigma$, we have
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(s+b_{j})-h(s)}\right|=\tilde{O}(N)$ (B.156)
for all $s=\tilde{O}(1)$. Because $|\boldsymbol{\theta}^{\top}\boldsymbol{x}^{t}|=\tilde{O}(1)$ with high probability, we have
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1})$ (B.157)
with high probability. Finally, because $\|\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta}\|\leq\varepsilon$, we have
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})}\right|=\tilde{O}\big{(}(\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta})^{\top}\boldsymbol{x}\big{)}=\tilde{O}(\varepsilon).$ (B.158)
Combining (B.157) and (B.158), we obtain the assertion. ∎

### v2

@@
 Abstract
-We study the problem of gradient descent learning of a single-index target function $f_{*}(\boldsymbol{x})=\textstyle\sigma_{*}\mathopen{}\mathclose{{}\left(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle}\right)$ under isotropic Gaussian data in $\mathbb{R}^{d}$, where the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ is an unknown degree $q$ polynomial with information exponent $p$ (defined as the lowest degree in the Hermite expansion).
-Prior works showed that gradient-based training of neural networks can learn this target with $n\gtrsim d^{\Theta(p)}$ samples, and such statistical complexity is predicted to be necessary by the correlational statistical query lower bound.
-Surprisingly, we prove that a two-layer neural network optimized by an SGD-based algorithm learns $f_{*}$ of arbitrary polynomial link function with a sample and runtime complexity of $n\asymp T\asymp C(q)\cdot d\mathrm{polylog}d$, where constant $C(q)$ only depends on the degree of $\sigma_{*}$, regardless of information exponent; this dimension dependence matches the information theoretic limit up to polylogarithmic factors.
+We study the problem of gradient descent learning of a single-index target function $f_{*}(\boldsymbol{x})=\textstyle\sigma_{*}\mathopen{}\mathclose{{}\left(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle}\right)$ under isotropic Gaussian data in $\mathbb{R}^{d}$, where the unknown link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ has information exponent $p$ (defined as the lowest degree in the Hermite expansion).
+Prior works showed that gradient-based training of neural networks can learn this target with $n\gtrsim d^{\Theta(p)}$ samples, and such complexity is predicted to be necessary by the correlational statistical query lower bound.
+Surprisingly, we prove that a two-layer neural network optimized by an SGD-based algorithm (on the squared loss) learns $f_{*}$ with a complexity that is not governed by the information exponent.
+Specifically, for arbitrary polynomial single-index models, we establish a sample and runtime complexity of $n\simeq T=\Theta(d\!\cdot\!\mathrm{polylog}d)$, where $\Theta(\cdot)$ hides a constant only depending on the degree of $\sigma_{*}$; this dimension dependence matches the information theoretic limit up to polylogarithmic factors.
+More generally, we show that $n\gtrsim d^{(p_{*}-1)\vee 1}$ samples are sufficient to achieve low generalization error, where $p_{*}\leq p$ is the generative exponent of the link function.
 Core to our analysis is the reuse of minibatch in the gradient computation, which gives rise to higher-order information beyond correlational queries.
@@
 Single-index models are a classical class of functions that capture low-dimensional structure in the learning problem.
-To efficiently estimate such functions, the learning algorithm should extract the relevant (one-dimensional) subspace from high-dimensional observations; hence this problem setting has been extensively studied in deep learning theory [BL20, BES+22, BBSS22, MHPG+23, WWF24], to examine the adaptivity to low-dimensional targets and benefit of representation learning in neural networks (NNs) optimized by gradient descent (GD).
-In this work we study the learning of a single-index target function with polynomial link function under isotropic Gaussian data:
+To efficiently estimate such functions, the learning algorithm should extract the relevant (one-dimensional) subspace from high-dimensional observations; hence this problem setting has been extensively studied in deep learning theory [BL20, BES+22, BBSS22, MHPG+23, MZD+23, WWF24], to examine the adaptivity to low-dimensional targets and benefit of representation learning in neural networks (NNs) optimized by gradient descent (GD).
+In this work we study the learning of a single-index target function under isotropic Gaussian data:
 $y_{i}=f_{*}(\boldsymbol{x}_{i})+\varsigma_{i},\quad f_{*}(\boldsymbol{x}_{i})=\sigma_{*}(\langle\boldsymbol{x}_{i},\boldsymbol{\theta}\rangle),\quad\boldsymbol{x}_{i}\stackrel{{\scriptstyle\mathrm{\tiny{i.i.d.}}}}{{\sim}}\mathcal{N}(0,\boldsymbol{I}_{d}),$ (1.1)
 where $\varsigma_{i}$ is i.i.d.
-label noise, $\boldsymbol{\theta}\in\mathbb{R}^{d}$ is the direction of index features, and we assume the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ is a degree-$q$ polynomial with information exponent $p$ defined as the index of the first non-zero coefficient in the Hermite expansion (see Definition 1).
-Equation (1.1) requires the estimation of the one-dimensional link function $\sigma_{*}$ and the relevant direction $\boldsymbol{\theta}$; it is known that learning is information theoretically possible with $n\gtrsim d$ training examples [Bac17, DPVLB24].
-Indeed, when $\sigma_{*}$ is polynomial, such statistical complexity can be achieved up to logarithmic factors by a tailored algorithm that exploit the specific structure of the low-dimensional target [CM20].
+label noise, $\boldsymbol{\theta}\in\mathbb{R}^{d}$ is the direction of index features, and we assume the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ has information exponent $p\in\mathbb{N}_{+}$ defined as the index of the first non-zero coefficient in the Hermite expansion (see Definition 1).
+Equation (1.1) requires the estimation of the one-dimensional link function $\sigma_{*}$ and the relevant direction $\boldsymbol{\theta}$; it is known that learning is information theoretically possible with $n\gtrsim d$ training examples [DH24, DPVLB24].
+Indeed, when $\sigma_{*}$ is polynomial, such statistical complexity can be achieved up to logarithmic factors by a tailored algorithm that exploit the structure of low-dimensional target [CM20].
 On the other hand, for gradient-based training of two-layer NNs, existing works established a sample complexity of $n\gtrsim d^{\Theta(p)}$ [BAGJ21, BBSS22, DNGL23], which presents a gap between the information theoretic limit and what is computationally achievable by (S)GD.
-Such a gap is also predicted by the correlational statistical query (SQ) lower bound [DLS22, AAM23], which roughly states that for a CSQ algorithm to learn (isotropic) Gaussian single-index models using less than exponential compute, a sample size of $n\gtrsim d^{p/2}$ is necessary.
-Although CSQ lower bounds are frequently cited to imply a fundamental barrier of learning via SGD (with the squared loss), strictly speaking, the CSQ model does not include empirical risk minimization with gradient descent, due to the non-adversarial noise and existence of non-correlational terms in the gradient computation.
-Very recently, [DTA+24] exploited higher-order terms in the gradient update arising from the reuse of the same training data, and showed that for certain link functions with high information exponent ($p>2$), two-layer NNs may still achieve weak recovery (i.e., nontrivial overlap with $\boldsymbol{\theta}$) after two GD steps with $O(d)$ batch size.
+Such a gap is also predicted by the correlational statistical query (CSQ) lower bound [DLS22, AAM23], which roughly states that for a CSQ algorithm to learn (isotropic) Gaussian single-index models using less than exponential compute, a sample size of $n\gtrsim d^{p/2}$ is necessary.
+Although CSQ lower bounds are frequently cited to imply a fundamental barrier of learning via SGD (with the squared/correlation loss), strictly speaking, the CSQ model does not include empirical risk minimization with gradient descent, due to the non-adversarial noise and existence of non-correlational terms in the gradient computation.
+Very recently, [DTA+24] exploited higher-order terms in the gradient update arising from the reuse of the same training data, and showed that for certain link functions with high information exponent ($p>2$), two-layer NNs may still achieve weak recovery (i.e., nontrivial overlap with $\boldsymbol{\theta}$) after two GD steps with $\Theta(d)$ batch size.
 While this presents evidence that GD-trained NNs can learn $f_{*}$ beyond the sample complexity suggested by the CSQ lower bound, the weak recovery statement in [DTA+24] may not translate to statistical guarantees; moreover, the class of functions where SGD can achieve vanishing generalization error is not fully characterized, as only a few specific examples of link functions are discussed.
@@
 (b) Same-batch GD (generalization error).
-Figure 1: We train a two-layer ReLU NN (3.1) with $N=1024$ neurons using SGD with step size $\eta=1/d$ to learn a single-index target $f_{*}(\boldsymbol{x})=\mathsf{He}_{3}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)$; heatmaps are values averaged over 10 runs.
+Figure 1: We train a ReLU NN (3.1) with $N=1024$ neurons using SGD (squared loss) with step size $\eta=1/d$ to learn a single-index target $f_{*}(\boldsymbol{x})=\mathsf{He}_{3}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)$; heatmaps are values averaged over 10 runs.
 $(a)$ online SGD with batch size $B=8$; $(b)$ GD on the same batch of size $n$ for $T=2^{14}$ steps.
-We only report weak recovery (i.e., overlap between parameters $\boldsymbol{w}$ and target $\boldsymbol{\theta}$, averaged across neurons) for online SGD since the generalization error does not drop.
-Given the existence of (non-NN) algorithms that learn any single-index polynomials with $n=\tilde{O}(d)$ samples [CM20, DPVLB24], it is natural to ask if gradient-based training of NNs can achieve similar statistical efficiency for the same function class.
-Motivated by observations in [DTA+24] that SGD with reused batch may break the “curse of information exponent”, we aim to address the following question:
-Can a two-layer NN optimized by SGD with reused batch learn arbitrary polynomial single-index models near the information-theoretic limit $n\asymp d$, regardless of the information exponent?
+For online SGD we only report weak recovery (i.e., averaged overlap between neuron $\boldsymbol{w}$ and target $\boldsymbol{\theta}$) since the test error does not drop.
+Given the existence of (non-NN) algorithms that learn any single-index polynomials in $n=\tilde{O}(d)$ samples [CM20] regardless of the information exponent $p$, and more generally, non-CSQ algorithms with a sample complexity surpassing the CSQ lower bound [DPVLB24], it is natural to ask if gradient-based training of NNs can achieve similar statistical efficiency for this function class.
+Motivated by observations in [DTA+24] that SGD with reused data may break the “curse of information exponent”, we aim to address the question:
+Can NN optimized by SGD with reused batch learn single-index $f_{*}$ beyond the CSQ lower bound? And for polynomial $\sigma_{*}$, can learning succeed near the information-theoretic limit $n\simeq d$?
 Empirically, the separation between one-pass (online) and multi-pass SGD is clearly observed in Figure 1, where we trained the same two-layer ReLU neural network to learn a single-index polynomial with information exponent $p=3$.
-We see that SGD with reused batch (Figure 1(b)) reaches low generalization error using roughly $n\asymp d$ samples, whereas online SGD fails to achieve even weak recovery with much larger sample size $n=\Omega(d^{2})$.
-Our main contribution is to establish this near-optimal sample complexity for two-layer NNs trained by a variant of SGD with reused training data.
+We see that SGD with reused data (Figure 1(b)) reaches low test error using roughly $n\simeq d$ samples, whereas online SGD fails to achieve even weak recovery with much larger sample size $n=\Omega(d^{2})$.
+Our main contribution is to establish this improved statistical complexity for two-layer NNs trained by a variant of SGD with reused training data.
 1.1 Our Contributions
-We answer the above question in the affirmative by showing that for (1.1) with arbitrary polynomial link function, SGD training on a natural class of shallow NNs can achieve small generalization error using polynomial compute and $n=\tilde{O}(d)$ training examples, if we employ a layer-wise optimization procedure (analogous to that in [BES+22, DLS22, AAM23]) and reuse of the same minibatch:
+We answer the above question in the affirmative by showing that SGD training (with the squared loss) on a natural class of shallow NNs can achieve small generalization error using polynomial compute and a sample complexity that is not governed by the information exponent, if we employ a layer-wise optimization procedure (analogous to that in [BES+22, DLS22, AAM23]) and reuse of the same minibatch.
+The core insight is that SGD can implement a full statistical query (SQ) algorithm that goes beyond CSQ, despite the correlational structure of the squared loss.
+Our main finding is summarized by the following theorem.
 Theorem (informal).
-A shallow NN with $N=\tilde{O}_{d}(1)$ neurons can learn arbitrary single-index polynomials up to small population loss: $\mathbb{E}_{\boldsymbol{x}}[(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]=o_{d,\mathbb{P}}(1)$, using $n=\tilde{O}_{d}(d)$ samples, and an SGD-based algorithm (with reused training data) minimizing the squared loss objective in $T=\tilde{O}_{d}(d)$ gradient steps.
+A shallow NN with $N=\tilde{\Theta}_{d}(1)$ neurons can learn arbitrary single-index models up to small population loss: $\mathbb{E}_{\boldsymbol{x}}[(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]=o_{d,\mathbb{P}}(1)$, if we employ an SGD-based algorithm (with reused training data) to minimize the squared loss objective, with a sample and runtime complexity of $n,T=\tilde{\Theta}_{d}(d^{(p_{*}-1)\vee 1})$, where $p_{*}$ is the generative exponent of the link function $\sigma_{*}$.
+Note that the generative exponent [DPVLB24] is defined as the minimum information exponent of the link function $\sigma_{*}$ after arbitrary $L^{2}$ transformation, and hence by definition $p_{*}\leq p$ (equality is achieved by the identity transformation).
 We make the following remarks on our main result.
-  • The theorem suggests that NN + SGD with reused batch can match the statistical efficiency of SQ algorithms tailored for low-dimensional polynomial regression [CM20].
-Furthermore, the sample complexity is information theoretically optimal up to polylogarithmic factors, and surpasses the CSQ lower bound for $p>2$ (see Figure 2); this disproves a conjecture in [AAM23] stating that $n\asymp d^{p/2}$ is the optimal sample complexity for empirical risk minimization with SGD.
+  • We know that $p_{*}\leq 2$ for arbitrary polynomial link functions.
+Therefore, the theorem suggests that NN + SGD with reused batch can learn single-index polynomials with a sample complexity $n=\tilde{O}_{d}(d)$ which is information theoretically optimal up to polylogarithmic factors, hence matching the efficiency of SQ algorithms tailored for low-dimensional polynomial regression [CM20].
+  • For non-polynomial $\sigma_{*}$ with high generative exponent $p_{*}>2$, our sample complexity $n\gtrsim d^{p_{*}-1}$ can be interpreted as an SQ version of the online SGD result in [BAGJ21].
+Since the information exponent $p$ can be arbitrarily larger than the generative exponent $p_{*}$, our main theorem disproves a conjecture in [AAM23] stating that $n\asymp d^{p/2}$ is the optimal sample complexity for empirical risk minimization with SGD on the squared loss / correlation loss.
   • A key observation in our analysis is that with suitable activation function, SGD with reused batch can go beyond correlational queries and implement (a subclass of) SQ algorithms.
-This enables polynomial transformations to the labels that reduce the information exponent to (at most) 2, and hence optimization can escape the high-entropy “equator” at initialization in polylogarithmic time.
-Upon completion of this work, we became aware of the preprint [ADK+24] showing weak recovery with similar sample complexity, also by exploiting the reuse of training data.
-Our work was conducted independently and simultaneously, and their contribution is not reflected in the present manuscript (beyond this paragraph).
+This enables polynomial transformations to the labels that reduce the information exponent, and therefore optimization can escape the high-entropy “equator” at initialization in polylogarithmic time.
+Upon completion of this work, we became aware of the preprint [ADK+24] showing weak recovery (for polynomial targets with $p_{*}\leq 2$) with similar sample complexity, also by exploiting the reuse of training data.
+Our work was conducted independently and simultaneously.
 Information theoretic limitSGD + batch reuse [This work] SQ algorithm [CM20]Smoothed SGD [DNGL23] CSQ lower bound [DLS22]One-pass SGD [BAGJ21]Kernel methods [GMMM21]$d$$\tilde{\Theta}(d)$$\tilde{\Theta}(d^{p/2})$$\tilde{\Theta}(d^{p-1})$$\Theta(d^{q})$
-Figure 2: Sample complexity of learning single-index model where the link function $\sigma_{*}$ has degree $q$ and information exponent $p$.
+Figure 2: Complexity of learning single-index model where the link function $\sigma_{*}$ is a degree-$q$ polynomial with information exponent $p$.
 For the CSQ lower bound, we translate the tolerance to sample complexity using the i.i.d.
@@
 2.1 Complexity of Learning Single-index Models
-We aim to learn a single-index model (1.1) where the link function $\sigma_{*}$ is a degree-$q$ polynomial with information exponent $p$ defined as follows.
-Definition 1 (Information exponent [BAGJ21]).
+We aim to learn a single-index model (1.1) where the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ has information exponent $p$ defined as follows [DH18, BAGJ21].
+Definition 1 (Information exponent).
 Let $\{\mathsf{He}_{j}\}_{j=0}^{\infty}$ denote the normalized Hermite polynomials.
-The information exponent of $g\in L^{2}(\gamma)$, denoted by $\mathrm{IE}(g):=p\in\mathbb{N}_{+}$, is the index of the first non-zero Hermite coefficient of $g$, that is, given the Hermite expansion $g(z)=\sum_{i=0}^{\infty}\alpha_{i}\mathsf{He}_{i}(z)$, $p:=\min\{i>0:\alpha_{i}\neq 0\}$.
-By definition we always have $p\leq q$.
-Note that $f_{*}$ contains $\Theta(d)$ parameters to be estimated, and hence information theoretically $n\gtrsim d$ samples are both sufficient and necessary for learning [MM18, BKM+19, DPVLB24]; however the sample complexity achieved by different (polynomial time) learning algorithms depends on structure of the link function.
+The information exponent of $g\in L^{2}(\gamma)$, denoted by $\mathrm{IE}(g):=p\in\mathbb{N}_{+}$, is the index of the first non-zero Hermite coefficient of $g$, that is, given $g(z)=\sum_{i=0}^{\infty}\alpha_{i}\mathsf{He}_{i}(z)$, $p:=\min\{i\!>\!0:\alpha_{i}\!\neq\!0\}$.
+By definition, when $\sigma_{*}$ is a degree-$q$ polynomial, we always have $p\leq q$.
+Note that $f_{*}$ contains $\Theta(d)$ parameters to be estimated, and hence information theoretically $n\gtrsim d$ samples are both sufficient and necessary for learning [MM18, BKM+19, DPVLB24]; however, the sample complexity achieved by different (polynomial time) algorithms depends on structure of the link function.
   • Kernel Methods.
@@
   • Gradient-based Training of NNs.
-While NNs can easily approximate a single-index model [Bac17], the statistical complexity of gradient-based learning established in prior works scales as $n\gtrsim d^{\Theta(p)}$: in the well-specified setting, [BAGJ21] proved a sample complexity of $n=\tilde{\Theta}(d^{p-1})$ for online SGD, which is later improved to $\tilde{\Theta}(d^{p/2})$ by a smoothed objective [DNGL23]; as for the misspecified setting, [BBSS22, DKL+23] showed that $n\gtrsim d^{p}$ samples suffice, and in some cases a $\tilde{\Theta}(d^{p-1})$ complexity is achievable [AAM23, OSSW24].
+While NNs can easily approximate a single-index model [Bac17], the sample complexity of gradient-based learning established in prior works typically scales as $n\gtrsim d^{\Theta(p)}$: in the well-specified setting, [BAGJ21] proved a sample complexity of $n=\tilde{\Theta}(d^{p-1})$ for online SGD, which is later improved to $\tilde{\Theta}(d^{p/2})$ by a smoothed objective [DNGL23]; as for the misspecified setting, [BBSS22, DKL+23] showed that $n\gtrsim d^{p}$ samples suffice, and in some cases a $\tilde{\Theta}(d^{p-1})$ complexity is achievable [AAM23, OSSW24a].
 Consequently, at the information-theoretic limit ($n\asymp d$), existing results can only cover the learning of low information exponent targets [AAM22, BMZ23, BES+23].
-It is worth noting that this exponential dependence on $p$ also appears in the CSQ lower bounds [DLS22, AAM22], which is often considered to be indicative of the performance of SGD learning with the squared loss (see Section 2.2).
+This exponential dependence on $p$ also appears in the CSQ lower bounds [DLS22, AAM22], which is often considered to be indicative of the performance of SGD learning with the squared loss (see Section 2.2).
 Statistical Query Learners.
-If we do not restrict ourselves to correlational queries, then (1.1) can be efficiently solved near the information-theoretic limit.
-Specifically, [CM20] proposed an SQ algorithm that learns any single-index polynomials using $n=\tilde{O}(d)$ samples; the key ingredient is to construct nonlinear transformations to the labels that lowers the information exponent to $2$ (similar preprocessing also appeared in context of phase retrieval [MM18, BKM+19]).
-This is consistent with the recently established SQ lower bound [DPVLB24], which suggests a statistical complexity independent of the information exponent $p$ when $\sigma_{*}$ is polynomial.
+If we do not restrict ourselves to correlational queries, the sample complexity of learning (1.1) can be drastically improved.
+Specifically, for polynomial $\sigma_{*}$, [CM20] gave an SQ algorithm that achieves low generalization error in $n=\tilde{O}(d)$ samples, which is near the information-theoretic limit; the key ingredient is to construct nonlinear transformations to the labels that lowers the information exponent to $2$; similar preprocessing also appeared in context of phase retrieval [MM18, BKM+19].
+Such transformations do not belong to CSQ, but can be utilized by a full SQ learner to enhance the statistical efficiency.
+Recently, [DPVLB24] introduced the generative exponent which governs the complexity of SQ algorithms.
+Definition 2 (Generative exponent).
+The generative exponent (GE) of $g\in L^{2}(\gamma)$ is defined as the lowest information exponent (IE) after arbitrary $L^{2}$ transformation, that is,
+$p_{*}=:\mathrm{GE}(g)=\inf_{\mathcal{T}\in L^{2}(P_{y})}\mathrm{IE}(\mathcal{T}\circ g).$ (2.1)
+The generative exponent is the smallest information exponent obtained by all possible label transformations.
+By definition we always have $p^{*}\leq p$, and the gap between the two indices can be arbitrarily large; for example, for the Hermite polynomials we have $\mathrm{IE}(\mathsf{He}_{k})=k$ whereas $\mathrm{GE}(\mathsf{He}_{k})\leq 2$.
+[DPVLB24] established a sample complexity lower bound of $n=\Omega(d^{\nicefrac{{p_{*}}}{{2}}\vee 1})$ for full SQ learners with polynomial compute (assuming $\tau\approx n^{-1/2}$), and obtained matching upper bound by a tensor partial-trace algorithm.
+Our goal is to show that SGD training of two-layer neural network can also achieve a sample and runtime complexity that scales with $n\simeq d^{\Theta(p_{*})}$, where the dimension dependence is governed by the generative exponent $p_{*}$ instead of the information exponent $p$.
 2.2 Can Gradient Descent Go Beyond Correlational Queries?
@@
 SGD with reused data.
-As discussed in Section 2.1, the gap between SQ and CSQ algorithms primarily stems from the existence of label transformations that decrease the information exponent.
+As previously discussed, the gap between SQ and CSQ algorithms primarily stems from the existence of label transformations that decrease the information exponent.
 While such transformation cannot be utilized by a CSQ learner, [DTA+24] argued that they may arise from two consecutive gradient updates using the same minibatch.
-For illustrative purposes, consider an example where one neuron $f_{\boldsymbol{w}}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}\rangle)$ is updated by two GD steps using the same training example $(\boldsymbol{x},y)$, starting from zero initialization $\boldsymbol{w}^{0}=\mathbf{0}$ (we focus on the correlational term in the loss for simplicity):
+For illustrative purposes, consider one neuron $f_{\boldsymbol{w}}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}\rangle)$ updated by two GD steps using the same data point $(\boldsymbol{x},y)$, starting from zero initialization $\boldsymbol{w}^{0}=\mathbf{0}$ (we focus on the correlational term in the loss for simplicity):
 $\!\!\!\boldsymbol{w}^{2}=$ $~{}\boldsymbol{w}^{1}+\eta\cdot y\sigma^{\prime}(\langle\boldsymbol{x},\boldsymbol{w}^{1}\rangle)\boldsymbol{x}=\eta\sigma^{\prime}(0)\underbrace{y\cdot\boldsymbol{x}}_{\text{\!\!\!CSQ term\!\!\!}}+~{}\eta\underbrace{y\sigma^{\prime}(\eta\sigma^{\prime}(0)\mathopen{}\mathclose{{}\left\|{\boldsymbol{x}}}\right\|^{2}\cdot y)\boldsymbol{x}}_{\text{non-CSQ term}}.$ (2.3)
@@
 2.3 Challenges in Establishing Statistical Guarantees
-Importantly, the analysis in [DTA+24] does not lead to concrete learnability guarantees for the class of single-index polynomials for the following reasons: $(i)$ it is not clear if an appropriate nonlinear transformation that lowers the information exponent can always be extracted from SGD with reused batch, and $(ii)$ the weak recovery guarantee may not translate to a sample complexity for the trained NN to achieve small generalization error.
+Importantly, the analysis in [DTA+24] does not lead to concrete learnability guarantees for the class of single-index polynomials for the following reasons: $(i)$ it is not clear if an appropriate nonlinear transformation that lowers the information exponent can always be extracted from SGD with reused data, and $(ii)$ the weak recovery guarantee may not translate to a sample complexity for the trained NN to achieve small generalization error.
 We elaborate these technical challenges below.
 SGD decreases information exponent.
-To show weak recovery, [DTA+24, Definition 3.1] assumed that the student activation $\sigma$ can reduce the information exponent of the labels to $1$; while a few examples are given (e.g., odd-degree Hermite polynomials), the existence of such transformations in SGD is not guaranteed:
+To show weak recovery, [DTA+24, Definition 3.1] assumed that the student activation $\sigma$ can reduce the information exponent of the labels to $1$; while a few examples are given, the existence of such transformations in SGD is not guaranteed:
   • The label transformation employed in prior SQ algorithms [CM20] is based on thresholding, which reduces the information exponent to $2$ for any polynomial $\sigma_{*}$; however, isolating such function from SGD updates on the squared loss is challenging.
-Instead, we show in Proposition 5 that monomial transformation suffices, and it can be extracted from SGD via Taylor expansion.
-  • If the link function $\sigma_{*}$ is even, then its information exponent after arbitrary nonlinear transformation is at least $2$; such functions are predicted not be not learnable by SGD in the $n\asymp d$ regime [DTA+24].
-To handle this setting, we analyze the SGD update up to $\mathrm{polylog}(d)$ time, at which a nontrivial overlap can be established by a Grönwall-type argument similar to [BAGJ21].
-This recovers results on phase retrieval when $\sigma_{*}(z)=z^{2}$ which requires $n=\Omega(d\log d)$.
+Instead, we make use of monomial transformation which can be extracted from SGD via Taylor expansion.
+  • If the link function satisfies $p_{*}\geq 2$, its information exponent after arbitrary nonlinear transformation is at least $2$; such functions are predicted not be not learnable by SGD in the $n\asymp d$ regime [DTA+24].
+To handle this setting, we analyze the SGD update up to $\mathrm{poly}(d)$ time, at which a nontrivial overlap can be established by a Grönwall-type argument similar to [BAGJ21].
+For $p_{*}=2$, this recovers results on phase retrieval when $\sigma_{*}(z)=z^{2}$ which requires $n=\Omega(d\log d)$ samples.
 From weak recovery to sample complexity.
@@
 Therefore, we need to show that starting from a nontrivial overlap, subsequent gradient steps can achieve strong recovery of the index features (i.e., $|\langle\boldsymbol{w},\boldsymbol{\theta}\rangle|>1-\varepsilon$), despite the link misspecification.
-After the first-layer parameters align with the target function, we can train the second-layer parameters with SGD to learn the link function $\sigma_{*}$ with the aid of random bias units.
-3 SGD Achieves Almost-linear Sample Complexity
-We consider the learning of single-index polynomials with degree $q$ and information exponent $p$:
+After the first-layer parameters align with the target function, we train the second-layer parameters with SGD to learn the link function $\sigma_{*}$ with the aid of random bias units [DLS22].
+3 Learning Polynomial $f_{*}$ in Linear Sample Complexity
+We first consider the setting where $\sigma_{*}$ is polynomial with degree $q$ specified as follows.
 Assumption 1.
 The target function is given as $f_{*}(\boldsymbol{x})=\sigma_{*}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)$, where the link function $\sigma_{*}:\mathbb{R}\to\mathbb{R}$ admits the Hermite decomposition $\sigma_{*}(z)=\sum_{i=p}^{q}\alpha_{i}\mathsf{He}_{i}(z)$.
+For single-index polynomials, we do not expect a computational-to-statistical gap under the SQ class [CM20] — indeed, we will establish learning guarantees near the information theoretic limit $n\asymp d$.
 3.1 Training Algorithm
-We train the following two-layer neural network with $N$ neurons using SGD to minimize the squared loss:
+We train the following two-layer network with $N$ neurons using SGD to minimize the squared loss:
 $f_{\boldsymbol{\Theta}}(\boldsymbol{x})=\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}(\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle+b_{j}),$ (3.1)
@@
   • Algorithm 1 employs a layer-wise training strategy common in the recent feature learning theory literature [DLS22, BES+22, BBSS22, AAM23, MHWSE23], where in the first stage, we optimize the first-layer parameters $\{\boldsymbol{w}_{j}\}_{j=1}^{N}$ with normalized SGD to learn the low-dimensional latent representation (index features $\boldsymbol{\theta}$), and in the second phase, we train the second-layer $\{a_{j}\}_{j=1}^{N}$ to fit the unknown link function $\sigma_{*}$.
-  • The most important part in Phase I of Algorithm 1 is the reuse of the same minibatch in the gradient computation.
+  • The most crucial part in Phase I of Algorithm 1 is the reuse of the same minibatch in the gradient computation.
 Specifically, we sample a fresh batch of training examples in every two GD steps; this enables us to extract non-CSQ terms from two consecutive gradient updates outlined in (2.3).
-  • We introduce an interpolation step between the current and previous iterates with hyperparameter $\xi$ to stabilize the training dynamics; this resembles a negative momentum term often seen in optimization algorithms [AZ18, ZLBH19]; the role of this interpolation step is discussed in Section 4.2.
-We also use a projected gradient update ~$\tilde{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})=(\boldsymbol{I}_{d}-\boldsymbol{w}^{2t}{\boldsymbol{w}^{2t}}^{\top})\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})$ for steps $2t$ and $12t+$ 1, where $\nabla_{\boldsymbol{w}}$ is the Euclidean gradient.
-Similar use of projection also appeared in [DNGL23, AAM23].
+  • We introduce an interpolation step between the current and previous iterates with hyperparameter $\xi$ to stabilize the training dynamics; this resembles a negative momentum often seen in optimization algorithms [AZ18, ZLBH19]; the role of this interpolation is discussed in Section 4.2.
+We use a projected gradient update ~$\tilde{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})=(\boldsymbol{I}_{d}-\boldsymbol{w}^{2t}{\boldsymbol{w}^{2t}}^{\top})\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})$ for steps $2t$ and $12t+$ 1, where $\nabla_{\boldsymbol{w}}$ is the Euclidean gradient; similar use of projection also appeared in [DNGL23, AAM23].
 1
-Input : Learning rates $\eta^{t}$, momentum parameters $\{\xi_{j}^{t}\}_{j=1}^{N}$, number of steps $T_{1},T_{2}$,
-$\ell_{2}$ regularization $\lambda$.
+Input : Step sizes $\eta^{t}$; momentum parameters $\xi^{t}$; training time $T_{1},T_{2}$; $\ell_{2}$ regularization $\lambda$.
 2
-3Initialize $\boldsymbol{w}^{0}_{j}\sim\mathbb{S}^{d-1}(1)$, $a_{j}\sim\mathrm{Unif}\{\pm r_{a}\}$.
+3Initialize $\boldsymbol{w}^{0}_{j}\sim\mathbb{S}^{d-1}(1)$, $a_{j}\sim\mathrm{Unif}\{\pm c_{a}\}$.
 4Phase I: normalized SGD on first-layer parameters
 5       for $t=0$ to $T_{1}$ do
-6             if $t>0$ is even then
-7                   Draw i.i.d. sample $(\boldsymbol{x},y)$.
-8                   Interpolate ${\boldsymbol{w}}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\xi_{j}^{t}(\boldsymbol{w}^{t}_{j}-\boldsymbol{w}^{t-2}_{j})$.
-9                   Normalize $\boldsymbol{w}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}/\|\boldsymbol{w}^{t}_{j}\|$.
-10            $\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2},\quad(j=1,\dots,N)$.
-11
-12
-13Initialize $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$.
-14Phase II: SGD on second-layer parameters
-15       $\hat{\boldsymbol{a}}\leftarrow\mathrm{argmin}_{\boldsymbol{a}\in\mathbb{R}^{N}}\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\mathopen{}\mathclose{{}\left(f_{\boldsymbol{\Theta}}(\boldsymbol{x}_{i})-y_{i}}\right)^{2}+\lambda\|\boldsymbol{a}\|^{2}.$
-16
+6             if $t$ is even then
+                   $\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d}),~{}y=f_{*}(\boldsymbol{x})+\varsigma$ ;
+                    // Draw i.i.d. data $(\boldsymbol{x},y)$
+                   ${\boldsymbol{w}}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\xi_{j}^{t}(\boldsymbol{w}^{t}_{j}-\boldsymbol{w}^{t-2}_{j})$,   (when $t>0$) ;
+                    // Interpolation step
+                   $\boldsymbol{w}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}/\|\boldsymbol{w}^{t}_{j}\|$ ;
+                    // Normalization
+7
+            $\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2},~{}~{}(j=1,\dots,N)$ ;
+              // SGD step
+8
+9
+10
+11Initialize $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$.
+12Phase II: SGD on second-layer parameters
+       $\hat{\boldsymbol{a}}\leftarrow\mathrm{argmin}_{\boldsymbol{a}\in\mathbb{R}^{N}}\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\mathopen{}\mathclose{{}\left(f_{\boldsymbol{\Theta}}(\boldsymbol{x}_{i})-y_{i}}\right)^{2}+\lambda\|\boldsymbol{a}\|^{2}$ ;
+        // Ridge regression
+13
+14
 Output : Prediction function $\boldsymbol{x}\mapsto f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})$ with $\hat{\boldsymbol{\Theta}}=(\hat{a}_{j},\boldsymbol{w}^{T_{1}}_{j},b_{j})_{j=1}^{N}$.
 Algorithm 1 Gradient-based training of two-layer neural network
-3.2 Convergence and Statistical Complexity
+3.2 Convergence and Sample Complexity
 Weak Recovery Guarantee.
@@
 We require the activation function to be a polynomial $\sigma(z)=\sum_{i=0}^{C_{\sigma}}\beta_{i}\mathsf{He}_{i}(z)$ and its degree $C_{\sigma}$ to be sufficiently large so that $C_{\sigma}\geq C_{q}$ holds ($C_{q}$ is defined in Proposition 6).
-For all $2\leq\ell\leq C_{\sigma}$ and $k=0,1$, we assume that $H\big{(}\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k\big{)}\neq 0$.
-More precisely, a given $\sigma_{*}$, the above condition only needs to be met for one pair of $(k,\ell)$, as specified in Appendix B.1.
-We show that this condition is satisfied for a wide range of polynomials with degree $C_{\sigma}$.
+For all $2\leq\ell\leq C_{\sigma}$ and $k=0,1$, we assume that $H\big{(}\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k\big{)}>0$.
+As discussed in Appendix B.1, for a given $\sigma_{*}$, the above assumption only needs to be met for one pair of $(k,\ell)$.
+Appendix B.1.3 states that $H\big{(}\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k\big{)}\neq 0$ also suffices if we set the momentum parameter $\xi$ differently.
+Now we verify this condition for a wide range of polynomial activations.
 Lemma 3.
-Given any $\ell\geq 2$ and $k\geq 0$.
-For $C_{\sigma}\geq\frac{2\ell+k-1}{\ell}$, if we choose $\{\beta_{i}\}_{i=0}^{C_{\sigma}}$ where each $\beta_{i}$ is randomly drawn from some interval $[a_{i},b_{i}]$ with $a_{i}<b_{i}$, then $H(\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k)\neq 0$ with probability one.
-The following theorem states that $n=\tilde{\Theta}(d)$ samples are sufficient to achieve weak recovery.
+Given $\ell\geq 2$ and $k\geq 0$.
+For $C_{\sigma}\geq\frac{2\ell+k-1}{\ell}$, if we choose $\{\beta_{i}\}_{i=0}^{C_{\sigma}}$ where $\beta_{i}$ is randomly drawn from some non-empty interval $[a_{i},b_{i}]$, then $H(\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k)\neq 0$ with probability 1.
+The next theorem states that $n=\tilde{\Theta}(d)$ samples are sufficient for SGD to achieve weak recovery.
 Theorem 1.
-Under Assumptions 1 and 2, for suitable choices of hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $\xi_{j}^{t}=O(1)$, there exists constant $C(q)$ such that after Phase I of Algorithm 1 is run for $2T_{1,1}=C(q)\cdot d\mathrm{polylog}(d)$ steps, with high probability, there exists a subset of neurons $\boldsymbol{w}_{j}^{2T_{1}}\in\mathcal{W}$ with $|\mathcal{W}|=\tilde{\Theta}(N)$ such that $\big{|}\langle\boldsymbol{w}_{j}^{2T_{1}},\boldsymbol{\theta}\rangle\big{|}>c$ for some $c\gtrsim 1/\mathrm{polylog}(d)$.
+Under Assumptions 1 and 2, for suitable choices of hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $1-\xi^{t}=o_{d}(1)$, there exists constant $C(q)$ such that after Phase I of Algorithm 1 is run for $2T_{1,1}=C(q)\cdot d\mathrm{polylog}(d)$ steps, with high probability, there exists a subset of neurons $\boldsymbol{w}_{j}^{2T_{1}}\in\mathcal{W}$ with $|\mathcal{W}|=\tilde{\Theta}(N)$ such that $\big{|}\langle\boldsymbol{w}_{j}^{2T_{1}},\boldsymbol{\theta}\rangle\big{|}>c$ for some $c\gtrsim 1/\mathrm{polylog}(d)$.
 Recall that at random initialization we have $\langle\boldsymbol{w},\boldsymbol{\theta}\rangle\approx d^{-1/2}$ with high probability.
@@
 Assumption 3.
-Recall the Hermite expansions $\sigma_{*}(z)=\sum_{i=p}^{q}\alpha_{i}\mathsf{He}_{i}(z)$, $\sigma_{j}(z)=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, we assume the coefficients satisfy $\alpha_{i}\beta_{j,i}\geq 0$ for $p\leq i\leq q$.
+Given the Hermite expansions $\sigma_{*}(z)=\sum_{i=p}^{q}\alpha_{i}\mathsf{He}_{i}(z)$, $\sigma_{j}(z)=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, we assume the coefficients satisfy $\alpha_{i}\beta_{j,i}\geq 0$ for $p\leq i\leq q$.
 This assumption is easily verified in the well-specified setting $\sigma_{*}=\sigma$ [BAGJ21] since $\alpha_{i}=\beta_{i}$, and under link misspecification, it has been directly assumed in prior work [MHWSE23].
-We follow [OSSW24] and show that by randomizing the Hermite coefficients of the activation function, a subset of neurons satisfy the above assumption for any degree-$q$ polynomial link function $\sigma_{*}$.
+We follow [OSSW24a] and show that by randomizing the Hermite coefficients of the activation function, a subset of neurons satisfy the above assumption for any degree-$q$ polynomial link function $\sigma_{*}$.
 Lemma 4.
 If we set $\sigma_{j}(z)=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, where for each neuron we sample $\beta_{j,i}\stackrel{{\scriptstyle\mathrm{\tiny{i.i.d.}}}}{{\sim}}\mathrm{Unif}(\{\pm r_{i}\})$ with appropriate constant $r_{i}$, then Assumption 2 and 3 are satisfied in $\operatorname{exp}\mathopen{}\mathclose{{}\left({-\Theta(q)}}\right)$-fraction of neurons.
-The proof is found in Appendix B.1.
-Note that in our construction of activation functions for both assumptions, we do not exploit knowledge of the link function $\sigma_{*}$ other than its degree $q$ which decides the constant $C_{\sigma}$.
-The next theorem shows that by running Phase I for $\tilde{\Theta}(d)$ more steps, a subset of neurons can achieve almost perfect overlap with the index features.
+Note that in our construction of activation functions for both assumptions, we do not exploit knowledge of the link function $\sigma_{*}$ other than its degree $q$ which decides the constant $C_{\sigma}$; see Appendix B.1 for more discussion of Assumption 3 and Lemma 4.
+The next theorem shows that by running Phase I for $\tilde{\Theta}(d)$ more steps, a subset of neurons achieves sufficiently large overlap with the index features.
 Theorem 2.
-For student neurons with activation satisfying Assumptions 2 and 3 and parameter $\boldsymbol{w}_{j}$ starting from a nontrivial overlap $c$ specified in Theorem 1, for suitable hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $\xi_{j}^{t}=1-\tilde{O}(\varepsilon)$, if we continue to run Phase I of Algorithm 1 for $2T_{1,2}=\tilde{\Theta}_{d}(d\varepsilon^{-2})$ steps, then we achieve $\big{\langle}\boldsymbol{w}_{j}^{2(T_{1,1}+T_{1,2})},\boldsymbol{\theta}\big{\rangle}>1-\varepsilon$ with high probability.
-Combining Theorem 1,2 and Lemma 3, we know that after $T_{1}=2(T_{1,1}+T_{1,2})$ steps, there exists some constant $C(q)$ depending only on the degree $q$ such that $1/(C(q)\cdot\mathrm{polylog}(d))$ fraction of the neurons become $\varepsilon$-close to the target direction $\boldsymbol{\theta}$.
+For student neurons satisfying Assumptions 2, 3 and parameter $\boldsymbol{w}_{j}$ starting from nontrivial overlap $c>0$ specified in Theorem 1, if Phase I of Algorithm 1 continues for $2T_{1,2}=\tilde{\Theta}_{d}(d\varepsilon^{-2})$ steps with hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1}\varepsilon)$, $\xi^{t}=1$, we achieve $\big{\langle}\boldsymbol{w}_{j}^{2(T_{1,1}+T_{1,2})},\boldsymbol{\theta}\big{\rangle}>1-\varepsilon$ with high probability.
 The following proposition shows that after strong recovery, training the second-layer parameters in Phase II is sufficient for the NN model (3.1) to achieve small generalization error.
@@
 After Phase I terminates, for suitable $\lambda>0$, the output of Phase II satisfies
-$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon.$
-with probability 1 as $d\to\infty$, if we set $T_{2}=C(q)N^{4}\mathrm{polylog}(d)\varepsilon^{-2}$, $N=C(q)\mathrm{polylog}(d)\varepsilon^{-1}$ for some constant $C(q)$ depending on the target degree $q$.
+$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon^{2}.$
+with probability 1 as $d\to\infty$, if we set $T_{2}=C(q)N^{4}\mathrm{polylog}(d)\varepsilon^{-4}$, $N=C(q)\mathrm{polylog}(d)\varepsilon^{-1}$ for some constant $C(q)$ depending on the target degree $q$.
 Putting things together.
 Combining the above theorems, we conclude that in order for two-layer NN (3.1) trained by Algorithm 1 to achieve $\varepsilon$ population squared loss, it is sufficient to set
-$n=T_{1}+T_{2}\asymp C(q)\cdot(d\varepsilon^{-2}\vee\varepsilon^{-6})\mathrm{polylog}(d),\quad N\asymp C(q)\cdot\varepsilon^{-1}\mathrm{polylog}(d),$ (3.2)
+$n=T_{1}+T_{2}\asymp C(q)\cdot(d\varepsilon^{-2}\vee\varepsilon^{-8})\mathrm{polylog}(d),\quad N\asymp C(q)\cdot\varepsilon^{-1}\mathrm{polylog}(d),$ (3.2)
 where constant $C(q)$ only depends on the target degree $q$ (although exponentially).
 Hence we may set $\varepsilon^{-1}\asymp\mathrm{polylog}d$ to conclude an almost-linear sample and computational complexity for learning arbitrary single-index polynomials up to $o_{d}(1)$ population error.
-This establishes the informal theorem in Section 1.
 4 Proof Sketch
@@
 If we ignore desideratum $(ii)$, then for polynomial link functions, transformations that decrease the information exponent to at most $2$ have been constructed in [CM20, Section 2.1].
-However, prior results are based on the thresholding function with specific offset, and it is not clear if such function naturally arises from SGD with batch reuse.
-The following proposition shows that the effect of thresholding can also be achieved by a simple monomial transformation.
+However, prior results are based on the thresholding function, and it is not clear if such function naturally arises from SGD with batch reuse.
+The following proposition shows that the effect of thresholding can also be achieved by a simple monomial transformation where the required degree can be uniformly upper bounded.
 Proposition 6.
-Let $f:\mathbb{R}\to\mathbb{R}$ be any polynomial with degree up to $p$ and $\mathopen{}\mathclose{{}\left\|{f}}\right\|_{L^{2}(\gamma)}^{2}=1$, then
+Let $g:\mathbb{R}\to\mathbb{R}$ be any polynomial with degree up to $p$ and $\mathopen{}\mathclose{{}\left\|{g}}\right\|_{L^{2}(\gamma)}^{2}=1$, then
   • (i)
-    There exists some $i\leq C_{q}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(f^{i})\leq 2$, where constant $C_{q}$ only depends on $q$.
+    There exists some $i\leq C_{q}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(g^{i})\leq 2$, where constant $C_{q}$ only depends on $q$.
   • (ii)
-    Let $f^{\mathrm{odd}}:\mathbb{R}\to\mathbb{R}$ be the odd part of $f$ with $\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{\mathrm{odd}}(t)^{2}]\geq\rho>0$.
-Then there exists some $i\leq C_{q,\rho}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(f^{i})=1$, where constant $C_{q,\rho}$ only depends on $q$ and $\rho$.
-We make the following remarks.
-  • The proposition implies that for any polynomial link function that is not even, there exists some power $i\in\mathbb{N}_{+}$ only depending on the degree of $\sigma_{*}$ such that raising the function to the $i$-the power reduces the information exponent to $1$.
-For even link functions, the information exponent after arbitrary transformation is at least $2$, and this lowest value can also be achieved by monomial transformation.
+    Let $g^{\mathrm{odd}}:\mathbb{R}\to\mathbb{R}$ be the odd part of $g$ with $\|g^{\mathrm{odd}}\|_{L^{2}(\gamma)}^{2}\geq\rho>0$.
+Then there exists some $i\leq C_{q,\rho}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(g^{i})=1$, where constant $C_{q,\rho}$ only depends on $q$ and $\rho$.
+The proof can be found in Appendix A. We make the following remarks.
+  • The proposition implies that for any polynomial link function that is not even, there exists $i\in\mathbb{N}_{+}$ only depending on the degree of $\sigma_{*}$ such that raising the function to the $i$-th power reduces the information exponent to $1$ ($p_{*}=1$).
+For even $\sigma_{*}$, the information exponent after arbitrary transformation is at least $2$ ($p_{*}=2$), which can also be attained by monomial transformation.
 Furthermore, we provide a uniform upper-bound on the required degree of transformation $i$ via a compactness argument.
@@
 $\langle\boldsymbol{\theta},\boldsymbol{w}^{2}\rangle=\mathopen{}\mathclose{{}\left\langle\boldsymbol{\theta},\big{[}\boldsymbol{w}^{1}+\eta\cdot y\sigma^{\prime}(\langle\boldsymbol{x},\boldsymbol{w}^{1}\rangle)\boldsymbol{x}\big{]}}\right\rangle=\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle+$ (4.1)
-$\eta\bigg{[}y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle+\sum_{i=0}^{C_{\sigma}-1}\underbrace{(\eta\|\boldsymbol{x}\|^{2})^{i}y^{i+1}(i!)^{-1}(\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{i}\sigma^{(i+1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle}_{=:\psi_{i}}\bigg{]}.$ (4.2)
-We take $\eta\leq c_{\eta}d^{-1}$ with a small constant $c_{\eta}$ so that $\eta\|\boldsymbol{x}\|^{2}\ll 1$.
-Crucially, the strength of each term in (4.2) can vary depending on properties of the link function $\sigma_{*}$, which is unknown.
-Hence a careful analysis is required to ensure that the suitable monomial transformation is always singled out from the gradient update.
-We therefore divide our analysis into four cases (for simplicity we present the noiseless setting below).
-  • (I)
-    If $\mathrm{IE}(\sigma_{*})=1$.
-All terms in the summation in (4.2) with $i\geq 2$ decay as fast as $\eta\|\boldsymbol{x}\|^{2}\ll 1$.
-On the other hand, the expectation of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ is roughly $\alpha_{1}\beta_{1}\gtrsim 1$.
-Therefore we may isolate the informative term $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$.
-This case is discussed in Section B.3.3.
-  • (II)
-    Else if $\mathrm{IE}((\sigma_{*})^{I})=1$ for some $2\leq I\leq C_{\sigma}$.
-Let $I$ be the lowest degree of monomial transformation such that $\mathrm{IE}((\sigma_{*})^{I})=1$.
-Since $\sigma_{*},\cdots,\sigma_{*}^{I-1}$ have information exponent larger than $1$, expectations of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ and $\psi_{i}\ (i=2,\cdots,I-2)$ scales as $\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle\asymp d^{-1/2}$.
-For $i=I-1$, because $\sigma_{*}^{I}$ has information exponent $1$,
-    $\mathbb{E}[\psi_{I-1}]=\,$ $\mathbb{E}\big{[}(\eta\|\boldsymbol{x}\|^{2})^{I-1}y^{I}((I-1)!)^{-1}(\sigma^{(1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{I-1}\sigma^{(I)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle\big{]}$ (4.3)
+$\eta\bigg{[}\textstyle y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle+\sum_{i=0}^{C_{\sigma}-1}\underbrace{\textstyle(\eta\|\boldsymbol{x}\|^{2})^{i}y^{i+1}(i!)^{-1}(\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{i}\sigma^{(i+1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle}_{=:\psi_{i}}\bigg{]}.$ (4.2)
+We take $\eta\leq c_{\eta}d^{-1}$ with a small constant $c_{\eta}$ so that $\eta\|\boldsymbol{x}\|^{2}\ll 1$ with high probability.
+Crucially, the strength of each term in (4.2) can vary depending on properties of the unknown link function $\sigma_{*}$.
+Hence a careful analysis is required to ensure that a suitable monomial transformation is always singled out from the gradient.
+We establish the following lemma on the evolution of alignment.
+Lemma 7.
+Under the assumptions per Theorem 1, the following holds for generative exponent $p_{*}=1,2$:
+$\langle\boldsymbol{\theta},\boldsymbol{w}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}^{2t}\rangle+c_{\eta}^{I}c_{\xi}c_{\sigma}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\nu^{2t}.$ (4.3)
+See Lemma 16 for the formal version.
+For $p_{*}=1$, taking expectation immediately yields that weak recovery within $(\eta(1-\xi)\gamma)^{-1}=O(d)$ steps.
+For $p_{*}=2$, $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle=:\kappa^{t}$ can be approximated by a differential equation $\frac{\mathrm{d}\kappa^{t}}{\mathrm{d}t}=\eta(1-\xi)\gamma\kappa^{t}$.
+Solving this yields $\kappa^{t}=\kappa^{0}\exp(\eta(1-\xi)\gamma t)\approx d^{-\frac{1}{2}}\exp(\eta(1-\xi)\gamma t)$, and weak recovery is obtained within $t\lesssim(\eta(1-\xi)\gamma)^{-1}\cdot\log d=O(d\log d)$ steps, similar to the analysis in [BAGJ21].
+Why interpolation is needed.
 […]
-    For $i\geq I$, $\psi_{i}$ decays as $c_{\eta}^{I}$, which is smaller than the scale of $\psi_{I-1}\asymp c_{\eta}^{I-1}$.
-Hence the term $\psi_{I-1}\gtrsim c_{\eta}^{I-1}$ is singled out.
-This case is discussed in Section B.3.2.
-  • (III)
-    Else if $\mathrm{IE}(\sigma_{*})=2$.
-We have $\mathbb{E}[y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle]\approx 2\alpha_{2}\beta_{2}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$.
-Also, for $i\geq 2$, since $\sigma_{*}^{2},\cdots,\sigma_{*}^{C_{\sigma}}$ have information exponent at least $2$, expectation of $\psi_{i}$ is roughly of order $(\eta\|\boldsymbol{x}\|^{2})^{i}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$.
-Therefore, the term $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ is singled out, and the expectation scales as $\alpha_{2}\beta_{2}d^{-1/2}$ at initialization.
-This case is discussed in Section B.3.3.
-  • (IV)
-    Else $\mathrm{IE}((\sigma_{*})^{I})=2$ for some $2\leq I\leq C_{\sigma}$.
-In this case, since $\sigma_{*},\cdots,\sigma_{*}^{I-1}$ have information exponent larger than $2$, expectations of $y\sigma^{\prime}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle$ and $\psi_{i}\ (i=2,\cdots,I-2)$ are at most $(\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle)^{2}\asymp d^{-1}$.
-And at $i=I-1$, because $\sigma_{*}^{I}$ has information exponent $2$,
-    $\mathbb{E}[\psi_{I-1}]=\,$ $\mathbb{E}\big{[}(\eta\|\boldsymbol{x}\|^{2})^{I-1}y^{I}((I-1)!)^{-1}(\sigma^{(1)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle))^{I-1}\sigma^{(I)}(\langle\boldsymbol{\omega},\boldsymbol{x}\rangle)\langle\boldsymbol{\theta},\boldsymbol{x}\rangle\big{]}$ (4.5)
-[…]
-    As for $i\geq I$, because $\sigma_{*}^{i}$ has information exponent larger than $1$ for $I+1\leq i\leq C_{\sigma}$, $\psi_{i}$ decays as $c_{\eta}^{I}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$, which is smaller than $\psi_{I-1}\asymp c_{\eta}^{I-1}\langle\boldsymbol{\theta},\boldsymbol{\omega}\rangle$.
-Thus, the term $\psi_{I-1}$ is dominating, whose scale is roughly $c_{\eta}^{I-1}d^{-1/2}$ at initialization.
-This case is discussed in Section B.3.1.
-Why interpolation is required.
-[…]
-Combining the four cases yields the following lemma on the evolution of alignment.
-Lemma 6.
-Under the assumptions per Theorem 1, one of the following holds:
-  • (i)
-    $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\gamma+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\nu^{2t}_{j}.$
-  • (ii)
-    $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2(t+1)}\rangle\geq\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\gamma\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle+\eta^{2t}_{j}(1-\xi_{j}^{2(t+1)})\nu^{2t}_{j}.$
-Here $\gamma\gtrsim c_{\eta}^{I-1}$ is a constant that depends on $\sigma_{*}$ and $\nu^{2t}_{j}$ is a mean-zero noise.
-For (i), taking expectation immediately yields that weak recovery is achieved within $(\eta(1-\xi)\gamma)^{-1}=\tilde{O}(d)$ steps.
-For (ii), $\langle\boldsymbol{\theta},\boldsymbol{w}_{j}^{2t}\rangle=:\kappa^{t}$ can be approximated by a differential equation $\frac{\mathrm{d}\kappa^{t}}{\mathrm{d}t}=\eta(1-\xi)\gamma\kappa^{t}$.
-Solving this yields $\kappa^{t}=\kappa^{0}\exp(\eta(1-\xi)\gamma t)\approx d^{-\frac{1}{2}}\exp(\eta(1-\xi)\gamma t)$, and weak recovery is obtained within $t\lesssim(\eta(1-\xi)\gamma)^{-1}\cdot\log d=\tilde{O}(d)$ steps, similar to the analysis in [BAGJ21].
 4.3 Analysis of Phase II and Statistical Guarantees
@@
 Since the objective is strongly convex, gradient-based optimization can efficiently minimize the empirical loss.
-The learnability guarantee follows from standard analysis analogous to that in [AAM22, DLS22, BES+22], where we construct a “certificate” second-layer $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ that achieves small loss and small norm:
+In Appendix B.6, the learnability guarantee follows from standard analysis analogous to that in [AAM22, DLS22, BES+22], where we construct a “certificate” second-layer $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ that achieves small loss and small norm:
 $\mathbb{E}_{\boldsymbol{x}}\mathopen{}\mathclose{{}\left(f_{*}(\boldsymbol{x})-\frac{1}{N}\sum_{j=1}^{N}a^{*}_{j}\sigma_{j}\big{(}\langle{\boldsymbol{w}_{j}^{T_{1}}},\boldsymbol{x}\rangle+b_{j}\big{)}}\right)^{2}\leq\varepsilon^{*},\quad\|\boldsymbol{a}^{*}\|\lesssim r^{*},$
 from which the population loss of the regularized empirical risk minimizer can be bounded via standard Rademacher complexity argument.
-To construct such a certificate, we make use of the random bias units $\{b_{j}\}_{j=1}^{N}$ to approximate the link function $\sigma_{*}$ as done in [DLS22, BBSS22, OSSW24].
+To construct such a certificate, we make use of the random bias units $\{b_{j}\}_{j=1}^{N}$ to approximate the link function $\sigma_{*}$ as done in [DLS22, BBSS22, OSSW24a].
+5 Beyond Polynomial Link Functions
+Thus far we have shown that for polynomial single-index target functions (which satisfy $p_{*}\leq 2$), SGD with data reuse can implement a polynomial transformation to the labels that reduces the information exponent to at most 2; consequently, the trained two-layer neural network can achieve small generalization error with $n=d\operatorname{polylog}(d)$ samples.
+However, as shown in [DPVLB24], there exists (non-polynomial) $\sigma_{*}$ with generative exponent $p_{*}>2$ (i.e., label transformations cannot lower the information exponent to $2$) and thus not learnable by SQ algorithms in linear sample complexity.
+Nevertheless, for a single-index model with generative exponent $p_{*}$, we know there exists an “optimal” label transformation that reduces the information exponent to $p_{*}$.
+If SGD can make use of such transformation, then from the arguments in [BAGJ21], it is natural to conjecture that a sample size of $n\simeq d^{p_{*}-1}$ is sufficient.
+In this section we confirm this intuition by proving that SGD with data reuse (Algorithm 1) indeed matches this complexity.
+The following lemma is an analogue of Proposition 6 stating that polynomial transformations are sufficient to lower the information exponent.
+Lemma 8.
+Given link function $\sigma_{*}$ with generative exponent $p_{*}\in\mathbb{N}_{+}$.
+Suppose we can take an orthonormal polynomial basis $\{\phi_{k}\}_{k}$ for the space $L^{2}(P_{y})$ with inner product $\langle f,g\rangle=\mathbb{E}_{y=\sigma_{*}(z)}[f(y)g(y)]$.
+Then there exists some degree of transformation $I\in\mathbb{N}_{*}$ such that $\mathrm{IE}(\sigma_{*}^{I})=p_{*}$.
+We outline the differences and additional technical challenges to handle the $\mathrm{GE}(\sigma_{*})>2$ setting.
+  • For general $L^{2}$ link functions $\sigma_{*}$, we can no longer make use of the compactness argument (see proof of Proposition 6) to upper bound the degree of monomial transformation.
+Hence in Lemma 8 we do not state a uniform upper bound on the required degree $I$, unlike the polynomial setting.
+  • Any link function with $p_{*}>2$ cannot be polynomial, and hence we cannot achieve low generalization error using a neural network with polynomial nonlinearity.
+We therefore need to use an activation function with universal function approximation ability.
+5.1 Sample Complexity for Weak Recovery
+We first show that Algorithm 1 achieves weak recovery (i.e., nontrivial overlap with the ground truth $\boldsymbol{\theta}$) with a complexity governed by the generative exponent of the link function $p_{*}=\mathrm{GE}(\sigma_{*})$.
+Similar to Section 3.2, we make use of randomized activation functions to ensure the desired label transformation is encoded — we defer the conditions on the student activation to Appendix B.1.2.
+Similar to Theorem 1, we focus on the subset of neurons with large initial overlap, and activation satisfying the assumptions in Appendix B.1.2 (these conditions are met by $\tilde{\Omega}(1)$ fraction of neurons).
+Proposition 9.
+Suppose the link function $\sigma_{*}$ has generative exponent $p_{*}$, and let $I\in\mathbb{N}_{+}$ be the smallest degree of monomial transformation that lowers the information exponent to $p_{*}$ (i.e., $\mathrm{IE}(\sigma_{*}^{I})=p$).
+We can find a student activation function $\sigma$ depending only on $p,p_{*}$ and $I$, such that if we take $\eta^{2t},\eta^{2t+1}=c_{\eta}Nd^{-1}$, $\xi^{2(t+1)}=1-c_{\xi}d^{-\nicefrac{{(p_{*}-2)_{+}}}{{2}}}$ for small $c_{\eta},c_{\xi}=o_{d}(1)$, and set
+$T_{1,1}\simeq c_{\xi}^{-1}\begin{cases}d&(\text{if $p_{*}=1$})\\ d(\log d)&(\text{if $p_{*}=2$})\\ d^{p_{*}-1}&(\text{if $p_{*}\geq 3$}),\end{cases}$ (5.1)
+then if the initial alignment $\langle\boldsymbol{w}^{0},\boldsymbol{\theta}\rangle\geq 2c_{\eta}^{-1}d^{-\nicefrac{{1}}{{2}}}$, there exists $\tau_{*}\leq T_{1,1}$ such that for all $\tau\geq\tau_{*}$,
+$\langle\boldsymbol{w}^{2\tau},\boldsymbol{\theta}\rangle\geq\tilde{\Theta}(1),\quad\text{with probability $1-o_{d}(1)$.}$ (5.2)
+Proposition 9 is a generalization of Theorem 1 beyond polynomial $\sigma_{*}$ (the proof of both results are presented in Appendix B.3,B.4), and can be interpreted as an SQ counterpart to [BAGJ21]: we establish a sufficient sample size of $n\simeq d^{(p_{*}-1)\vee 1}$ for Algorithm 1 to exit the search phase, which is parallel to the $n\simeq d^{(p-1)\vee 1}$ rate for one-pass SGD (note that our rates are slightly sharper due to logarithmic factors removed, since $c_{\xi}^{-1}$ can grow arbitrarily slowly with $d$).
+For high generative exponent $\sigma_{*}$ with $p_{*}>2$, we no longer match the information theoretically optimal sample complexity $n\asymp d$, which is consistent with the computational-to-statistical gap observed in [DPVLB24].
+5.2 Generalization Error Guarantee
+After Phase I of Algorithm 1, we learn the unknown link function $\sigma_{*}$ via training the second-layer.
+To approximate non-polynomial functions, we introduce a ReLU component in the student nonlinearity $\sigma$ (see Lemma 12 for discussions), and make use of the approximation result for the (univariate) ReLU kernel in [BBSS22], which handles general $\sigma_{*}$ whose second derivative has bounded 4th moment.
+Combining the above, we arrive at the following end-to-end guarantee for learning single-index models with arbitrary generative exponent using SGD training of neural network.
+Proposition 10 (Informal).
+Suppose the link function $\sigma_{*}$ has generative exponent $p_{*}\in\mathbb{N}_{*}$ and satisfies $\sigma_{*},\sigma_{*}^{\prime\prime}\in L^{4}(\gamma)$.
+For appropriately chosen activation function $\sigma$ (see Appendix B.1.2), a neural network (3.1) with $N=\tilde{\Theta}(1)$ neurons optimized by Algorithm 1 achieves small population loss $\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]=o_{d,\mathbb{P}}(1)$ with a sample complexity of $n=\tilde{\Theta}(d^{(p_{*}-1)\vee 1})$.
+See Appendix B.6 for the full statement with $\varepsilon$ dependence.
+This proposition confirms that the sample complexity for weak recovery (Proposition 9) is the bottleneck in single-index learning, as the total sample size required for Algorithm 1 to achieve low test error also scales with $d^{(p_{*}-1)\vee 1}$.
 6 Conclusion and Future Directions
-In this work we showed that a two-layer neural network (3.1) trained by SGD with reused batch can learn arbitrary single-index polynomials up to $\varepsilon$ population error using $n=\tilde{O}(d\varepsilon^{-2})$ samples and compute.
-Our analysis is based on the observation that by reusing the same minibatch twice in the gradient computation, a non-correlational term arises in the SGD update that transforms the labels (despite the loss function is not modified).
-Specifically, following the definition in [DPVLB24], we know that polynomial $\sigma_{*}$ has generative exponent at most $2$, which implies the existence of nonlinear transformation $\mathcal{T}:\mathbb{R}\to\mathbb{R}$ such that the information exponent $p_{*}$ becomes at most $2$, i.e.,
-$\mathbb{E}[\mathcal{T}(y)\mathsf{He}_{i}(\langle\boldsymbol{x},\boldsymbol{\theta}\rangle)]\neq 0,\quad\text{for $i=1$ or $2$.}$
-We show that restricting $\mathcal{T}$ to be polynomial is sufficient, and such transformation can be extracted by Taylor-expanding the SGD update.
-Then we show via careful analysis of the trajectory that strong recovery and low population error can be achieved under suitable activation function.
-Future directions.
-First, our analysis only handles link functions with generative exponent $p_{*}\leq 2$; while this covers arbitrary polynomial $\sigma_{*}$ analogous to [CM20], it is interesting to examine whether SGD with reused batch can learn targets with $p_{*}\geq 3$ with a sample complexity matching the SQ lower bound.
-It is also possible that ERM algorithms on i.i.d.
-data can achieve a statistical complexity beyond the SQ lower bound due to non-adversarial noise [DH21, DH24]; such mechanism is not exploited in our current analysis.
-Additional interesting directions include extension to multi-index [BAGJ22, BBPV23, CWPPS23, Gla23], hierarchical polynomials [NDL23], and additive models [OSSW24].
-Lastly, the SGD algorithm that we employ requires a layer-wise training procedure and a specific batch reuse schedule; one may therefore ask if standard multi-pass SGD training of all parameters simultaneously (as reported in Figure 1) also achieves the same statistical efficiency.
+We showed that a two-layer neural network (3.1) trained by SGD with reused batch can learn single-index model (with generative exponent $p_{*}$) using $n\simeq d^{(p_{*}-1)\vee 1}$ samples and compute; in particular, when the link function $\sigma_{*}$ is polynomial, we established a sample complexity of $n=\tilde{O}(d\varepsilon^{-2})$ to achieve $\varepsilon$ population loss, which is almost information theoretically optimal.
+Our analysis is based on the observation that by reusing the same training data twice in the gradient computation, a non-correlational term arises in the SGD update that transforms the labels (despite the loss function not modified).
+We proved that monomial transformations that lower the information exponent of $\sigma_{*}$ can be extracted by Taylor-expanding the SGD update; then we showed via careful analysis of the trajectory that strong recovery and low population loss is achieved under suitable activation function.
+Interesting future directions include extension to multi-index models [BAGJ22, BBPV23, CWPPS23], hierarchical target functions [AZL19, NDL23], and in-context learning [OSSW24b].
+Also, the SGD algorithm that we employ requires a layer-wise training procedure and a specific batch reuse schedule; one may therefore ask if standard multi-pass SGD training of all parameters simultaneously [Gla23] (as reported in Figure 1) also achieves the same statistical efficiency.
 Acknowledgements
@@
 TS was partially supported by JSPS KAKENHI (24K02905) and JST CREST (JPMJCR2015).
+This research is unrelated to DW’s work at xAI.
 Appendix A Polynomial Transformation
@@
 Consider the following expectation:
-$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}.$ (A.1)
+$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}.$ (A.1)
 We evaluate the case when $i$ is even. (A.1) can be lower bounded as
 […]
-$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.3)
-$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.4)
-$\geq$ $~{}\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|\geq 2\tau(f)]\bigg{(}\frac{2\tau(f)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.5)
-$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.6)
-$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{\tau(f)}{2\tau(f)}\bigg{)}^{i}(0^{2}-1)\bigg{]}$ (A.7)
-$\geq$ $~{}3\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[|f(t)|\geq 2\tau(f)]-2^{-i}.$ (A.8)
+$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.3)
+$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(t^{2}-1)\bigg{]}$ (A.4)
+$\geq$ $~{}\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|f(t)|\geq 2\tau(f)]\bigg{(}\frac{2\tau(f)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.5)
+$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[\tau(f)\leq|f(t)|<2\tau(f)]\bigg{(}\frac{f(t)}{2\tau(f)}\bigg{)}^{i}(2^{2}-1)\bigg{]}$ (A.6)
+$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|f(t)|<\tau(f)]\bigg{(}\frac{\tau(f)}{2\tau(f)}\bigg{)}^{i}(0^{2}-1)\bigg{]}$ (A.7)
+$\geq$ $~{}3\mathbb{P}_{t\sim\mathcal{N}(0,1)}[|f(t)|\geq 2\tau(f)]-2^{-i}.$ (A.8)
 Note that $\mathbb{P}[|f(t)|\geq 2\tau(f)]$ is positive (since $f$ is polynomial) and independent of $i$, while $2^{-i}$ decays to 00 as $i$ increases.
@@
 (i-4): Conclusion via hypercontractivity. Because $f$ is a polynomial with degree at most $q$, Gaussian hypercontractivity [O’D14] yields that
-$2H(f^{i};2)^{2}\leq\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}(f(t))^{2i}\big{]}\leq(2i-1)^{iq}\big{(}\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}f(t)^{2}\big{]}\big{)}^{i}=(2i-1)^{iq}.$ (A.10)
+$2H(f^{i};2)^{2}\leq\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}(f(t))^{2i}\big{]}\leq(2i-1)^{iq}\big{(}\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}f(t)^{2}\big{]}\big{)}^{i}=(2i-1)^{iq}.$ (A.10)
 Therefore, for all polynomials in $\mathcal{F}_{q}$, a partial sum of (A.9) is uniformly bounded by
@@
 Consider the following expectation:
-$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\bigg{(}\frac{g(t)}{2\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.13)
+$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\bigg{(}\frac{g(t)}{2\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.13)
 (A.13) is decomposed as
 […]
-$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[2\tau(g)\leq|g(t)|<3\tau(g)]\bigg{(}\frac{g(t)}{3\tau(f)}\bigg{)}^{i}t\bigg{]}$ (A.15)
-$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.16)
+$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[2\tau(g)\leq|g(t)|<3\tau(g)]\bigg{(}\frac{g(t)}{3\tau(f)}\bigg{)}^{i}t\bigg{]}$ (A.15)
+$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.16)
 We first evaluate the first term.
@@
 Combining the above, the first term of (A.16) is bounded as
-$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|g(t)|\geq 3\tau(f)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.17)
-$=\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\beta\leq t\leq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.18)
-$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\leq\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.19)
-$=\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[\beta\leq t\leq-\alpha]t\big{]}+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}-\bigg{(}\frac{g(-t)}{3\tau(g)}\bigg{)}^{i}\bigg{)}t\bigg{]}$ (A.20)
-$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\beta\leq t\leq-\alpha\big{]}.$ (A.21)
+$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|g(t)|\geq 3\tau(f)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.17)
+$=\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[\beta\leq t\leq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.18)
+$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[t\leq\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.19)
+$=\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[\beta\leq t\leq-\alpha]t\big{]}+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}-\bigg{(}\frac{g(-t)}{3\tau(g)}\bigg{)}^{i}\bigg{)}t\bigg{]}$ (A.20)
+$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,1)}\big{[}\beta\leq t\leq-\alpha\big{]}.$ (A.21)
 Following the exact same reasoning, we know that the second term of (A.16) is positive. Finally, the third term which is bounded by
-$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\big{]}\geq-\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.22)
+$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\big{]}\geq-\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.22)
 Putting things together,
-(A.16)$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\beta\leq t\leq-\alpha\big{]}-\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.23)
+(A.16)$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,1)}\big{[}\beta\leq t\leq-\alpha\big{]}-\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.23)
 The first term is independent of $i$ and positive, while the second term goes to zero as $i$ grows. Therefore, there exists some $i$ such that $\mathrm{IE}(g^{i};1)=1$.
@@
 (ii-3): Lower bound of test function via compactness.
-Let $\mathcal{F}_{q}$ be a set of unit $L^{2}$-norm polynomials with degree up to $q$ and $\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{\mathrm{odd}}(t)^{2}]\geq c$.
+Let $\mathcal{F}_{q}$ be a set of unit $L^{2}$-norm polynomials with degree up to $q$ and $\mathbb{E}_{t\sim\mathcal{N}(0,1)}[f^{\mathrm{odd}}(t)^{2}]\geq c$.
 Since $\mathscr{H}(f)$ is always positive for $\mathcal{F}_{q}$, $\mathscr{H}(f)$ is continuous with respect to $f$, and $\mathcal{F}_{q}$ is a compact set, $\inf_{f\in\mathcal{F}_{q}}\mathscr{H}(f)$ has the minimum value $\mathscr{H}_{0}$ that is positive.
@@
 Because $\mathscr{H}_{0}$ depends on $c$, $C_{q,c}$ depends on $c$ as well as $q$. ∎
+A.3 Proof for Non-Polynomial Functions
+For non-polynomial link functions, we note that similar to [DPVLB24], the existence of polynomial basis is needed to exclude extreme cases, and we cannot upper bound the required degree $I$ because general link functions are not included in a compact space.
+Proof of Lemma 8.
+The derivation is analogous to [DPVLB24, Lemma F.14].
+Let $z\sim\mathcal{N}(0,1)$ and $y=\sigma_{*}(z)$.
+We define $\zeta_{p_{*}}(y)=\mathbb{E}[\frac{1}{\sqrt{p_{*}!}}\mathsf{He}_{p_{*}}(z)|y]$ and its basis expansion $\zeta_{p_{*}}(y)=\sum_{k=0}^{\infty}v_{k}\phi_{k}$.
+Let $K$ be a smallest integer such that $v_{k}\neq 0$.
+Then, there exists an integer with $I\leq K$ such that $\mathrm{IE}(y^{I})=p_{*}$.
+Indeed,
+$\mathbb{E}[\phi_{K}(y)\mathsf{He}_{p_{*}}(z)]$ $=\mathbb{E}_{y}[\Phi_{K}(y)\mathbb{E}_{z|y}[\mathsf{He}_{p_{*}}(z)|y]]$ (A.26)
+$=\mathbb{E}_{y}\bigg{[}\Phi_{K}(y)\sum_{k=0}^{K}v_{k}\phi_{k}(y)\bigg{]}=v_{K}\neq 0,$ (A.27)
+which means that at least one of $y,y^{2},\cdots,y^{K}$ yields a non-zero $p_{*}$-th Hermite coefficient. ∎
 Appendix B SGD with Reused Batch
-In this section we establish the statistical and computational complexity of Algorithm 1.
-Recall that the algorithm first trains the first-layer parameters with $T_{1}$ steps of SGD update, where we reuse the same sample for two consecutive steps.
-The analysis of first-layer training is divided into two phases: $(i)$ weak recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\geq\varepsilon$), and $(ii)$ strong recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\geq 1-\varepsilon$).
-We then train the second-layer parameters after strong recovery is achieved.
-The section is organized as follows.
-  • Section B.1 verifies the conditions on the activation function $\sigma$ to guarantee weak and strong recovery.
-  • Section B.2 isolates a (nearly) constant fraction of neurons at initialization with an alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ above a certain threshold.
-We focus on such neurons in Phase I of first-layer training.
-  • Section B.3 lower bounds the expected update of alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ of two gradient steps, and Section B.4 shows that neurons yield weak recovery within $2T_{1,1}=\tilde{O}(d)$ steps.
-  • Section B.5 discusses how to convert weak recovery to strong recovery using $2T_{1,2}=\tilde{O}(d\varepsilon^{-2})$ SGD steps.
-  • Finally, Section B.6 analyzes the second-layer training and concludes the proof.
-In the following proofs, we introduce constants $c_{i}$ and $C_{i}$, which depends on $d$ at most polylogarithmically. Specifically, the asymptotic strength of the constants is ordered as follows.
-$1\simeq C_{1}\lesssim c_{1}^{-1}\lesssim C_{2}\lesssim C_{3}\lesssim c_{2}^{-1}$ (B.1)
-$c_{\eta}$ in the main text can be taken as $c_{\eta}=c_{1}$, where $c_{1}$ should satisfy $\lim_{d\to\infty}c_{1}=0$, but the convergence can be arbitrarily slow.
-This requirement comes from the fact that we do not know the exact value of $H(\sigma_{*}^{I};k_{*})$, which might be very small.
-To ensure that the signal is isolated, taking $\eta\asymp c_{1}d^{-1}$ with arbitrarily slow $c_{1}$ suffices.
-$C_{2}$ can also be arbitrarily slow, as long as it satisfies $C_{2}=\mathrm{poly}(c_{1}^{-1})$.
-$C_{3}=\mathrm{polylog}(d)$ will be used to represent polylogarithmic factor that comes from high probability bounds.
-B.1 Conditions on the Activation Function
-B.1.1 Verifying Assumption 2
-In the following, we focus on the activation function of a single neuron and omit the subscript that distinguishes different neurons. Recall that we consider polynomial activation functions written as
-$\sigma(z):=\sum_{j=0}^{C_{\sigma}}\beta_{j}\mathsf{He}_{j}(z).$ (B.2)
-For weak recovery, we can use any polynomial that has degree $C_{\sigma}\geq C_{q}$ as long as the following condition holds: If $\mathrm{IE}(\sigma_{*})\geq 2$ and there exists some $i\leq C_{\sigma}$ such that $\mathrm{IE}(\sigma_{*}^{i})=1$, $\sigma$ should satisfy
-$H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}\neq 0.$ (B.3)
-If $\mathrm{IE}(\sigma_{*})\geq 3$ and there does not exist any $i\leq C_{\sigma}$ such that $\mathrm{IE}(\sigma_{*}^{i})=1$ (in this case there exists some $i\leq C_{q}$ such that $\mathrm{IE}(\sigma_{*}^{i})=2$), $\sigma$ should satisfy
-$H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};1\bigg{)}\neq 0.$ (B.4)
-Below we prove Lemma 2 which shows that the above conditions are met with probability $1$ for randomly drawn the Hermite coefficients.
+In this section we show that Algorithm 1 learns single-index models in $\tilde{O}(d^{1\lor(p_{*}-1)})$ samples with high probability.
+The algorithm trains the first layer for $T_{1}$ SGD steps, where we sample a new data point in every two steps.
+The first layer training is further divided into two phases: weak recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\gtrsim 1$) and strong recovery ($\|\boldsymbol{w}-\boldsymbol{\theta}\|\lesssim\varepsilon$).
+Then, we learn the second layer parameters.
+Specifically, Section B.2 shows that at initialization, a (nearly) constant fraction of neurons has alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ beyond a certain threshold.
+We focus on such neurons in the first phase of training.
+Section B.3 lower bounds the expected update of alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ of two gradient steps, and Section B.4 establishes that the neurons achieve weak recovery within $2T_{1,1}=\tilde{O}(d^{1\lor(p_{*}-1)})$ steps.
+Section B.5 discusses how to convert weak recovery to strong recovery using $2T_{1,2}=\tilde{O}(d\varepsilon^{-2})$ more steps.
+We let $T_{1}=2T_{1,1}+2T_{1,2}$.
+Finally, Section B.6 analyzes second layer training and concludes the proof.
+In the following proofs, we use several constants, which depends on $d$ at most at most polylogarithmically. Specifically, asymptotic strength of the constants is ordered as follows.
+$1\simeq c_{\sigma}\simeq C_{1}\lesssim\mathopen{}\mathclose{{}\left\{\begin{matrix}c_{\eta}^{-1}\simeq C_{2}\lesssim\mathrm{poly}(c_{\eta}^{-1})\lesssim\mathopen{}\mathclose{{}\left\{\begin{matrix}c_{1}^{-1}\simeq C_{3}\\ c_{2}^{-1}\\ \end{matrix}}\right\}\\ \delta^{-1}\\ \end{matrix}}\right\}$$\lesssim\mathopen{}\mathclose{{}\left\{\begin{matrix}\delta^{-1}\mathrm{poly}(c_{\eta}^{-1})\lesssim c_{\xi}^{-1}\\ \mathrm{poly}(c_{1}^{-1})\lesssim\bar{c}_{\eta}^{-1}\end{matrix}}\right\}$ (B.1)
+$\lesssim\mathrm{polylog}(d)=C_{4}.$ (B.2)
+Here, $c_{\eta}$ and $\delta$ should satisfy $\lim_{d\to\infty}c_{\eta}=\lim_{d\to\infty}\delta=0$, but the convergence can be arbitrarily slow, (e.g., as slow as $1/\log\log\log\cdots\log d$).
+This requirement comes from the fact that we do not know the exact value of $H(\sigma_{*}^{I};p_{*})$.
+To ensure that one signal term (from the Taylor series) is isolated, taking $\eta\asymp d^{-1}$ with a sufficiently small constant is insufficient but $\eta\asymp c_{\eta}d^{-1}$ with arbitrarily slow $c_{\eta}$ suffices.
+Also, to guarantee that the failure probability is $o_{d}(1)$, we require $\delta$ to be $o_{d}(1)$.
+$c_{\xi}$ can also decay arbitrarily slowly, as long as it satisfies $c_{\xi}\lesssim\delta\mathrm{poly}(c_{\eta}^{-1})$.
+$C_{4}=\mathrm{polylog}(d)$ will be used to represent any polylogarithmic factor that comes from high probability bounds.
+For the first-layer training, we can reduce the argument into training of one neuron using the correlation loss as follows. At each step, the gradient update (Line 8 of Algorithm 1) is written as
+$\boldsymbol{w}^{t+1}_{j}$ $\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\big{(}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2}\big{)}$ (B.3)
+$=\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}^{2}+2\eta^{t}_{j}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}$ (B.4)
+$=\boldsymbol{w}^{t}_{j}-\frac{2\eta^{t}c_{a}^{2}}{N}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}\big{(}\tilde{\nabla}_{\boldsymbol{w}}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}+\frac{2\eta^{t}c_{a}}{N}y\big{(}\tilde{\nabla}_{\boldsymbol{w}}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}.$ (B.5)
+While the second term scales with $\eta^{t}c_{a}^{2}N^{-1}$, the third term scales with $\eta^{t}c_{a}N^{-1}$.
+Thus, by setting $c_{a}$ sufficiently small, we can ignore the interaction between neurons.
+We will show that the strength of the signal in the direction of $\boldsymbol{\theta}$ is at least $(\kappa_{j}^{t})^{p_{*}-1}\gtrsim d^{-\frac{p_{*}-1}{2}}$ (up to a polylogarithmic factor, and $p_{*}=\mathrm{GE}(\sigma_{*})$).
+On the other hand, we can easily see that $\boldsymbol{\theta}^{\top}\big{(}\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}\big{(}\tilde{\nabla}_{\boldsymbol{w}}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}$ is bounded by $\tilde{O}(1)$ with high probability.
+Therefore, by simply letting $c_{a}=\tilde{\Theta}(d^{-\frac{p_{*}-1}{2}})$, we can ignore the effect of the second term in (B.5).
+Moreover, for simplicity, we will reparameterize $\frac{2\eta^{t}c_{a}}{N}$ as $\eta^{t}$ below.
+Consequently, we may analyze the following update
+$\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}+\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\big{(}y\sigma_{j}({\boldsymbol{w}^{t}_{j}}^{\top}\boldsymbol{x})\big{)},$ (B.6)
+instead of Line 8 of Algorithm 1. Since there is no interaction between neurons now, we omit the subscript $j$ when the context is clear.
+B.1 Assumptions on Link Function
+The analysis consists of three different phases: weak recovery and strong recovery of the first-layer weights, and approximation of the link function (ridge regression of the second-layer).
+Each phase requires different assumptions on the activation functions, depending on the link function.
+Before starting the analysis, we decompose Assumptions 2 and 3 and clarify which conditions are needed in each phase.
+We prove that instead of using a specific activation function tailored to different link functions, a randomized activation function satisfies all required assumptions with probability $\Omega(1)$.
+In the following, we write the student activation function as
+$\sigma_{j}(s):=\sum_{i=0}^{\infty}\beta_{j,i}\mathsf{He}_{i}(s)$ (B.7)
+with coefficients $\{\beta_{j,i}\}_{i=0}^{C_{\sigma}}$ (sometimes the subscript $j$, which is the index of the neurons, is omitted).
+B.1.1 For polynomial link functions
+In the following, we summarize the precise conditions to be satisfied by the activation functions (these conditions are weaker than Assumptions 2 and 3).
+For polynomial link functions, we focus on polynomial activation functions (with bounded degree) for simplicity, but non-polynomial activation functions would not change the proof significantly.
+Let $p$ and $q$ be the minimum and maximum degree of non-zero Hermite coefficients of $\sigma_{*}$.
+Note that $\mathrm{GE}(\sigma_{*})=1$ or $2$ holds (see Proposition 6).
+Let $I\leq C_{q}$ (according to Proposition 6) be the smallest integer such that $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}$ and $C_{\sigma}$ be the degree of the activation function.
+  • (I)
+    If $I=1\Leftrightarrow\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})=p_{*}$.
+    • Weak recovery: $\alpha_{p_{*}}\beta_{p_{*}}>0$ (covered by Assumption 3).
+    • Strong recovery: $\sum_{j=p_{*}}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ (covered by Assumption 2).
+    • Approximation (ridge regression): $\beta_{i}\neq 0$ for some $i\geq q$ (covered by Assumption 3).
+  • (II)
+    If $2\leq I=\{\min\ i\mid\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}\}\leq C_{\sigma}$.
+    • Weak recovery: $H((\sigma_{*})^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)>0$ (covered by Assumption 3).
+    • Strong recovery: $\sum_{j=p_{*}+1}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ (covered by Assumption 2).
+    • Approximation: $\beta_{i}\neq 0$ for some $i\geq q$ (covered by Assumption 3).
+Note that it is difficult to construct a deterministic activation function that satisfies all of the assumptions for any link function $\sigma_{*}$ (the simplest counterexample is to consider $-\sigma_{*}$ which flips the Hermite coefficients).
+Instead, we show the existence of randomized construction of such an activation function that satisfies all of the assumptions on the activation function simultaneously with constant probability, which entails that a subset of neurons can achieve strong recovery.
+The construction does not depend on properties of the link function itself except for its degree $q$.
+Lemma 11.
+There exists a randomized activation function sampled from a discrete set such that the above conditions hold with constant probability.
+Proof.
+Let $c$ be a sufficiently small constant only used in this proof and $C_{\sigma}$ be the minimum odd integer with $C_{\sigma}\geq\max\{C_{q}+1,q+2,3\}$, where $C_{q}$ was introduced in Proposition 6.
+With probability $\frac{1}{2}$, we let $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$, and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$.
+With probability $\frac{1}{2}$, we let $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}\sim\mathrm{Unif}(\{\pm 1\})$.
+We first consider (I).
+When $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$, and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$, it is easy to see $\mathrm{sign}(\alpha_{j})=\mathrm{sign}(\beta_{j})$ for all $j=1,\cdots,q$ hold with probability at least $2^{-q}$, which is sufficient to satisfy (I).
+We then consider (II).
+First focus on the case when $p_{*}=1$ and $I$ is even.
+When $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$ and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$, by taking $c$ sufficiently small, we have
+$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)=\underbrace{I!\beta_{I}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}).$ (B.8)
+When $I$ is even, by adjusting the sign of $\beta_{1}$, $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is non-zero and has the same sign as $H((\sigma_{*})^{I};1)$ with probability $\frac{1}{2}$.
+Note that the sign of $\beta_{1}$ is independent from whether $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds.
+This holds with probability at least $2^{-q+1}$.
+Thus we verified (II) for $p_{*}=1$ and even $I$.
+For $p_{*}=1$ and odd $I$, consider $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}\sim\mathrm{Unif}(\{\pm 1\})$.
+Note that $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ (this is the condition for strong recovery) and the condition for ridge regression also holds.
+Furthermore, the sign of $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};0)$ is ±$\pm 1$ with equiprobability, independent of $\beta_{2},\dots,\beta_{q}$.
+Therefore, by taking $c$ sufficiently small, we can obtain the desired sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$.
+Thus we proved (II) for $p_{*}=1$ and odd $I$.
+Regarding (II) for $p_{*}=2$ and even $I$, when $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$ and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$, we have
+$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)=\underbrace{(I+1)!\beta_{I+1}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}).$ (B.9)
+Thus, similar to (II) with $p_{*}=1$ and even $I$, we get (II) for $p_{*}=2$ and even $I$.
+Finally, consider (II) for $p_{*}=2$ and odd $I$.
+When $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}\sim\mathrm{Unif}(\{\pm 1\})$, the sign of $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};1)$ is ±$\pm 1$ with equiprobability when $I$ is odd, and this term dominates the others in $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)$.
+Thus, (II) for $p_{*}=2$ and odd $I$ holds similarly to (II) for $p_{*}=1$ and odd $I$.
+Now we have obtained the assertion for all cases. ∎
+B.1.2 For general link functions
+Now we consider non-polynomial link functions with potentially large generative exponent $p_{*}=\mathrm{GE}(\sigma_{*})\geq 2$.
+For weak and strong recovery to succeed, the conditions on the activation function are essentially the same as those for polynomial link functions:
+  • (I)
+    If $I=1\Leftrightarrow\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})=p_{*}$.
+    • Weak recovery: $\alpha_{p_{*}}\beta_{p_{*}}>0$.
+    • Strong recovery: $\sum_{j=p_{*}}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$,
+  • (II)
+    If $2\leq I=\{\min\ i\mid\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}\}\leq C_{\sigma}$.
+    • Weak recovery: $H((\sigma_{*})^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)>0$,
+    • Strong recovery: $\sum_{j=p_{*}+1}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
+Due to the proof strategy (which uses Taylor expansion), we also require that all differentials and sum of expectations appearing in the following proofs are well-defined and bounded.
+To approximate a non-polynomial $\sigma_{*}$, we introduce the following condition on the activation function.
+We sample $\sigma_{j}$ from a discrete set (with bounded cardinality).
+Let $J$ be an index set such that the coefficients of $\sigma_{j}\ (j\in J)$ satisfy the conditions above.
+Because we are selecting $\sigma_{j}$ from a discrete set, $|J|\simeq N$ holds.
+We introduce the following condition, which states that the target single-index model can be well-approximated by a linear combination of student neurons.
+Assumption 4.
+When $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$ where $(C_{b}=\mathrm{polylog}(d))$ and $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T_{2}}\sim\mathcal{N}(0,\boldsymbol{I}_{d})$, there exists a set of coefficients $a_{1},\dots,a_{|J|}$ such that
+$\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\bigg{(}\frac{1}{|J|}\sum_{j\in J}a_{j}\sigma_{j}(\boldsymbol{\theta}_{j}^{\top}\boldsymbol{x}_{i}+b_{j})-\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}_{i})\bigg{)}^{2}\lesssim\varepsilon^{2},$ (B.10)
+holds with coefficients of reasonable magnitudes $\sum_{j\in J}a_{j}^{2}=\Theta(|J|)$ with high probability (w.r.t.
+the randomness of $b_{j}$ and $\boldsymbol{x}_{i}$).
+Moreover, $\mathbb{E}_{\boldsymbol{x}}[\sigma_{j}(\boldsymbol{\theta}_{j}^{\top}\boldsymbol{x}+b_{j})^{4}]\leq\mathrm{polylog}(d)$ for all $j$ with high probability (w.r.t.
+the randomness of $b_{j}$).
+The following lemma states that we can design a randomized activation function that satisfies all of the above assumptions with probability $\Omega(1)$, as long as the link function $\sigma$ satisfies Assumption 4 for $\sigma=\mathrm{ReLU}$.
+In other words, we are able to cover the class of link functions $\sigma$ that can be efficiently approximated by a two-layer ReLU network.
+Since the general link functions are not included in a compact space, we do not have an upper bound of exponent to obtain $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})$ as we had $C_{q}$ in the polynomial case.
+Consequently, our student activation is not entirely agnostic to the link function $\sigma_{*}$, as we require knowledge of $p$ (information exponent), $p_{*}$ and $I$.
+Lemma 12.
+Suppose the target link function $\sigma_{*}$ satisfies Assumption 4 for $\sigma_{j}=\mathrm{ReLU}$.
+There exists a randomized activation sampled from a discrete set such that the above conditions hold with constant probability.
+Before we sketch the design of activation function, we present the following approximation result from [BBSS22], which establishes that Assumption 4 with $\sigma_{j}=\mathrm{ReLU}$ is satisfied for broad class of functions, according to Lemma 4.4 and 4.5 of [BBSS22].
+Specifically, taking $\tau=1/2$ and $\lambda=N^{-1}$ yields that $\mathbb{E}_{\boldsymbol{x}}[(\frac{1}{|J|}\sum_{j\in J}a_{j}\sigma_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})-\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{2}]\leq N^{-\frac{2}{7}}$.
+Although they sample $b_{j}$ from Gaussian $\mathcal{N}(0,2)$, the result translates to uniform sampling of biases from $[-C_{b},C_{b}]$ by introducing additional logarithmic factor.
+Lemma 13 (Lemma 4.4, 4.5 of [BBSS22]).
+Suppose that $\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma_{*}(z)^{4}]$, $\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma_{*}^{\prime\prime}(z)^{4}]<\infty$.
+Then, Assumption 4 with $\sigma_{j}=\mathrm{ReLU}$ holds with $\varepsilon=N^{-\frac{1}{7}}$ and $C_{b}\simeq\sqrt{\log d}$.
+Proof of Lemma 12.
+We show the existence of suitable $\sigma$ in two steps: first we construct a randomized polynomial activation function that satisfies conditions (I)(II) with constant probability; then we add a small ReLU perturbation so that the activation can approximate non-polynomial $\sigma_{*}$.
+Recall $p\in\mathbb{N}_{+}$ is the information exponent of $\sigma_{*}$.
+We first show that there exists a randomized polynomial activation that satisfies the conditions for weak and strong recovery with probability $\Omega(1)$.
+Note that the issue of differentiability and bounded moment is avoided when we focus on the polynomial activation functions.
+We specify the following two distributions.
+With probability $\frac{1}{2}$, let $\beta_{1}\sim\mathrm{Unif}(\{-1,1\})$, $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $j=1,\cdots,p_{*}+I-1$ and $\beta_{j}=0$ otherwise, where $c>0$ is a sufficiently small constant.
+With probability $\frac{1}{2}$, let $\beta_{1}=\mathrm{Unif}(\{-1,1\})$, $\beta_{2}=\mathrm{Unif}(\{-c,c\})$, $\beta_{j}=\mathrm{Unif}(\{-c^{2},c^{2}\})$ for all $2\leq j\leq(p_{*}+I)\lor p$ for a sufficiently small constant $c>0$, and $\beta_{j}=0$ otherwise.
+Regarding (I), consider the case when the coefficients are sampled from the first distribution, and $|\beta_{j}|\ll 1$ except for $j=p_{*}$.
+Then, $\sum_{j=p_{*}}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}\approx p_{*}!\alpha_{p_{*}}\beta_{p_{*}}s^{p_{*}}$.
+Choosing the sign of $\beta_{p_{*}}$, we have that the assumption holds with probability $\Omega(1)$.
+Regarding (II) with even $I$, consider coefficients sampled from the first distribution, and $\mathrm{Sign}(\beta_{j})=\mathrm{Sign}(\alpha_{j})$ for $j\leq(p_{*}+I-1)\lor p$.
+Then, $\sum_{j=p_{*}+1}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
+Also, similarly to (B.8),
+$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)=\underbrace{i!\beta_{I+p_{*}-1}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}).$ (B.11)
+By flipping the sign of $\beta_{1}$, we can change the sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$. Thus, (II) for even $I$ is satisfied by a randomized choice of $\beta_{1}$.
+For (II) with odd $I$, consider coefficients sampled from the second distribution, and $\mathrm{Sign}(\beta_{j})=\mathrm{Sign}(\alpha_{j})$ for $j\leq(p_{*}+I)\lor p$.
+Then, $\sum_{j=p_{*}+1}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
+$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)=\frac{1}{(p_{*}-1)!}\mathbb{E}[\sigma^{(I)}(\sigma^{(1)})^{I-1}\mathsf{He}_{p_{*}-1}]$ (B.12)
+$=\frac{1}{(p_{*}-1)!}\mathbb{E}[(I-1)(\beta_{p_{*}+I}\mathsf{He}_{p_{*}+I})^{(I)}(\beta_{2}\mathsf{He}_{2})^{(1)}(\beta_{1})^{I-2}\mathsf{He}_{p_{*}-1}]+O(c^{4}).$ (B.13)
+$=\underbrace{\frac{2(I-1)\beta_{p_{*}+I}\beta_{2}(\beta_{1})^{I-2}(p_{*}+I)!}{(p_{*}-1)!}}_{\text{$\asymp c^{3}$}}+O(c^{4}).$ (B.14)
+By flipping the sign of $\beta_{1}$, we can change the sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$. Thus, (II) for odd $I$ is satisfied by a randomized choice of $\beta_{1}$.
+Therefore, we have constructed a randomized polynomial activation $\sigma$ that satisfies all of the conditions for weak and strong recovery.
+Now we provide a sketch of reasoning that when the link function $\sigma_{*}$ is well-approximated by ReLU as specified in Assumption 4, we can find some $\sigma$ that additionally satisfies Assumption 4 by introducing a small ReLU component.
+Specifically, we add $c_{\mathrm{R}}\cdot\mathrm{ReLU}$ to the activation function with probability $\frac{1}{2}$, with a sufficiently small $c_{\mathrm{R}}=\tilde{\Omega}(1)$, e.g., $c_{\mathrm{R}}=(\text{log}d)^{-C}$ for some $C>0$.
+When a two-layer ReLU network approximates $\sigma_{*}$ that satisfies Assumption 4, by using the neurons with added ReLU component, $\sigma_{*}$ can be approximated up to some polynomial residual with degree $(p_{*}+I)\lor p$.
+And by using the remaining polynomial neurons, we can approximate the additional polynomial terms in $\sigma_{*}$ (see Lemma 22,23).
+Subtracting the latter from the former, we obtain the desired approximation result.
+When $c_{\mathrm{R}}$ is sufficiently small, this additional term does not impact the conditions for weak and strong recovery and the moment calculations; similarly, since $c_{\mathrm{R}}\ll 1$ we may discard this non-smooth term before Taylor expansion without affecting the analysis of optimization dynamics.
+We remark that to avoid such unnatural design of activation function, we can also train the first-layer parameters using a polynomial activation specified above, and then perturb it before the second-layer training to enhance the approximation ability — such strategy has also been employed in prior layer-wise training analysis [AAM22].
+∎
+B.1.3 More Discussion on Assumption 2
+Assumption 2 requires $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$ is not zero and has the same sign as $H(\sigma_{*}^{I};p_{*})$.
+We remark that if we allow a negative momentum parameter larger than $1$, i.e., setting $\xi^{2(t+1)}=1+c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$, we can negate the opposite sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$ (see Lemma 16), and the subsequent analysis still holds.
+Therefore, what we essentially need is $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)\neq 0$.
+Lemma 3 confirms that it is satisfied by almost all polynomials:
 Proof of Lemma 3.
@@
 If $k\leq i(C_{\sigma}-2)+1$, this choice of $\sigma$ yields $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)>0$, which confirms that $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)$ as a polynomial of $\{\beta_{j}\}_{j=0}^{C_{\sigma}}$ is not identically equal to zero.
-Now, the assertion follows from the Schwartz–Zippel Lemma [Sch80], or the fact that zeros of a non-zero polynomial form a measure-zero set. ∎
-B.1.2 Verifying Assumption 3
-On the other hand, for the strong recovery we require an additional condition on the activation function due to link misspecification, which is also introduced in [BAGJ21, MHWSE23]:
-$\sum_{j=p}^{p}j!\alpha_{j}\beta_{j}s^{j}>0\quad\text{for all $s>0$}.$ (B.5)
-In order to meet Assumption 2 and (3) simultaneously, we follow [OSSW24] and randomize the activation function. Specifically, the activation function should satisfy
-  • (I)
-    If $\mathrm{IE}(\sigma_{*})=1$. We require $\beta_{1}>0$ and $\sum_{j=1}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
-  • (II)
-    Else if $\mathrm{IE}((\sigma_{*})^{I})=1$ for some $2\leq I\leq C_{\sigma}$.
-We require $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};1)$.
-Also, $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
-  • (III)
-    Else if $\mathrm{IE}(\sigma_{*})=2$. We require $\beta_{2}>0$ and $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
-  • (IV)
-    Else $\mathrm{IE}((\sigma_{*})^{I})=2$ for some $2\leq I\leq C_{q}$.
-We require $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};2)$.
-Also, $\sum_{j=3}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
-Now we prove Lemma 3 which verifies the existence of an activation function that satisfies the assumptions above with non-zero probability.
-The construction does not depend on the link function itself, but only its degree $q$.
-Proof of Lemma 3.
-Let $c$ be a sufficiently small constant, and $C_{\sigma}$ be the minimum odd integer with $C_{\sigma}\geq\max\{C_{q}+1,q+2,3\}$.
-With probability $1/2$, we choose the coefficients as $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$, and $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $2\leq j\leq C_{\sigma}$.
-Then, it is easy to see (I) and (III) are met with probability at least $2^{-q}$, because they hold when $\mathrm{sign}(\alpha_{j})=\mathrm{sign}(\beta_{j})$ holds in the summation.
-By taking $c$ sufficiently small, we have
-$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)=\underbrace{i!\beta_{I}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c).$ (B.6)
-When $I$ is even, by adjusting the sign of $\beta_{1}$, $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};1)$ with probability $\frac{1}{2}$.
-Note that the sign of $\beta_{1}$ is independent from whether $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds.
-This holds with probability at least $2^{-q+1}$.
-Thus we verified (II) for even $I$.
-In the same vein, we can verify (IV) for even $I\leq C_{q}\leq C_{\sigma}-1$. We have
-$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)=\underbrace{(I+1)!\beta_{I+1}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}),$ (B.7)
-and (IV) can be verified using the same argument.
-Otherwise (also with probability $1/2$), we choose the coefficients as $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}=\pm 1$ to verify (II) and (IV) for odd $I$.
-It is easy to see that $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds for (I) and $\sum_{j=3}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ for (III).
-In addition, $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};0)>0$ and $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};1)>0$.
-Therefore, by taking $c$ sufficiently small, flipping the sign of $\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1}$ can change the sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};k)$ for both (II) and (IV) with odd $I$.
-Combining all cases yields the desired claim.
+Hence the assertion follows from so-called Schwartz–Zippel Lemma [Sch80], or the fact that zeros of a non-zero polynomial form a measure-zero set.
 ∎
-B.2 Random Initialization
-In Section B.3.1 we focus on the neurons with slightly larger initial alignment that satisfy $\kappa^{0}_{j}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}$ at initialization, where constant $C_{2}$ grows at most polylogarithmically in $d$.
-The following lemma states that roughly a constant portion of the neurons satisfies this initial alignment condition.
+B.2 Initialization
+We first consider the initial alignment.
+In the following sections, we focus on the neurons that satisfy $\kappa^{0}_{j}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{0}_{j}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$ at the initialization.
+The following lemma states that roughly a constant portion of the neurons satisfy the initial alignment condition upon random initialization.
+In particular, if we take $c_{\eta}=\Omega((\log\log d)^{-\frac{1}{2}})$, the fraction of neurons that satisfy the initial alignment condition is at least $e^{-16c_{\eta}^{-2}}=\tilde{\Omega}(1)$.
+Let us write $C_{2}=c_{\eta}^{-1}$ for simplicity in the following.
 Lemma 14.
@@
 $\frac{\sqrt{2e(\beta-1)}}{2\beta\sqrt{\pi}}e^{-\frac{\beta s^{2}}{2}}\leq\int_{s}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{t^{2}}{2}}\mathrm{d}t$ (B.16)
-Proof of Lemma 7. Because $\kappa^{0}=v^{\top}w\overset{\mathrm{d}}{=}\frac{{\sf e}_{1}^{\top}g}{g}$, where $g\sim\mathcal{N}(0,\boldsymbol{I}_{d})$,
-$\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$ $=\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\sf e}_{1}^{\top}g\geq 4C_{2}\land\|g\|\leq 2d^{\frac{1}{2}}\bigg{]}$ (B.10)
-$\geq\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\sf e}_{1}^{\top}g\geq 4C_{2}\bigg{]}-\mathbb{P}_{g\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\|g\|\geq 2d^{\frac{1}{2}}\bigg{]}$ (B.11)
-$\gtrsim\frac{\sqrt{2e(\beta-1)}}{2\beta\sqrt{\pi}}e^{-8\beta C_{2}^{2}}-e^{-\Omega(d)}.$ (B.12)
-By letting $\beta=2$, we have that $\mathbb{P}[\kappa^{0}_{j}\geq C_{2}d^{-\frac{1}{2}}]\gtrsim e^{-16C_{2}^{2}}.$ Due to symmetry, $\mathbb{P}[\kappa^{0}_{j}\leq 2C_{2}d^{-\frac{1}{2}}]=\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$.
+Proof of Lemma 14.
+Because $\kappa^{0}=\boldsymbol{v}^{\top}\boldsymbol{w}\overset{\mathrm{d}}{=}\frac{{\boldsymbol{e}}_{1}^{\top}\boldsymbol{g}}{\|\boldsymbol{g}\|}$, where $\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})$,
+$\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$ $=\mathbb{P}_{\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\boldsymbol{e}}_{1}^{\top}\boldsymbol{g}\geq 4C_{2}\land\|\boldsymbol{g}\|\leq 2d^{\frac{1}{2}}\bigg{]}$ (B.17)
+$\geq\mathbb{P}_{\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}{\boldsymbol{e}}_{1}^{\top}\boldsymbol{g}\geq 4C_{2}\bigg{]}-\mathbb{P}_{\boldsymbol{g}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\|\boldsymbol{g}\|\geq 2d^{\frac{1}{2}}\bigg{]}$ (B.18)
+$\gtrsim\frac{\sqrt{2e(\beta-1)}}{2\beta\sqrt{\pi}}e^{-8\beta C_{2}^{2}}-e^{-\Omega(d)},$ (B.19)
+where we used Lemma 15 for the final inequality.
+By letting $\beta=2$, we have that $\mathbb{P}[\kappa^{0}_{j}\geq C_{2}d^{-\frac{1}{2}}]\gtrsim e^{-16C_{2}^{2}}.$ Because of the symmetry, $\mathbb{P}[\kappa^{0}_{j}\leq 2C_{2}d^{-\frac{1}{2}}]=\mathbb{P}[\kappa^{0}_{j}\geq 2C_{2}d^{-\frac{1}{2}}]$.
 ∎
-B.3 Population Update
-We first analyze the training of first-layer parameters by evaluating the expected (population) update of two gradient steps with the same training example. At each step, the parameters are updated as
-$\boldsymbol{w}^{t+1}_{j}$ $\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\big{(}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2}\big{)}$ (B.13)
-$=\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}^{2}+2\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}.$ (B.14)
-While the second term scales with $\eta^{t}a_{j}^{2}=\eta^{t}c_{a}^{2}$, the third term scales with $\eta^{t}a_{j}=\eta^{t}c_{a}$.
-Thus, by setting the second-layer scale $c_{a}$ sufficiently small, we can ignore the interaction of neurons; similar mechanism also appeared in [BES+22, AAM23].
-Specifically, in the following, we show that the strength of the signal is $\boldsymbol{\theta}^{\top}\boldsymbol{w}_{j}^{t}\gtrsim d^{-\frac{1}{2}}$.
-Thus, by simply letting $c_{a}\lesssim C_{3}^{-1}d^{-\frac{1}{2}}$, we can ignore the effect of the squared term.
-Thus we may focus on the following correlational update:
-$\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}+\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\Bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\Bigg{)}.$ (B.15)
-Due to the absence of interaction between neurons, we omit the subscript $j$ for the index of neurons and ignore the prefactor of $N$ (which can be absorbed into the learning rate); multiplying $N$ to $\eta^{t}$ specified below recovers the scaling of $\eta^{t}$ presented in the main text.
-B.3.1 When $\mathrm{IE}(\sigma_{*}^{I})=2$ with $I\geq 2$
-First we consider the most technically difficult case, when $\mathrm{IE}[\sigma_{*}]\geq 3$ and the information exponent cannot be lowered to 1 for $i\leq C_{\sigma}$; in this case, from Proposition 5 we know that there exists some $2\leq i\leq C_{\sigma}$ such that $\mathrm{IE}[\sigma_{*}^{i}]=2$ and we let $I$ be the first such $i$.
+B.3 Weak Recovery: Population Update
+We divide the first layer training into the first phase (weak recovery) and the second phase (strong recovery).
+We first evaluate the expected update of two gradient steps with the same training example.
+Lemma 16.
+Let $\eta^{2t},\eta^{2t+1}=\eta=c_{\eta}d^{-1}$, $\xi^{2(t+1)}=\xi=1-c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$.
+Suppose that the link function satisfies $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}$ (we choose the smallest such $I$) and activation functions satisfy all of the assumptions in Section B.1 for weak recovery.
+Then, for $\boldsymbol{w}^{2t}$ with $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}\leq c_{\eta}^{I}$, the alignment $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}$ can be evaluated as,
+$\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}\geq\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}+c_{\eta}^{I}c_{\xi}c_{\sigma}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\nu^{2t}.$ (B.20)
+Here $c_{\sigma}=p_{*}!\alpha_{p_{*}}\beta_{p_{*}}$ (when $\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})$) or $c_{\sigma}=\frac{p_{*}!H(\sigma_{*}^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)}{2(I-1)!}$ (otherwise), and $\nu^{2t}$ is a mean-zero sub-exponential random variable.
+Proof.
+The expected alignment $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}$ after two gradient steps from $\boldsymbol{w}^{2t}=\boldsymbol{\omega}$ using the same sample $(\boldsymbol{x},y)$, step size $\eta^{2t}=\eta^{2t+1}=\eta=c_{\eta}d^{-1}$ and momentum parameter $\xi^{2(t+1)}=\xi=1-c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$ is evaluated as follows.
+With a projection matrix $\boldsymbol{P}_{\boldsymbol{\omega}}=\boldsymbol{I}-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}$, the first step updates the weight as
+$\boldsymbol{w}^{2t+1}\leftarrow\boldsymbol{w}^{2t}+\eta\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x},$ (B.21)
+and the next gradient step with the same sample is computed as
+~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t+1}}^{\top}\boldsymbol{x})$ $=y\sigma^{\prime}({\boldsymbol{w}^{2t+1}}^{\top}\boldsymbol{x})\boldsymbol{x}$ (B.22)
+$=y\sigma^{\prime}\big{(}(\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})^{\top}\boldsymbol{x}\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}$ (B.23)
+$=y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x},$ (B.24)
+here we used the notation $\|\boldsymbol{\theta}\|_{A}^{2}=\boldsymbol{\theta}^{\top}A\boldsymbol{\theta}$ for a vector $\boldsymbol{\theta}\in\mathbb{R}^{d}$ and a positive symmetric matric $A\in\mathbb{R}^{d\times d}$.
+From (B.21) and (B.24), the parameter after the two steps is obtained as
+$\boldsymbol{w}^{2(t+1)}$ $\leftarrow\boldsymbol{w}^{2t+1}+\eta\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t+1}}^{\top}\boldsymbol{x})$ (B.25)
+$=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+\eta y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.26)
+$=\boldsymbol{\omega}+\eta\boldsymbol{g}^{2t},$ (B.27)
+where
+$\boldsymbol{g}^{2t}=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.28)
+Finally, the normalization step yields
+$\boldsymbol{w}^{2(t+1)}$ $\leftarrow\frac{\boldsymbol{w}^{2(t+1)}-\xi^{2(t+1)}(\boldsymbol{w}^{2(t+1)}-\boldsymbol{w}^{2t})}{\|\boldsymbol{w}^{2(t+1)}-\xi^{2(t+1)}(\boldsymbol{w}^{2(t+1)}-\boldsymbol{w}^{2t})\|}=\frac{\boldsymbol{\omega}+\eta\xi\boldsymbol{g}^{2t}}{\|\boldsymbol{\omega}+\eta\xi\boldsymbol{g}^{2t}\|}=\frac{\boldsymbol{\omega}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{g}^{2t}}{\|\boldsymbol{\omega}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{g}^{2t}\|}.$ (B.29)
+Therefore, by writing $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}=\kappa^{2t}$, the update of the alignment is
+$\kappa^{2(t+1)}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}$ (B.30)
+$=\frac{\kappa^{2t}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}}{\|\boldsymbol{\omega}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{g}^{2t}\|}$ (B.31)
+$\geq\kappa^{2t}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}-\frac{1}{2}\kappa^{2t}c_{\eta}^{2}c_{\xi}^{2}d^{-p_{*}\lor 2}\|\boldsymbol{g}^{2t}\|^{2}-\frac{1}{2}c_{\eta}^{3}c_{\xi}^{3}d^{-\frac{3p_{*}}{2}\lor 3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}|\|\boldsymbol{g}^{2t}\|^{2}.$ (B.32)
+We can easily see that $\mathbb{E}[\|\boldsymbol{g}^{2t}\|^{2}]\lesssim d$ and $\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}|\|\boldsymbol{g}^{2t}\|^{2}]\lesssim d$, which implies that the expectation of the last two terms of (B.32) is bounded by $\lesssim\kappa^{2t}c_{\eta}^{2}c_{\xi}^{2}d^{-(p_{*}-1)\lor 1}\lor c_{\eta}^{3}c_{\xi}^{3}d^{-(\frac{3p_{*}}{2}-1)\lor 2}\leq c_{\eta}^{2}c_{\xi}^{2}d^{-(p_{*}-1)\lor 1}\kappa^{2t}$.
+Now we bound $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}]$ by $\gtrsim c_{\eta}^{I-1}\kappa^{p_{*}-1}$.
+Let $C_{\sigma}$ be the maximum degree of the activation function with non-zero coefficients of Hermite expansion, which may be infinity when we consider general link functions, and there appear some infinite sums.
+For these cases we simply assume the sums converge – we discuss the validity of this condition in Section B.1.2.
+We omit the subscript $2t$ in the following for simplicity.
+We divide the analysis into the two cases.
+(I) If $I=1\Leftrightarrow\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})=p_{*}$.
+For the first term of $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$, we have
+$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ $=\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}\bigg{(}\sum_{j=p_{*}}^{\infty}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\bigg{)}\boldsymbol{x}\bigg{]}$ (B.33)
+$=\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\sum_{j=p_{*}}^{\infty}\bigg{[}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j}\boldsymbol{\omega}\bigg{]}$ (B.34)
+$=\sum_{j=p_{*}}^{C_{\sigma}}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}$ (B.35)
+$=p_{*}!\alpha_{p_{*}}\beta_{p_{*}}\kappa^{p_{*}-1}+O(\kappa^{p_{*}}).$ (B.36)
+For the second term of $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$, the following decomposition can be made.
+$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.37)
+$=\sum_{i=1}^{C_{\sigma}-1}(i!)^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}y\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}y\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}\boldsymbol{x}\bigg{]}+\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.38)
+$=\sum_{i=1}^{C_{\sigma}-1}(i!)^{-1}\eta^{i}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2i}y^{i+1}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}\boldsymbol{x}\bigg{]}$ (B.39)
+$\quad+p_{*}!\alpha_{p_{*}}\beta_{p_{*}}\kappa^{p_{*}-1}+O(\kappa^{p_{*}}).$ (B.40)
+We evaluate each term in the summation.
+We need to show that although $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ is a function of $\boldsymbol{\theta}^{\top}\boldsymbol{x}$ and $\boldsymbol{\omega}^{\top}\boldsymbol{x}$, it is mostly independent from the two quantities.
+To verify this, let $\boldsymbol{e}=\frac{\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}}{\|\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}\|}$ be the orthogonal component of $\boldsymbol{\theta}$ to $\boldsymbol{\omega}$.
+Then, we have that
+$\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ $=\boldsymbol{x}^{\top}(I-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}-\boldsymbol{e}\boldsymbol{e}^{\top})\boldsymbol{x}+(\boldsymbol{e}^{\top}\boldsymbol{x})^{2}$ (B.41)
+$=\underbrace{\boldsymbol{x}^{\top}(I-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}-\boldsymbol{e}\boldsymbol{e}^{\top})\boldsymbol{x}}_{\sim\chi_{d-2}^{2},\text{ independent from $\boldsymbol{\omega}^{\top}\boldsymbol{x}$ and $\boldsymbol{\theta}^{\top}\boldsymbol{x}$}}+\bigg{(}\frac{\boldsymbol{\theta}^{\top}\boldsymbol{x}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}^{\top}\boldsymbol{x}}{\|\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}\|}\bigg{)}^{2}.$ (B.42)
+With $P_{\boldsymbol{\omega},\boldsymbol{\theta}}=\boldsymbol{I}_{d}-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}-\boldsymbol{e}\boldsymbol{e}^{\top}$, (B.39) is expanded as
 […]
+$\quad\quad\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}(\boldsymbol{x}^{\top}P_{\boldsymbol{\omega},\boldsymbol{\theta}}\boldsymbol{x})^{i-j}\varsigma^{i+1-l}(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l}(\boldsymbol{\theta}^{\top}\boldsymbol{x}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}^{\top}\boldsymbol{x})^{2j}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}\boldsymbol{x}\bigg{]}$ (B.44)
+$=\sum_{i=1}^{C_{\sigma}-1}\sum_{j=0}^{i}\sum_{l=0}^{i+1}\sum_{k=0}^{2j}\frac{{i\choose j}{i+1\choose l}{2j\choose k}\eta^{i}\kappa^{k}(-1)^{k}\mathbb{E}[\varsigma^{i+1-l}]\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{i-j}]}{i!\|\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}\|^{2j}}$ (B.45)
+$\quad\quad\mathbb{E}\bigg{[}(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{2j-k}(\boldsymbol{\omega}^{\top}\boldsymbol{x})^{k}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}(\boldsymbol{\theta}^{\top}\boldsymbol{x}-\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\boldsymbol{\omega}^{\top}\boldsymbol{x})\bigg{]}$ (B.46)
+For a general differentiable function $g(\boldsymbol{x})$, we have $\mathbb{E}[\mathsf{He}_{t}(x_{1})g(\boldsymbol{x})]=\mathbb{E}[\frac{\mathrm{d}^{t}}{\mathrm{d}x_{1}^{t}}g(\boldsymbol{x})]$.
+If $g(\boldsymbol{x})$ is a polynomial (with a bounded coefficients) of $x_{1}$ and $\boldsymbol{u}^{\top}\boldsymbol{x}$ and its degree with respect to $x_{1}$ is at most $s(\leq t)$, $|\mathbb{E}[\mathsf{He}_{t}(x_{1})g(\boldsymbol{x})]|\lesssim|u_{1}|^{t-s}$, because differentiation of $g(\boldsymbol{x})=\bar{g}(x_{1},\boldsymbol{u}^{\top}\boldsymbol{x})$ is taken with respect to the first variable at most $s$ times.
+Each term of (B.46) is an expectation of $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l}$, multiplied by the polynomial of $\boldsymbol{\theta}^{\top}\boldsymbol{x}$ and $\boldsymbol{\omega}^{\top}\boldsymbol{x}$, where its degree with respect to $\boldsymbol{\theta}^{\top}\boldsymbol{x}$ is at most $2j-k$.
+Thus each term of (B.46) is evaluated as (here we omit the constants)
+$\frac{\eta^{i}\kappa^{k}(-1)^{k}\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{i-j}]}{i!\|\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}\|^{2j}}\mathbb{E}\bigg{[}\underbrace{(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l}}_{\mathrm{IE}\geq p_{*}}\underbrace{(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{2j-k}(\boldsymbol{\omega}^{\top}\boldsymbol{x})^{k}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}(\boldsymbol{\theta}^{\top}\boldsymbol{x}-\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\boldsymbol{\omega}^{\top}\boldsymbol{x})}_{\text{degree w.r.t.
+$\boldsymbol{\theta}^{\top}\boldsymbol{x}$ is at most $2j-k+1$}}\bigg{]}$ (B.47)
+$\lesssim c_{\eta}^{i}d^{-i}d^{i-j}\kappa^{k}\kappa^{((p_{*}-2j+k-1)\lor 0)}\lesssim c_{\eta}d^{-j}\kappa^{((p_{*}-2j-1)\lor 0)}\leq c_{\eta}\kappa^{p_{*}-1}(d/\kappa^{2})^{-j}\leq c_{\eta}\kappa^{p_{*}-1}.$ (B.48)
+The lower bound follows in the same fashion. Therefore,
+[…]
+Now, $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$ can be evaluated as
+[…]
+(II) If $I=\{\min\ i\mid\mathrm{IE}(\sigma_{*}^{i})=\mathrm{GE}(\sigma_{*})=p_{*}\}\geq 2$.
+Note that $\alpha_{j}=0$ for all $j\leq p_{*}$ from the assumption. Following (B.36), the first term of $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$ is evaluated as
+$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=\sum_{j=p}^{C_{\sigma}}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}=O(\kappa^{p_{*}}).$ (B.51)
+For the second term of $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$, similarly to (B.46), the following decomposition can be made.
+$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.52)
+$=\sum_{i=0}^{C_{\sigma}-1}\sum_{j=0}^{i}\sum_{l=0}^{i+1}\sum_{k=0}^{2j}\frac{{i\choose j}{i+1\choose l}{2j\choose k}\eta^{i}\kappa^{k}(-1)^{k}\mathbb{E}[\varsigma^{i+1-l}]\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{i-j}]}{i!\|\boldsymbol{\theta}-(\boldsymbol{\theta}^{\top}\boldsymbol{\omega})\boldsymbol{\omega}\|^{2j}}$ (B.53)
+$\quad\quad\mathbb{E}\bigg{[}(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{2j-k}(\boldsymbol{\omega}^{\top}\boldsymbol{x})^{k}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}(\boldsymbol{\theta}^{\top}\boldsymbol{x}-\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\boldsymbol{\omega}^{\top}\boldsymbol{x})\bigg{]}.$ (B.54)
+Each term of (B.54) (omitting constants) is evaluated as
+$\eta^{i}\kappa^{k}\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{i-j}]\mathbb{E}\bigg{[}(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{2j-k}(\boldsymbol{\omega}^{\top}\boldsymbol{x})^{k}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}(\boldsymbol{\theta}^{\top}\boldsymbol{x}-\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\boldsymbol{\omega}^{\top}\boldsymbol{x})\bigg{]}$ (B.55)
+$=c_{\eta}^{i}\kappa^{k}d^{-j}\mathbb{E}\bigg{[}\underbrace{(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l}}_{\mathrm{IE}\geq\begin{cases}p_{*}\ (l\geq I)\\ p_{*}+1\ (l<I)\end{cases}}\underbrace{(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{2j-k}(\boldsymbol{\omega}^{\top}\boldsymbol{x})^{k}\sigma^{(i+1)}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{(1)}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{i}(\boldsymbol{\theta}^{\top}\boldsymbol{x}-\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\boldsymbol{\omega}^{\top}\boldsymbol{x})}_{\text{degree w.r.t.
+$\boldsymbol{\theta}^{\top}\boldsymbol{x}$ is at most $2j-k+1$}}\bigg{]}$ (B.56)
+$\lesssim c_{\eta}^{i}\kappa^{k}d^{-j}\kappa^{(\mathrm{IE}(\sigma_{*}^{l})-2j+k-1)\lor 0}$ (B.57)
+When $i\leq I-2$ and $\eta=c_{\eta}d^{-1}$, we have $l\leq i+1<I$ and $\mathrm{IE}((\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l})\geq p_{*}+1$. Thus
+(B.57)$\lesssim\kappa^{p_{*}}$ (B.58)
+When $i\geq I$, $\mathrm{IE}((\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{l})\geq p_{*}$ and we get
+(B.57)$\lesssim c_{\eta}^{I}\kappa^{p_{*}-1}.$ (B.59)
+Now the case of $i=I-1$. When $i=I-1$ and $j\neq 0$, and using the assumption that $\kappa\leq c_{\eta}$,
+(B.57)$\lesssim c_{\eta}^{I-1}\kappa^{p_{*}-1}(\kappa^{-2}/d)\leq c_{\eta}^{I}\kappa^{p_{*}-1}.$ (B.60)
+When $i=I-1$, $j=0$, and $k\neq 0$,
+(B.57)$\lesssim c_{\eta}^{I-1}\kappa^{p_{*}}.$ (B.61)
+When $i=I-1$, $j=0$, $k=0$, and $l\leq I-1$,
+(B.57)$\lesssim c_{\eta}^{I-1}\kappa^{p_{*}}.$ (B.62)
+Therefore, except for $i=I-1$, $j=0$, $k=0$, and $l\leq I-1$, we can bound (B.57) by $\lesssim c_{\eta}^{I}\kappa^{p_{*}-1}+\kappa^{p_{*}}$.
+The lower bound follows in the same way.
+Finally, consider the case of $i=I-1$, $j=0$, $k=0$, and $l=I$.
+[…]
+$=\eta^{I-1}\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{I-1}]\sum_{m=p_{*}}^{C_{\sigma}I}m!H(\sigma_{*}^{I};m)H(\sigma^{(I)}(\sigma^{(1)})^{I-1};m-1)(1-\kappa^{2})\kappa^{m-1}$ (B.64)
+$=c_{\eta}^{I-1}p_{*}!d^{-(I-1)}\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{I-1}]H(\sigma_{*}^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)(1-\kappa^{2})\kappa^{p_{*}-1}+O(c_{\eta}^{I-1}\kappa^{p_{*}}).$ (B.65)
+Putting it all together (recovering the constants omitted in (B.55) again),
+[…]
+$=c_{\eta}^{I-1}\frac{p_{*}!d^{-(I-1)}\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{I-1}]}{(I-1)!}H(\sigma_{*}^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)\kappa^{p_{*}-1}+O(c_{\eta}^{I}\kappa^{p_{*}-1}+\kappa^{p_{*}}),$ (B.67)
+and
+[…]
+$=c_{\eta}^{I-1}\underbrace{\frac{p_{*}!d^{-(I-1)}\mathbb{E}_{z\sim\chi^{2}_{d-2}}[z^{I-1}]}{(I-1)!}}_{=\Theta(1)}H(\sigma_{*}^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)\kappa^{p_{*}-1}+O(c_{\eta}^{I}\kappa^{p_{*}-1}+\kappa^{p_{*}}).$ (B.69)
+Combining (i) and (ii), we have
+$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]\geq 2c_{\eta}^{I-1}c_{\sigma}\kappa^{p_{*}-1}+O(c_{\eta}^{I}\kappa^{p_{*}-1}+\kappa^{p_{*}})$ (B.70)
+for a positive constant $c_{\sigma}=\Theta(1)$.
+Here $c_{\sigma}>0$ satisfies $2c_{\sigma}=2p_{*}!\alpha_{p_{*}}\beta_{p_{*}}$ (for (i)) or $2c_{\sigma}=\frac{p_{*}!H(\sigma_{*}^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)}{(I-1)!}$ (for (ii)).
+Going back to (B.32), by setting $\nu^{2t}=(\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}])$, we have
+$\kappa^{2(t+1)}$ $\geq\kappa^{2t}+2c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}]+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}(\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}^{2t}])+O(c_{\eta}^{2}c_{\xi}^{2}d^{-(p_{*}-1)\lor 1}\kappa^{2t})$ (B.71)
+$=\kappa^{2t}+2c_{\eta}^{I}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}c_{\sigma}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\nu^{2t}$ (B.72)
+$\quad\quad+O\mathopen{}\mathclose{{}\left(c_{\eta}^{2}c_{\xi}^{2}d^{-(p_{*}-1)\lor 1}(\kappa^{2t})^{2t}+c_{\eta}^{I+1}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}}}\right).$ (B.73)
+When $c_{\xi}\leq c_{\eta}^{I}$ and $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\kappa\leq c_{\eta}^{I}$, terms in the big-$O$ notation is smaller than $c_{\eta}^{I}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}c_{\sigma}(\kappa^{2t})^{p_{*}-1}$ and we have
+$\kappa^{2(t+1)}\geq\kappa^{2t}+c_{\eta}^{I}c_{\xi}c_{\sigma}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\nu^{2t}.$ (B.74)
+It is straightforward to check $\nu^{2t}$ has sub-Weibull tail. ∎
+B.4 Weak Recovery: Stochastic Update
+This subsection proves weak recovery using the results on population update from the previous section. Specifically, from the previous section, we know that
+$\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(t+1)}\geq\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}+c_{\eta}^{I}c_{\xi}c_{\sigma}d^{-\frac{p_{*}}{2}\lor 1}(\kappa^{2t})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}\lor 1}\nu^{2t},$ (B.75)
+with the mean-zero sub-Weibull random variable $\nu^{2t}$ and a positive $c_{\sigma}=\Theta(1)$.
+For notational simplicity we write $c_{\eta}^{I}c_{\sigma}=c_{1}$.
+The following lemma is a detailed version of Proposition 9.
 Lemma 17.
-Starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $C_{2}d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot\eta\tilde{\eta}c_{1}^{I-1}H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};1\bigg{)}H\big{(}(\sigma_{*})^{I};2\big{)}\kappa+\eta\tilde{\eta}\nu,$ (B.16)
-where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
-Proof.  We first compute one gradient step from $\boldsymbol{w}=\boldsymbol{\omega}$ with a fresh sample $(\boldsymbol{x},y)$.
-~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{x}.$ (B.17)
-Then, with a projection matrix $\boldsymbol{P}_{\boldsymbol{\omega}}=I-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}$, the updated parameter becomes
-$\boldsymbol{w}\leftarrow\boldsymbol{w}+\boldsymbol{P}_{\boldsymbol{\omega}}\eta y\sigma^{\prime}(\boldsymbol{w}^{\top}\boldsymbol{x})\boldsymbol{x}=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x},$ (B.18)
-and the next gradient step with the same sample is computed as
-~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})$ $=\eta y\sigma^{\prime}(\boldsymbol{w}^{\top}\boldsymbol{x})\boldsymbol{x}$ (B.19)
-$=y\sigma^{\prime}\big{(}(\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})^{\top}\boldsymbol{x}\big{)}\boldsymbol{x}$ (B.20)
-$=y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{x},$ (B.21)
-here we used the notation $\|\boldsymbol{v}\|_{\boldsymbol{A}}^{2}=\boldsymbol{v}^{\top}\boldsymbol{A}\boldsymbol{v}$ for a vector $\boldsymbol{v}\in\mathbb{R}^{d}$ and a positive symmetric matrix $\boldsymbol{A}\in\mathbb{R}^{d\times d}$.
-From (B.18) and (B.21), the parameter after the two steps is obtained as
-$\boldsymbol{w}$ $\leftarrow\boldsymbol{w}+\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})$ (B.22)
-$=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+\eta y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}$ (B.23)
-$=\boldsymbol{\omega}+\eta\boldsymbol{g},$ (B.24)
-where we defined
-$\boldsymbol{g}:=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.25)
-Finally, normalization yields
-$\boldsymbol{w}$ $\leftarrow\frac{\boldsymbol{w}-\xi(\boldsymbol{w}-\boldsymbol{\omega})}{\|\boldsymbol{w}-\xi(\boldsymbol{w}-\boldsymbol{\omega})\|}=\frac{\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}.$ (B.26)
-Therefore, the update of the alignment is
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}$ $=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{(1+\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2})^{\frac{1}{2}}}$ (B.27)
-$\geq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}.$ (B.28)
-On the other hand, we have
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}.$ (B.29)
-We evaluate the expectation of (B.28). For the $j$-th Hermite polynomial $\mathsf{He}_{j}$ and $\boldsymbol{u}\in\mathbb{S}^{d-1}$, we have that
-$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{i}({\sf e}_{1}^{\top}\boldsymbol{x})f(\boldsymbol{u}^{\top}\boldsymbol{x}){\sf e}_{1}^{\top}\boldsymbol{x}]$ (B.30)
-$=j({\sf e}_{1}^{\top}x)^{j-1}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j-1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]+({\sf e}_{1}^{\top}\boldsymbol{x})^{j+1}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})],$ (B.31)
-$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{j}(x_{1})f(\boldsymbol{u}^{\top}\boldsymbol{x}){\sf e}_{2}^{\top}\boldsymbol{x}]=({\sf e}_{1}^{\top}\boldsymbol{x})^{j}({\sf e}_{2}^{\top}\boldsymbol{x})\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})].$ (B.32)
-Therefore,
-$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{j}({\sf e}_{1}^{\top}\boldsymbol{x})f(\boldsymbol{u}^{\top}\boldsymbol{x})\boldsymbol{x}]$ (B.33)
-$=\begin{pmatrix}j({\sf e}_{1}^{\top}\boldsymbol{x})^{j-1}\\ 0\\ \vdots\\ 0\end{pmatrix}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j-1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]{\sf e}_{1}+({\sf e}_{1}^{\top}\boldsymbol{x})^{j}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]\boldsymbol{u}.$ (B.34)
-Hence the first term of $\boldsymbol{g}$ (B.25) can be exapanded as
-$\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ $=\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}\bigg{(}\sum_{j=p}^{q}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}j\sum_{j=0}^{C_{q}}\beta_{j}\mathsf{He}_{j}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\bigg{)}\boldsymbol{x}\bigg{]}$ (B.35)
-$=\boldsymbol{P}_{\boldsymbol{\omega}}\sum_{j=p}^{q}\bigg{[}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j}\boldsymbol{\omega}\bigg{]}$ (B.36)
-$=\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}.$ (B.37)
-The coefficient is evaluated as
-$\bigg{|}\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\bigg{|}\lesssim\kappa^{p-1}\leq\kappa^{2}.$ (B.38)
-For the second term of $\boldsymbol{g}$ (B.25), we first bound the difference in replacing $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ with $d$,
-$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}$ (B.39)
-$=2\big{|}\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)h(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2},\boldsymbol{\omega}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})]\big{|},$ (B.40)
-where $h$ is a polynomial with degree at most $(C_{q}+q)^{C_{q}-1}+q+1$ and coefficients are all $O(1)$. (B.40) is further upper bounded as
-(B.40)$\leq 2\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)^{2}]^{\frac{1}{2}}\mathbb{E}[(h(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2},\boldsymbol{\omega}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}))^{2}]^{\frac{1}{2}},$ (B.41)
-by Cauchy-Schwarz inequality.
-$\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)^{2}]^{\frac{1}{2}}=\eta(2d-1)^{\frac{1}{2}}$, and the expectation of $\mathbb{E}[h^{2}]$ is $O(1)$ when $\eta\leq d^{-1}$.
-Therefore, (B.40) is bounded by $C_{1}\eta d^{\frac{1}{2}}$.
-Now, we consider $\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$.
-The following decomposition can be made.
-$\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.42)
-$=\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.43)
-$=\sum_{k=0}^{j-1}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose k}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.44)
-We evaluate each term of (B.44) except for $k=I-1$. Each term of (B.44) is a constant multiple of $\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}$. and we can evaluate the constant by
-$\bigg{|}\boldsymbol{\theta}^{\top}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.45)
-$=(\eta d)^{k}\bigg{|}\mathbb{E}\bigg{[}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{k}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}^{k+1}\mathsf{He}_{j-k-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}.$ (B.46)
-When $k\leq I-2$, $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{k+1}$ has information exponent larger than $2$. Therefore, we have
+Take $\eta^{2t},\eta^{2t+1}=\eta=c_{\eta}d^{-1}$, $\xi^{2(t+1)}=\xi=1-c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$.
+Suppose that the link function satisfies $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}$ (we choose the smallest such $I$) and activation functions satisfy all of the assumptions in Section B.1 for weak recovery.
+Let
+$T_{1,1}=C_{3}c_{\xi}^{-1}\begin{cases}d&(\text{if $p_{*}=\mathrm{GE}(\sigma_{*})=1$})\\ d(\log d)&(\text{else if $p_{*}=\mathrm{GE}(\sigma_{*})=2$})\\ d^{p_{*}-1}&(\text{else $p_{*}=\mathrm{GE}(\sigma_{*})\geq 3$}),\end{cases}$ (B.76)
+and take $c_{\xi}\lesssim\delta\mathrm{poly}(c_{\eta})$, $c_{2}\gtrsim\mathrm{poly}(c_{\eta})$, and $C_{3}\simeq c_{1}^{-1}$.
+If $\kappa^{0}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$, there exists some $\tau_{*}\leq T_{1,1}$ such that
+$\kappa^{2\tau_{*}}\geq 2c_{2},$ (B.77)
+with probability at least $1-\delta$, and $\kappa^{2\tau}\geq 2c_{2}$ for all $\tau_{*}\leq\tau\leq T_{1,1}$, with high probability.
+We may take $\delta=o_{d}(1)$ with arbitrarily slow decay.
+The proof is adapted from [BAGJ21], but our bound on $T_{1,1}$ is slightly sharper (by a $\log d$ factor for $p_{*}=2$ and by a $(\log d)^{2}$ factor for $p_{*}\geq 3$).
+For $p_{*}=2$, this is because of a trick that we carefully “restart” the dynamics, whose failure probability exponentially decays.
+Proof.  We divide the proof into the following cases.
+(i) When $p_{*}=1$.
+Note that $\{\sum_{s=0}^{\tau}\nu^{2s}\}_{\tau}$ is Martingale with $\mathbb{E}[(\nu^{2s})^{2}]\lesssim 1$.
+By Doob’s maximal inequality and Markov’s inequality, with probability $1-\delta$, we have
+$\max_{0\leq\tau\leq T}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}\bigg{|}^{2}\leq\delta^{-1}\mathbb{E}[(\sum_{s=0}^{T}\nu^{2s})^{2}]\leq\delta^{-1}\sum_{s=0}^{T}\mathbb{E}[(\nu^{2s})^{2}]\leq C_{1}\delta^{-1}(T+1)$ (B.78)
+for any fixed $T\geq 0$, with a sufficiently large constant $C_{1}=\Theta(1)$. In the following we consider the case when (B.78) holds for $T=c_{1}^{-1}c_{\xi}^{-1}d-1$.
+If $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\kappa^{2t}\leq c_{\eta}^{I}$ for all $t=0,1,\cdots,\tau$, we have
+$\kappa^{2(\tau+1)}$ $\geq\kappa^{2\tau}+c_{1}c_{\xi}d^{-1}+c_{\eta}c_{\xi}d^{-1}\nu^{2\tau}$ (B.79)
+$\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}+c_{1}c_{\xi}(\tau+1)d^{-1}\gamma-c_{\eta}c_{\xi}d^{-1}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}\bigg{|}.$ (B.80)
+Now, applying (B.78) to get
 […]
-$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{k}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{k-l+1}\mathsf{He}_{j-k-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.48)
-$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\sum_{m=3}^{\infty}m!\kappa^{q(k-l+1)}H\big{(}(\sigma_{*})^{k-l+1};m\big{)}H\big{(}(\sigma^{\prime})^{k}\mathsf{He}_{j-k-1};m-1\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}\bigg{|}$ (B.49)
-$\lesssim(\eta d)^{k}\kappa^{2}.$ (B.50)
-When $k\geq I$, we know that $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{k+1},\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ have information exponent larger than $1$, and $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1},\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))$ have information exponent larger than $2$.
-Thus, the expectation in (B.46) is bounded by
-[…]
-$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\sum_{m=2}^{\infty}m!\kappa^{q(k-l+1)}H\big{(}(\sigma_{*})^{k-l+1};m-1\big{)}H\big{(}(\sigma^{\prime})^{k}\mathsf{He}_{j-k-1};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}\bigg{|}$ (B.52)
-$\lesssim(\eta d)^{k}\kappa.$ (B.53)
-Now, consider the case when $k=I-1$:
-$\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}\big{)}^{I}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.54)
-$=\sum_{l=0}^{I}\sum_{j=1}^{C_{\sigma}}(\eta d)^{k}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.55)
-Note that $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1}$ has information exponent larger than $2$.
-Therefore, for $l\geq 1$, we have
-$\bigg{|}\boldsymbol{\theta}^{\top}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.56)
-$\lesssim(\eta d)^{I-1}\bigg{|}\mathbb{E}\bigg{[}\sum_{m=3}^{\infty}m!\kappa^{m-1}H\big{(}\mathsf{He}_{j-I}(\sigma^{\prime})^{I-1};m-1\big{)}H\big{(}\big{(}\sigma_{*}\big{)}^{I-l};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.57)
-$\lesssim(\eta d)^{I-1}\kappa^{2}.$ (B.58)
-And for $l=0$, $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ has information exponent $2$ and we have
-$\sum_{j=1}^{C_{\sigma}}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}\mathsf{He}_{j-I}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.59)
-$=(\eta d)^{I-1}\sum_{m=2}^{\infty}m!\kappa^{m-1}H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};m-1\bigg{)}H\big{(}(\sigma_{*})^{I};m\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$ (B.60)
-If $H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};1\bigg{)}\neq 0$, we have
-(B.60)$=(1+O(\kappa))\cdot(\eta d)^{I-1}\underbrace{2H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};1\bigg{)}H\big{(}(\sigma_{*})^{I};2\big{)}}_{=:\gamma}\kappa\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}.$ (B.61)
-We have that
-$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.62)
-$=(1+O(\kappa))\cdot(\eta d)^{I-1}\gamma\kappa\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+O(\kappa^{2}+C_{1}\eta d^{\frac{1}{2}}+(\eta d)^{I-2}\kappa^{2}+(\eta d)^{I}\kappa+(\eta d)^{I-1}\kappa^{2})$ (B.63)
-When $d^{-\frac{1}{2}}C_{3}\leq\kappa\leq c_{2}$ and $\eta=c_{1}d^{-1}$, we have
-$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa$ (B.64)
-Together with the bound on the first term ((B.37) and (B.38)), the expectation of $\boldsymbol{g}$ (B.25) is evaluated as
-$\boldsymbol{\theta}^{\top}\mathbb{E}[\boldsymbol{g}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa.$ (B.65)
-By using this, the expected update of the alignment becomes
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\mathbb{E}[\|\boldsymbol{g}\|^{2}]-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}].$ (B.66)
-Note that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\kappa\lesssim c_{1}$. Thus,
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa-C_{1}\eta\tilde{\eta}^{2}(\kappa\eta d+\eta^{2}\tilde{\eta}d).$ (B.67)
-When $\tilde{\eta}\leq c_{2}$, $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ is evaluated as
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa.$ (B.68)
-In the same way, using (B.29), we also have the opposite bound:
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\leq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa$ (B.69)
-Regarding the noise, recall that
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}.$ (B.70)
-$\eta^{-1}\tilde{\eta}^{-1}\big{(}\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}[\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}]\big{)}=\ \boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$ is a mean-zero polynomial of Gaussian inputs, with all coefficients and variances of inputs bounded by $O(1)$.
-Notice that normalization does not increase the absolute value of the noise.
-Thus, regarding $\nu=\eta^{-1}\tilde{\eta}^{-1}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{w}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\big{)}$, we have that
-$\mathbb{P}\big{[}\big{|}\eta^{-1}\tilde{\eta}^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{w}-\mathbb{E}[\eta^{-1}\tilde{\eta}^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{w}]\big{|}>t\big{]}\leq\mathbb{P}\bigg{[}\bigg{|}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}\bigg{[}\boldsymbol{\theta}^{\top}\boldsymbol{g}\bigg{]}\bigg{|}>t\bigg{]}\leq\exp(-t^{1/C_{1}}/C_{1}).$ (B.71)
-This completes the proof. ∎
-B.3.2 When $\mathrm{IE}(\sigma_{*}^{I})=1$ with $I\geq 2$
-[…]
+when $\tau\leq c_{1}^{-1}c_{\xi}^{-1}d-1$.
+By letting $c_{\xi}\leq c_{\eta}^{-4}c_{1}C_{1}^{-1}\delta$, we have $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq c_{\eta}c_{\xi}^{\frac{1}{2}}c_{1}^{-\frac{1}{2}}C_{1}^{\frac{1}{2}}\delta^{-\frac{1}{2}}d^{-\frac{1}{2}}$, and
+$\kappa^{2(\tau+1)}\geq c_{\eta}^{-1}d^{-\frac{1}{2}}+c_{1}c_{\xi}(\tau+1)d^{-1},$ (B.82)
+which verifies $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\kappa^{2t}$ for $t=\tau+1$. Thus, there exists some $\tau_{*}\leq c_{1}^{-1}c_{\xi}^{-1}d$ such that
+$\kappa^{2\tau_{*}}\geq 4c_{2},$ (B.83)
+for $c_{1}\leq\frac{1}{4}c_{\eta}^{I}$, with probability $1-\delta$.
+Now we prove that $\kappa^{2t}\geq 2c_{2}$ holds for all $\tau_{*}\leq t\leq T_{1,1}=C_{3}c_{\xi}^{-1}d$.
+Because $\nu^{2t}$ are mean-zero sub-Weibull random variables, we also have that $|\sum_{s=\tau}^{\tau+\tau^{\prime}-1}\nu^{2s}|\leq C_{4}\sqrt{\tau^{\prime}}$ for all $0\leq\tau,\tau^{\prime}\leq T_{1,1}$ with high probability.
+Also, because $\eta^{t}\ll d^{-1}$ and $|1-\xi^{t}|\ll 1$, we can easily see that $|\kappa^{2(\tau+1)}-\kappa^{2\tau}|=\tilde{O}(d^{-1})$ for all $\tau=0,1,\cdots,T_{1,1}-1$, with high probability.
+Thus, when there exists $\tau\geq\tau_{*}$ such that $\kappa^{2(\tau-1)}\geq 4c_{2}$ and $\kappa^{2\tau}<4c_{2}$, we have $\kappa^{2\tau}\geq 3c_{2}$ with high probability.
+Moreover, following the above argument, we can inductively show that
+$\kappa^{2(\tau+\tau^{\prime})}$ $\geq 3c_{2}+c_{1}c_{\xi}\tau^{\prime}d^{-1}-c_{\eta}c_{\xi}d^{-1}C_{4}\sqrt{\tau^{\prime}}$ (B.84)
+$\geq 3c_{2}+c_{1}c_{\xi}\tau^{\prime}d^{-1}-\begin{cases}c_{2}&(\tau^{\prime}\leq c_{\eta}^{-2}c_{\xi}^{-2}C_{4}^{-2}c_{2}^{2}d^{2})\\ c_{1}c_{\xi}\tau^{\prime}d^{-1}&(\tau^{\prime}\geq c_{\eta}^{2}c_{1}^{-2}C_{4}^{2})\end{cases}.$ (B.85)
+$\geq 2c_{2},$ (B.86)
+for $\tau^{\prime}\leq T_{1,1}=C_{3}c_{\xi}^{-1}d$ or until $\kappa^{2(\tau+\tau^{\prime})}\geq 4c_{2}$ holds again.
+By repeating this argument (if there are multiple such $\tau$), we see that $\kappa^{2t}\geq 2c_{2}$ holds for all $\tau_{*}\leq t\leq T_{1,1}=C_{3}c_{\xi}^{-1}d$ with high probability.
+(ii) When $p_{*}=2$.
+We define $\iota_{0}=0,\iota_{1}=\log_{(1+c_{1}c_{\xi}d^{-1})}(4),\iota_{2}=2\log_{(1+c_{1}c_{\xi}d^{-1})}(4),\dots$.
+We show that, for each $i$, if $\kappa^{2\iota_{i}}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$, we have $\kappa^{2(\iota_{i+1})}\geq 2\kappa^{2\iota_{i}}$, with probability at least $1-\delta 4^{-i}$, or there exists some $t\ (\iota_{i}<t\leq\iota_{i+1})$ with $\kappa^{2t}>c_{\eta}^{I}$.
+Assume that the above statement holds until some $i-1\geq 0$ (we do not need to assume anything for $i=0$).
+Then, we have $\kappa^{2\iota_{i}}\geq 2^{i}\kappa^{0}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$.
+Similarly to (B.80), if $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\kappa^{2t}\leq c_{\eta}^{I}$ for all $t=\iota_{i},\iota_{i}+1,\cdots,\tau$, we have
+$\kappa^{2(\tau+1)}\geq\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}-c_{\eta}c_{\xi}d^{-1}\bigg{|}\sum_{s=\iota_{i}}^{\tau}\nu^{2s}\bigg{|}.$ (B.87)
+Applying (B.78) with $\delta=\delta/4^{i}$ and $T=\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}-1$ to get
+$\kappa^{2(\tau+1)}$ $\geq\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}-c_{\eta}c_{\xi}d^{-1}C^{\frac{1}{2}}\delta^{-\frac{1}{2}}\sqrt{\tau+1-\iota_{i}}$ (B.88)
+$\geq\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}-\frac{1}{2}\kappa^{2\iota_{i}}$ (B.89)
+when $\tau\leq\iota_{i}+\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}-1$, which verifies $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\frac{1}{2}\kappa^{2\iota_{i}}\leq\kappa^{2t}$ for $t=\tau+1$.
+This implies that, with probability $1-\delta/4^{i}$, we have
+$\kappa^{2(\tau+1)}\geq\frac{1}{2}\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}$ (B.90)
+for all $\tau=\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}-1$, which is equivalent to
+$\kappa^{2\tau}\geq(1+c_{1}c_{\xi}d^{-1})^{\tau-\iota_{i}}\frac{1}{2}\kappa^{2\iota_{i}}$ (B.91)
+for all $\tau=\iota_{i},\iota_{i}+1,\cdots,\iota_{i}+\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}$.
+By taking $c_{\xi}\ll c_{1}c_{\eta}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d$, we have $\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}\geq\log_{(1+c_{1}c_{\xi}d^{-1})}(4)$, and we get
+$\kappa^{2\iota_{i+1}}\geq 2\kappa^{2\iota_{i}}$ (B.92)
+with probability $1-\delta/4^{i}$ (or there exists $t\leq\iota_{i+1}$ such that $\kappa^{2t}>c_{\eta}^{I}$).
+Thus, by induction, for all $i$, we have that
+$\kappa^{2\iota_{i}}\geq 2^{i}\kappa^{0},$ (B.93)
+or that there exists some $t\leq\iota_{i}$ such that $\kappa^{2t}$ is larger than $c_{\eta}^{I}$, with probability $1-\delta$.
+The LHS of (B.93) becomes larger than $c_{\eta}^{I}$ for some $i\leq\log d$.
+Because $\iota_{i}=\Theta(ic_{1}^{-1}c_{\xi}^{-1}d)$, within $O(c_{1}^{-1}c_{\xi}d\log d)$ steps, there exists at least one $\tau_{*}=O(c_{1}^{-1}c_{\xi}^{-1}d\log d)$ such that $\kappa^{2\tau_{*}}\geq 4c_{2}$ for $c_{2}\leq\frac{1}{4}c_{\eta}^{I}$, with probability $1-\delta$.
+Once such $\tau_{*}$ is obtained, following the last paragraph of (i), we can see that $\kappa^{2t}\geq 2c_{2}$ holds until $t=T_{1,1}$ with high probability.
+(iii) When $p_{*}\geq 3$. We apply (B.78) with $T=\frac{1}{p_{*}-2}c_{1}^{-1}c_{\xi}^{-1}d^{\frac{p_{*}}{2}}(\kappa^{0})^{-(p_{*}-2)}$ to obtain that
+$c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}\bigg{|}\leq c_{\eta}c_{\xi}^{\frac{1}{2}}c_{1}^{-\frac{1}{2}}C^{\frac{1}{2}}\delta^{-\frac{1}{2}}d^{-\frac{p_{*}}{4}}(\kappa^{0})^{-\frac{p_{*}-2}{2}}$ (B.94)
+for all $\tau=0,1,\cdots,T-1$, with probability $1-\delta$.
+We take $c_{\xi}\ll c_{\eta}^{-2}c_{1}C^{-1}\delta d^{\frac{p_{*}}{4}}(\kappa^{0})^{\frac{p_{*}}{2}}$ so that (B.94) is bounded by $c_{1}^{-1}d^{-\frac{1}{2}}$. Then,
+$\kappa^{2(\tau+1)}$ $\geq\kappa^{0}+c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}\sum_{s=0}^{\tau}(\kappa^{2s})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}}\sum_{s=0}^{\tau}\nu^{2s}$ (B.95)
+$\geq c_{\eta}^{-1}d^{-\frac{1}{2}}+c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}\sum_{s=0}^{\tau}(\kappa^{2\tau})^{p_{*}-1}.$ (B.96)
+It is easy to see that $\kappa^{2(\tau+1)}$ is lower bounded by $a^{\tau+1}$, where $a^{0}=c_{\eta}^{-1}d^{-\frac{1}{2}}$ and $a^{\tau+1}=a^{\tau}+c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}(a^{\tau})^{p_{*}-1}$.
+By applying Lemma 18, we have
+$\kappa^{2\tau}\geq\frac{\kappa^{0}}{\mathopen{}\mathclose{{}\left(1-c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}(p_{*}-2)(\kappa^{0})^{(p_{*}-2)}t}\right)^{\frac{1}{p_{*}-2}}}.$ (B.97)
+Thus, until $\tau\leq\big{(}c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}(p_{*}-2)(\kappa^{0})^{(p_{*}-2)}\big{)}^{-1}\leq T+1\ll d^{p_{*}-1}$, with probability at least $1-\delta$, there exists some $\tau_{*}$ such that
+$\kappa^{2\tau_{*}}\geq 4c_{2}\geq c_{\eta}^{I}$ (B.98)
+when $c_{2}\leq\frac{1}{4}c_{\eta}^{I}$.
+Once such $\tau_{*}$ is obtained, following the last paragraph of (i), we can see that $\kappa^{2t}\geq 2c_{2}$ holds until $t=T_{1,1}$ with high probability. ∎
+In the above proof we used the (discrete version of) Bihari–LaSalle inequality from [BAGJ22].
 Lemma 18.
-For the case of $\mathrm{IE}(\sigma_{*}^{I})=1$, starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot\eta\tilde{\eta}c_{1}^{I-1}H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}H\big{(}(\sigma_{*})^{I};1\big{)}+\eta\tilde{\eta}\nu,$ (B.72)
-where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
-Proof.  Similarly to Lemma 9, update of the alignment is evaluated as
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}\geq\kappa+\eta c_{2}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}$ (B.73)
-and
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta c_{2}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2},$ (B.74)
+For $p\geq 3$ and $c>0$, consider a positive sequence $(a^{t})_{t\geq 0}$ such that
+$a^{t+1}=a^{t}+c(a^{t})^{p-1}.$ (B.99)
+Then, we have
+$a^{t}\geq\frac{a^{0}}{\mathopen{}\mathclose{{}\left(1-c(p-2)(a^{0})^{(p-2)}t}\right)^{\frac{1}{p-2}}}.$ (B.100)
+Proof.  From definition, we have
+$c=\frac{a^{t+1}-a^{t}}{(a^{t})^{p-1}}\leq\int_{t=a^{t}}^{a^{t+1}}\frac{1}{x^{p-1}}\leq\frac{1}{p-2}\mathopen{}\mathclose{{}\left[\frac{1}{(a^{t})^{p-2}}-\frac{1}{(a^{t+1})^{p-2}}}\right].$ (B.101)
+Taking the summation and re-arranging the terms yield
+$(a^{t})^{-(p-2)}\leq(a^{0})^{-(p-2)}-c(p-2)t,$ (B.102)
+$\therefore a^{t}\geq\frac{a^{0}}{\mathopen{}\mathclose{{}\left(1-c(p-2)(a^{0})^{(p-2)}t}\right)^{\frac{1}{p-2}}},$ (B.103)
+which gives the lower bound. ∎
+B.5 From Weak Recovery to Strong Recovery
+In the previous subsection, we proved that after $t=2T_{1,1}=\tilde{\Theta}(d)$ steps, with probability $\tilde{\Omega}(1)$ over the randomness of initialization, we obtain nontrivial alignment $\kappa^{2T_{1,1}}_{j}\geq 2c_{2}$.
+This subsection discusses how to convert the weak recovery into the strong recovery.
+Lemma 19.
+Suppose the neuron satisfies $\kappa^{2T_{1,1}}\geq 2c_{2}$.
+Take $\eta^{2t}=\eta=\bar{c}_{\eta}\varepsilon d^{-1}$, $\eta^{2t+1}=0$, $\xi^{2(t+1)}=0$ for all $t\geq T_{1,1}$, where $\bar{c}_{\eta}\lesssim\mathrm{poly}(c_{1})$.
+If the activation functions satisfy all of the assumptions in Section B.1 for strong recovery, then we have
+$\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(T_{1,1}+\tau_{*})}\geq 1-\varepsilon,$ (B.104)
+with high probability, where $\tau_{*}\leq T_{1,2}=C_{3}d\varepsilon^{-2}$.
+Moreover, $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(T_{1,1}+t)}\geq 1-\varepsilon$ for all $\tau_{*}\leq t\leq T_{1,2}=C_{3}d\varepsilon^{-2}$, with high probability.
+Proof.
+Consider the Hermite expansions of $\sigma_{*}$ and $\sigma$.
+Let $p$ be the smallest degree that both $\sigma_{*}$ and $\sigma$ have non-zero coefficients.
+First we compute the population gradient (of the correlation term) as
+$\mathbb{E}\big{[}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}$ $=\mathbb{E}\bigg{[}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}\sum_{j=p}^{\infty}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}\sum_{j=0}^{\infty}\beta_{j}\mathsf{He}_{j}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\bigg{)}\bigg{]}$ (B.105)
+$=\sum_{j=p}^{\infty}\big{[}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j}\boldsymbol{w}^{2t}\big{]}.$ (B.106)
+Applying $P_{\boldsymbol{w}^{2t}}$, we have
+$\mathbb{E}\big{[}P_{\boldsymbol{w}^{2t}}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}=(\boldsymbol{\theta}-({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{\theta})\boldsymbol{w}^{2t})\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}.$ (B.107)
+Thus, the update of the alignment $\kappa^{2t}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}$ is
+$\kappa^{2(t+1)}\geq\kappa^{2t}+\eta\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\eta^{2}\kappa^{2t}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2},$ (B.108)
 where
-$\boldsymbol{g}=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.75)
-From (B.37) and (B.38) we know that the first term of $\boldsymbol{g}$ (B.75) is evaluated as
-$\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta},$ (B.76)
-where the sum of coefficients are bounded by $\lesssim\kappa$.
-We then consider the second term of $\boldsymbol{g}$ (B.75). We can replace $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ by $d$ with the following bound similarly to (B.40):
-$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}\leq C_{1}\eta d^{\frac{1}{2}}.$ (B.77)
-For the term $\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$.
-the following decomposition can be made.
-$\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.78)
-$=\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.79)
-$=\sum_{k=0}^{j-1}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose k}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.80)
-We evaluate each term of (B.80) except for $k=I-1$. Similarly to the bounds on (B.46), each term is a constant multiple of $\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$, and we want to bound
-$\bigg{|}\boldsymbol{\theta}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}.$ (B.81)
-When $k\leq I-2$, $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{k+1}$ has information exponent larger than $1$. Therefore, we have
-(B.81)$\lesssim(\eta d)^{k}\kappa.$ (B.82)
-When $k\geq I$, we have
-(B.81)$\lesssim(\eta d)^{k}\leq c_{1}^{I}.$ (B.83)
-Now we consider the case when $k=I-1$:
-$\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}\big{)}^{I}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.84)
-$=\sum_{l=0}^{I}\sum_{j=1}^{C_{\sigma}}(\eta d)^{k}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.85)
-Note that $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1}$ has information exponent larger than $1$.
-Therefore, for $l\geq 1$, we have
-$\bigg{|}\boldsymbol{\theta}^{\top}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.86)
-$\lesssim(\eta d)^{I-1}\bigg{|}\mathbb{E}\bigg{[}\sum_{m=2}^{\infty}m!\kappa^{m-1}H\big{(}\mathsf{He}_{j-I}(\sigma^{\prime})^{I-1};m-1\big{)}H\big{(}\big{(}\sigma_{*}\big{)}^{I-l};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.87)
-$\lesssim(\eta d)^{I-1}\kappa.$ (B.88)
-And for $l=0$, $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ has information exponent $1$ and we have
-$\sum_{j=1}^{C_{\sigma}}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}\mathsf{He}_{j-I}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.89)
-$=(\eta d)^{I-1}\sum_{m=1}^{\infty}m!\kappa^{m-1}H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};m-1\bigg{)}H\big{(}(\sigma_{*})^{I};m\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$ (B.90)
-If $H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}\neq 0$, we have
-(B.90)$=(1+O(\kappa))\cdot(\eta d)^{I-1}\underbrace{H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}H\big{(}(\sigma_{*})^{I};1\big{)}}_{=:\gamma}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}.$ (B.91)
-Note that
-$H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};0\bigg{)}=H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}.$ (B.92)
-Now we have that
-$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.93)
-$=(1+O(\kappa))\cdot(\eta d)^{I-1}\gamma\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+O(\kappa+C_{1}\eta d^{\frac{1}{2}}+(\eta d)^{I-2}\kappa+(\eta d)^{I}+(\eta d)^{I-1}\kappa).$ (B.94)
-When $d^{-\frac{1}{2}}C_{3}\leq\kappa\leq c_{2}$ and $\eta=c_{1}d^{-1}$, we have
-$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma$ (B.95)
-Together with the bound on the first term ((B.76)), the expectation of $\boldsymbol{g}$ is evaluated as
-$\boldsymbol{\theta}^{\top}\mathbb{E}[\boldsymbol{g}]=(1+O(c_{1}))\cdot c_{1}^{I-1}.$ (B.96)
-By using this and (B.73), the expected update of the alignment becomes
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\mathbb{E}[\|\boldsymbol{g}\|^{2}]-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}].$ (B.97)
-Note that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\kappa\lesssim c_{1}$.
-Taking $\tilde{\eta}\leq c_{2}$ yields that
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma.$ (B.98)
-The upper bound can be obtained in similar fashion,
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\leq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma.$ (B.99)
-Finally, the noise can be handled in the exact same way as that of Lemma 9, the details of which we omit. ∎
-B.3.3 When $\mathrm{IE}(\sigma_{*}^{I})=1,2$ with $I=1$
-We finally consider the case when $\mathrm{IE}[\sigma_{*}]=1$ or when $\mathrm{IE}[\sigma_{*}]=2$ and $\mathrm{IE}[\sigma_{*}^{i}]\geq 2$ for $i=2,\cdots,C_{\sigma}$.
-Lemma 11.
-Starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then for $\mathrm{IE}[\sigma_{*}]=1$, the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot 2\eta\tilde{\eta}\alpha_{1}\beta_{1}+\eta\tilde{\eta}\nu,$ (B.100)
-and when $\mathrm{IE}[\sigma_{*}]=2$ and $\mathrm{IE}[\sigma_{*}^{i}]\geq 2$ for all $i=2,\cdots,C_{q}$,
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot 4\eta\tilde{\eta}\alpha_{2}\beta_{2}\kappa+\eta\tilde{\eta}\nu,$ (B.101)
-where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
-Proof.  Similar to Lemma 9, the update of the alignment is
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}\geq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}$ (B.102)
-and
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}$ (B.103)
-where
-$\boldsymbol{g}=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.104)
-We first consider the case when $\mathrm{IE}[\sigma_{*}]=1$. We have
-$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=\alpha_{1}\beta_{1}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}=\alpha_{1}\beta_{1}(1-\kappa^{2}).$ (B.105)
-Because we take $\eta=c_{1}d^{-1}$ with a vanishing constant $c_{1}$ and assume that $\kappa\leq c_{2}$, we have
-$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}\lesssim\eta d\beta_{1}^{2}\lesssim c_{1}.$ (B.106)
-and
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]=2(1+O(c_{1}))\cdot\alpha_{1}\beta_{1}.$ (B.107)
-The first claim follows from the fact that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\eta=c_{1}d^{-1}$.
-Next we consider the case when $\mathrm{IE}[\sigma_{*}^{i}]=2$ for $i=1,2,\cdots,C_{q}$. We have
-$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=2\alpha_{2}\beta_{2}\kappa\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}=2\alpha_{2}\beta_{2}(1-\kappa^{2})\kappa.$ (B.108)
-On the other hand, because $y,\cdots,y^{C_{\sigma}}$ has information exponent larger than $1$, we have
-$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}\lesssim\eta d\kappa=c_{1}\kappa.$ (B.109)
-Thus, when $\kappa\leq c_{1}$,
-$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]=(1+O(c_{1}))\cdot 4\alpha_{2}\beta_{2}\kappa,$ (B.110)
-which establishes the second claim. ∎
-B.4 Stochastic Update
-As a result of the previous subsection, by choosing either of $\xi_{j}=1-\tilde{\eta}_{j}$ or $\xi_{j}=1+\tilde{\eta}_{j}$, we obtained that
-  • (i)
-    When $\mathrm{IE}[\sigma_{*}^{i}]=1$ for some $i\leq C_{\sigma}$, the update can be written as
-    $\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\gamma+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}\leq\kappa^{2(t+1)}_{j}\leq\kappa^{2t}_{j}+2\eta^{2t}_{j}\tilde{\eta}_{j}\gamma+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}.$ (B.111)
-  • (ii)
-    Otherwise, the update can be written as
-    $\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\gamma\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}\leq\kappa^{2(t+1)}_{j}\leq\kappa^{2t}_{j}+2\eta^{2t}_{j}\tilde{\eta}_{j}\gamma\kappa^{2t}_{j}+\eta^{2t}_{j}\tilde{\eta}_{j}\nu^{2t}_{j}.$ (B.112)
-Here $\eta_{j}^{2t}=\eta_{j}^{2t+1}=c_{1}d^{-1}$, $\tilde{\eta}_{j}\leq c_{2}$, and $\gamma\geq c_{2}$ that depends on $\sigma_{*}$.
-$\nu^{2t}_{j}$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu^{2t}_{j}|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
-We assumed that $d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for the former and $C_{2}d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for the latter.
-For each neuron $j$, we sample $T_{1,1,j}\sim\mathrm{Unif}(\{1,\cdots,T_{1,1}\})$ and let $\eta^{t}=c_{1}d^{-1}$.
-For $t=0,\cdots,2(T_{1,1,j}-1)$, we let $\xi_{j}^{t}\equiv 1-\tilde{\eta}_{j}$ or $\xi_{j}^{t}\equiv 1+\tilde{\eta}_{j}$ with equiprobability, where $\tilde{\eta}_{j}=c_{2}$, and $\xi_{j}^{t}=1$ for $t=2T_{1,1,j},\cdots,2(T_{1,1}-1)$.
-The goal of this subsection is to prove the following lemma.
-Lemma 12.
-Let $T_{1}=\tilde{\Theta}(d)$.
-If the initial alignment satisfies $\kappa_{j}^{0}\geq C_{2}d^{-\frac{1}{2}}$ (for (i)) or $\kappa_{j}^{0}\geq 2C_{2}d^{-\frac{1}{2}}$ (for (ii)), then we have $\frac{1}{4}c_{2}\leq\kappa^{2T_{1,1}}_{j}\leq c_{2}$ for at least $1/\mathrm{polylog}(d)$ fraction of neurons, with high probability.
-Proof.
-(i) the case of $\mathrm{IE}=1$.
-If $d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for all $t=0,1,\cdots,\tau\ (0\leq\tau\leq T_{1,j}-1)$, we have
-$\kappa^{2(\tau+1)}_{j}$ $\geq\kappa^{2\tau}_{j}+\eta^{\tau}\tilde{\eta}_{j}\gamma+\eta^{\tau}\tilde{\eta}_{j}\nu^{2\tau}_{j}$ (B.113)
-$=\kappa^{0}_{j}+\sum_{s=0}^{\tau}\eta^{s}\tilde{\eta}_{j}\gamma+\sum_{s=0}^{\tau}\eta^{s}\tilde{\eta}_{j}\nu^{2s}_{j}$ (B.114)
-$\geq 2d^{-\frac{1}{2}}+c_{1}c_{2}d^{-1}\tilde{\varepsilon}\gamma(s+1)-c_{1}c_{2}^{2}d^{-1}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}_{j}\bigg{|}.$ (B.115)
-With high probability, $\sum_{s=0}^{\tau}\nu^{2s}_{j}$ is bounded by $C_{3}\sqrt{\tau+1}$.
-If $\tau+1\leq 4\gamma^{-2}C_{3}^{2}$, by letting $c_{2}^{2}\leq\frac{1}{2}\gamma c_{1}^{-1}C_{3}^{-2}d^{\frac{1}{2}}$, we have $c_{1}c_{2}d^{-1}C_{3}\sqrt{\tau+1}\leq d^{-\frac{1}{2}}$.
-If $\tau+1\geq 4\gamma^{-2}C_{3}^{2}$, we have $C_{3}\sqrt{\tau+1}\leq\frac{1}{2}\gamma(\tau+1)$.
-Thus, in either case,
-(B.115)$\geq d^{-\frac{1}{2}}+\frac{1}{2}c_{1}c_{2}^{2}d^{-1}\gamma(\tau+1).$ ) . (B.116)
-This verifies that $\kappa^{2t}_{j}\geq C_{1}d^{-\frac{1}{2}}$ holds for $t=\tau+1$. By induction,
-$\kappa^{2t}_{j}$ $\geq d^{-\frac{1}{2}}+\frac{1}{2}c_{1}c_{2}^{2}d^{-1}\gamma(\tau+1)$ (B.117)
-holds for $t$ until $\kappa^{2t}$ gets larger than $c_{2}$.
-By letting $T_{1,1}\geq c_{1}^{-1}c_{2}^{-2}\gamma^{-1}d$, we have $\kappa^{2t}\geq\frac{1}{2}c_{2}$ for some $t\leq T_{1,1}$, with high probability.
-Now together with (B.113), we have
-$\kappa^{2(t+1)}_{j}$ $\leq\kappa^{2t}_{j}+\eta^{t}\tilde{\eta}_{j}\cdot 2\gamma+\eta^{t}\tilde{\eta}_{j}\nu^{2t}_{j}.$ (B.118)
-Hence we obtain the following bound with high probability,
-$|\kappa^{2(t+1)}_{j}-\kappa^{2t}_{j}|\leq\eta^{t}_{j}\tilde{\eta}_{j}(2\gamma+C_{3})\lesssim c_{1}c_{2}^{2}C_{3}d^{-1}=:\Delta_{1}.$ (B.119)
-When $\kappa^{2\tau}\geq\frac{1}{2}c_{1}$ holds for some $\tau$, with high probability, we have
-$\frac{1}{4}c_{2}\leq\kappa^{2\tau+s}\leq c_{2}$ (B.120)
-for all $0\leq s\leq\Delta_{1}/4c_{2}$.
-Because $1/4\Delta_{1}c_{2}=\tilde{\Theta}(d)$ and $T_{1,1}$ is also $\tilde{\Theta}(d)$, with probability $\tilde{\Theta}(1)$, $2T_{1,1,j}$ satisfies $2\tau\leq 2T_{1,1,j}\leq 2\tau+1/4\Delta_{1}c_{2}$.
-This establishes the first assertion.
-(ii) the case of $\mathrm{IE}=2$:
-If $C_{2}d^{-\frac{1}{2}}\leq\kappa^{2t}_{j}\leq c_{2}$ for all $t=0,1,\cdots,\tau\ (0\leq\tau\leq T_{1,j}-1)$, we have
-$\kappa^{2(\tau+1)}_{j}$ $\geq\kappa^{2\tau}_{j}+\eta^{\tau}\tilde{\eta}_{j}\kappa_{j}^{2\tau}\gamma+\eta^{\tau}\tilde{\eta}_{j}\nu^{2s}_{j}$ (B.121)
-$=\kappa^{0}_{j}+c_{1}c_{4}d^{-1}\gamma\sum_{s=0}^{\tau}\kappa_{j}^{2s}+c_{1}c_{2}d^{-1}\sum_{s=0}^{\tau}\nu^{2s}_{j}.$ (B.122)
-With high probability, $\sum_{s=0}^{\tau}\nu^{2s}_{j}$ is bounded by $C_{3}\sqrt{\tau+1}$.
-If $\tau+1\leq 4\gamma^{-2}C_{2}^{-2}C_{3}^{2}d$, by letting $c_{2}\leq\frac{1}{2}\gamma c_{1}^{-1}C_{2}^{2}C_{3}^{-2}$, we have $c_{1}c_{2}d^{-1}C_{3}\sqrt{\tau+1}\leq C_{2}d^{-\frac{1}{2}}$.
-If $\tau+1\geq 4\gamma^{-2}C_{2}^{-2}C_{3}^{2}d$, we have $C_{3}\sqrt{\tau+1}\leq\frac{1}{2}\gamma(\tau+1)C_{2}d^{-\frac{1}{2}}$.
-Thus, in either case,
-(B.122)$\geq\kappa_{j}^{0}/2+\frac{1}{2}c_{1}c_{4}d^{-1}\tilde{\varepsilon}\gamma\sum_{s=0}^{\tau}\kappa_{j}^{2s}.$ (B.123)
-This verifies that $\kappa^{2t}_{j}\geq C_{1}d^{-\frac{1}{2}}$ holds for $t=\tau+1$. By induction,
-$\kappa^{2t}$ $\geq\kappa_{j}^{0}/2+\frac{1}{2}c_{1}c_{2}d^{-1}\gamma\sum_{s=0}^{\tau}\kappa_{j}^{2s}$ (B.124)
-holds for $t$ until $\kappa^{2t}$ gets larger than $c_{2}$.
-By letting $T_{1,1}\geq\log_{(1+\frac{1}{2}c_{1}c_{2}d^{-1})}\frac{c_{2}}{C_{1}d^{-\frac{1}{2}}}$, we have $\kappa^{2t}\geq\frac{1}{2}c_{2}$ for some $t\leq T_{1,1}$, with high probability.
-Similarly to the case of (i), we can verify that $\frac{1}{4}c_{2}\leq\kappa^{2t}\leq c_{2}$ for $\tilde{\Theta}(d)$ steps.
-Therefore, we obtain the second assertion.
-∎
-B.5 From Weak Recovery to Strong Recovery
-In the previous subsection, we proved that after $t=2T_{1,1}=\tilde{\Theta}(d)$ steps, with probability $\tilde{\Omega}(1)$ over the randomness of initialization, $T_{1,1,j}$, and $\tilde{\eta}_{j}$, neurons achieve small alignment with the target direction $\frac{1}{4}c_{2}\leq\kappa^{2T_{1,1}}_{j}\leq c_{2}$.
-This subsection discusses how to convert the weak recovery into the strong recovery.
-We focus on the neurons that satisfy $\alpha_{j}\beta_{j}\geq 0$ for all $j$ as specified in Assumption 3.
-For each neuron $j$, we let $c_{a}\eta^{t}=\eta=c_{1}d^{-1}$ if $t$ is even and $\eta^{t}=0$ if $t$ is odd, for $t=2T_{1,1},\cdots,2(T_{1,1}+T_{1,2}-1)$.
-The momentum is defined as $\xi_{j}^{t}=1-\tilde{\eta}$, where $\tilde{\eta}=c_{2}\varepsilon$.
-In the following, we show that the strength of the signal is greater than some constant $\varepsilon$.
-Thus, for second-layer initialization $c_{a}\lesssim\varepsilon$, the effect of the interaction term $\nabla_{\boldsymbol{w}}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}^{2}$ can be ignored, and we drop the subscript that distinguishes $N$ neurons.
-Lemma 13.
-Consider the neuron that satisfies $\frac{1}{4}c_{2}\leq\kappa^{2T_{1,1}}_{j}\leq c_{2}$. We have
-$\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(T_{1,1}+T_{1,2})}\geq 1-\varepsilon,$ (B.125)
-with high probability, where $T_{1,2}=\Theta{O}_{d}(d\varepsilon^{-2})$.
-Proof.  The expected gradient (of the correlation term) can be computed as
-$\mathbb{E}\big{[}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}$ $=\mathbb{E}\bigg{[}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}\sum_{j=p}^{q}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}\sum_{j=0}^{C_{q}}\beta_{j}\mathsf{He}_{j}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\bigg{)}\bigg{]}$ (B.126)
-$=\sum_{j=p}^{q}\big{[}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j}\boldsymbol{w}^{2t}\big{]}.$ (B.127)
-Applying $P_{\boldsymbol{w}^{2t}}$, we have
-$\mathbb{E}\big{[}P_{\boldsymbol{w}^{2t}}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}=(\boldsymbol{\theta}-({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{\theta})\boldsymbol{w}^{2t})\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}.$ (B.128)
-The update of the alignment is
-$\kappa^{2(t+1)}\geq\kappa^{2t}+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2},$ (B.129)
-where
-$\boldsymbol{g}=P_{\boldsymbol{w}^{2t}}yc_{a}\sigma^{\prime}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\boldsymbol{x}.$ (B.130)
+$\boldsymbol{g}=P_{\boldsymbol{w}^{2t}}y\sigma^{\prime}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\boldsymbol{x}.$ (B.109)
 From (B.106), the expectation of (B.107) is bounded by
-$\mathbb{E}[\kappa^{2(t+1)}]$ $\geq\kappa^{2t}+\eta\tilde{\eta}(1-(\kappa^{2t})^{2})\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}-C_{3}\eta^{2}\tilde{\eta}^{2}d(\kappa^{2t}+\eta\tilde{\eta}).$ (B.131)
-$\geq\kappa^{2t}+\eta\tilde{\eta}\tilde{\varepsilon}p!\alpha_{p}\beta_{p}(\kappa^{2t})^{p-1}-C_{3}\eta^{2}\tilde{\eta}^{2}d(\kappa^{2t}+\eta\tilde{\eta}).$ (B.132)
-By letting $\eta\leq c_{1}d^{-1}$ and $\tilde{\eta}\leq c_{2}\varepsilon$, we have
-$\mathbb{E}[\kappa^{2(t+1)}]\geq\kappa^{2t}+\frac{1}{2}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(\kappa^{2t})^{p-1}.$ (B.133)
-Because the noise $\nu^{2t}:=\eta^{-1}\tilde{\eta}^{-1}(\kappa^{2(t+1)}-\mathbb{E}[\kappa^{2(t+1)}])$ satisfies $\mathbb{P}[|\nu^{2t}|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$,
-$\kappa^{2(T{1,1}+t)}\geq\kappa^{2T_{1,1}}+\frac{1}{2}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}\sum_{s=T_{1,1}}^{T_{1,1}+t-1}(\kappa^{2s})^{p-1}+\eta\tilde{\eta}\sum_{s=T_{1,1}}^{T_{1,1}+t-1}\nu^{2s},$ (B.134)
+$\mathbb{E}[\kappa^{2(t+1)}]$ $\geq\kappa^{2t}+\eta(1-(\kappa^{2t})^{2})\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}-\eta^{2}C_{4}d(\kappa^{2t}+\eta)$ (B.110)
+$\geq\kappa^{2t}+\eta(1-(\kappa^{2t})^{2})\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\kappa^{2t})^{p-1}-\eta^{2}C_{4}d(\kappa^{2t}+\eta).$ (B.111)
+By letting $\eta\leq c_{1}^{p-1}\varepsilon d^{-1}$, when $\kappa^{2t}\leq 1-\varepsilon$, we have
+$\mathbb{E}[\kappa^{2(t+1)}]\geq\kappa^{2t}+\frac{1}{2}\eta\varepsilon\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\kappa^{2t})^{p-1}\geq\kappa^{2t}+\eta\varepsilon c_{1}^{p}.$ (B.112)
+It is easy to see that the noise $\nu^{2t}$ has sub-Weibull tail and we obtain that
+$\kappa^{2(t+1)}\geq\kappa^{2t}+\frac{1}{2}\eta\varepsilon\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\kappa^{2t})^{p-1}+\eta\nu^{2t}\geq\kappa^{2t}+\eta\varepsilon c_{1}^{p}+\eta\nu^{2t}.$ (B.113)
+Suppose that $2c_{2}\leq\kappa^{2(T{1,1}+\tau)}\leq 1-\varepsilon$ for all $t=0,1,\dots,\tau-1$. By taking the summation of (B.113), we have
+$\kappa^{2(T{1,1}+\tau)}\geq\kappa^{2T{1,1}}+\eta\varepsilon tc_{1}^{p}+\eta\sum_{s=T{1,1}}^{T{1,1}+\tau-1}\nu^{2t}\geq 2c_{2}+\eta\varepsilon\tau c_{1}^{p}-C_{4}\eta\sqrt{\tau},$ (B.114)
 with high probability.
-The third term is bounded by $\eta\tilde{\eta}C_{3}t\leq\frac{1}{8}c_{2}$ when $t\leq\frac{1}{8}\eta^{-1}\tilde{\eta}^{-1}C_{3}^{-1}c_{2}$ and by $\eta\tilde{\eta}C_{3}\sqrt{t}\leq 4\eta^{\frac{3}{2}}\tilde{\eta}^{\frac{3}{2}}c_{2}^{-\frac{1}{2}}C_{3}^{\frac{3}{2}}t\leq\frac{1}{2}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(c_{2}/8)^{p-1}$ when $t\leq\frac{1}{8}\eta^{-1}\tilde{\eta}^{-1}C_{3}^{-1}c_{2}$ and $\varepsilon=\tilde{O}(d^{-1})$.
-Therefore, if $\kappa^{2s}\geq\frac{1}{8}c_{2}$ holds for all $s=T_{1,1},\cdots,T_{1,1}+t-1$, we have
-$\kappa^{2(T_{1,1}+t)}\geq\frac{1}{8}c_{2}+\frac{1}{4}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(c_{2}/8)^{p-1}t,$ (B.135)
-with high probability.
-Thus, by induction, $\kappa^{2(T{1,1}+t)}\geq\frac{1}{8}c_{2}$ holds and (B.135) holds for all $t$, until we get $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$.
-Because of (B.135), we have $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ until $t\leq T_{1,2}$, where $T_{1,2}\geq\big{(}\frac{1}{4}\eta\tilde{\eta}\varepsilon p!\alpha_{p}\beta_{p}(c_{2}/8)^{p-1}\big{)}^{-1}=\tilde{O}\big{(}(\eta\tilde{\eta}\varepsilon)^{-1}\big{)}=\tilde{O}(d\varepsilon^{-2})$.
-After we achieve the strong recovery $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ for some $t$, $\kappa^{2(T{1,1}+s)}$ may get smaller than $1-\varepsilon$.
-However, by letting $s^{\prime}$ be the first such step, because at each step the alignment only moves at most $\tilde{O}(\eta\tilde{\eta})=\tilde{O}(d^{-1}\varepsilon)$, $s^{\prime}$ should still satisfies $\kappa^{2(T{1,1}+s^{\prime})}\geq 1-2\varepsilon\geq c_{2}$.
-Thus, (B.135) holds again until $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$.
-Therefore, we can guarantee $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ after $t\geq T_{1,2}$, with high probability.
-∎
+The third term is bounded by $C_{4}\eta\sqrt{\tau}\leq c_{2}$ when $\tau\leq c_{2}^{2}C_{4}^{-2}\eta^{-2}=c_{2}^{2}C_{4}^{-2}\bar{c}_{\eta}^{-2}\varepsilon^{-2}d^{2}$ and by $\frac{1}{2}\eta\varepsilon\tau c_{1}^{p}$ when $\tau\geq 4\varepsilon^{-2}c_{1}^{-2p}C_{4}^{2}$.
+Because $c_{2}^{2}C_{4}^{-2}\bar{c}_{\eta}^{-2}\varepsilon^{-2}d^{2}\geq 4\varepsilon^{-2}c_{1}^{-2p}C_{4}^{2}$, we can bound (B.114) by
+$\kappa^{2(T{1,1}+\tau)}\geq c_{2}+\frac{1}{2}\eta\varepsilon\tau c_{1}^{p},$ (B.115)
+which verifies $2c_{2}\leq\kappa^{2(T{1,1}+\tau)}$.
+Therefore, by induction, until $\kappa^{2t}\geq 1-\varepsilon$, we have the lower bound (B.115), whose RHS exceeds $1-\varepsilon$ when $\tau\geq 2\eta^{-1}\varepsilon^{-1}c_{1}^{-p}\leq C_{3}d\varepsilon^{-2}$.
+Thus, there exists $\tau_{*}\leq T_{1,2}=C_{3}d\varepsilon^{-2}$ such that $\kappa^{2(T{1,1}+\tau_{*})}\geq 1-\varepsilon$, with high probability.
+Now, what remains is to prove that $\kappa^{2(T{1,1}+\tau)}\geq 1-3\varepsilon$ holds for all $\tau_{*}\leq t\leq T_{1,2}=C_{3}d\varepsilon^{-2}$.
+Because $\nu^{2t}$ are mean-zero sub-Weibull random variables, we have that $|\sum_{s=\tau}^{\tau+\tau^{\prime}-1}\nu^{2s}|\leq C_{4}\sqrt{\tau^{\prime}}$ for all $0\leq\tau,\tau^{\prime}\leq T_{1,1}$ with high probability.
+Also, because $\eta^{t}\ll\varepsilon d^{-1}$, we can easily see that $|\kappa^{2(\tau+1)}-\kappa^{2\tau}|=\tilde{O}(\varepsilon d^{-1})$ for all $\tau=0,1,\cdots,T_{1,1}-1$, with high probability.
+Thus, when there exists $\tau\geq\tau_{*}$ such that $\kappa^{2(T_{1,1}+\tau-1)}\geq 1-\varepsilon$ and $\kappa^{2(T_{1,1}+\tau)}<1-\varepsilon$, we have $\kappa^{2(T_{1,1}+\tau)}\geq 1-2\varepsilon$ with high probability.
+Moreover, following the above argument, we can inductively show that
+$\kappa^{2(T{1,1}+\tau+\tau^{\prime})}$ $\geq 1-2\varepsilon+\eta\varepsilon\tau^{\prime}c_{1}^{p}-C_{4}\eta\sqrt{\tau^{\prime}}$ (B.116)
+$\geq 1-2\varepsilon+\eta\varepsilon\tau^{\prime}c_{1}^{p}-\begin{cases}\varepsilon&(\tau^{\prime}\leq\bar{c}_{\eta}^{-2}C_{4}^{-2}d^{2})\\ \eta\varepsilon\tau^{\prime}c_{1}^{p}&(\tau^{\prime}\geq\varepsilon^{-2}C_{4}^{2}c_{1}^{-2p})\end{cases}.$ (B.117)
+$\geq 1-3\varepsilon,$ (B.118)
+for $\tau^{\prime}\leq T_{1,2}$ or until $\kappa^{2(T{1,1}+\tau+\tau^{\prime})}\geq 1-\varepsilon$ holds again.
+Note that the last inequality follows from $\bar{c}_{\eta}^{-2}C_{4}^{-2}d^{2}\geq\varepsilon^{-2}C_{4}^{2}c_{1}^{-2p}$.
+By repeating this argument (if there are multiple such $\tau$), we can see that $\kappa^{2(T{1,1}+t)}\geq 1-\varepsilon$ holds for all $\tau_{*}\leq t\leq T_{1,2}=C_{3}d\varepsilon^{-2}$ with high probability.
+Adjusting hidden constants to remove a factor of $3$ from $3\varepsilon$ yields the desired result. ∎
 B.6 Second Layer Training
-This subsection proves the generalization error after training the second layer.
+From the previous analysis, we know that at least $\Omega(1)$ portion of the neurons will satisfy the weak and strong recovery conditions (Appendix B.1), at least $\tilde{\Omega}(1)$ portion of the neurons (independent from the choice of $\sigma_{j}$) satisfy initial alignment conditions (Appendix B.2), and at least $1-o(1)$ fraction of them achieves strong recovery.
+This subsection proves a generalization error bound after second-layer training.
 Let $f_{\boldsymbol{a}}(\boldsymbol{x})=f_{\boldsymbol{\Theta}}(\boldsymbol{x})$ for $\boldsymbol{\Theta}=(\hat{\boldsymbol{w}}_{j},a_{j},\hat{b}_{j})_{j=1}^{N}$ where $\boldsymbol{a}\in\mathbb{R}^{N}$ and $(\hat{\boldsymbol{w}}_{j},\hat{b}_{j})_{j=1}^{N}$ are the parameters trained in the first stage.
-Here we let $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ be the “certificate” such that $\|\boldsymbol{a}^{*}\|=\tilde{O}(N)$ that is shown to exist in Lemma 15 (here we suppress dependence on constants $p,q$).
+Let $\boldsymbol{a}^{*}\in\mathbb{R}^{N}$ be the “certificate” with $\|\boldsymbol{a}^{*}\|^{2}=\tilde{O}(N)$ that is shown to exist in Lemma 22.
+Polynomial Link Functions.
 The following lemma is a complete version of Proposition 5.
 Lemma 20.
-There exists a $4q$-th order polynomial $Q(R_{w},b,q^{\prime})$ of $R_{\boldsymbol{w}}=\max_{j}\mathopen{}\mathclose{{}\left\|{\boldsymbol{w}_{j}}}\right\|$ and $b=(b_{j})_{j=1}^{N}$ such that, if we set $\lambda=\Theta\mathopen{}\mathclose{{}\left(\sqrt{\frac{2}{T_{2}\delta_{0}}N^{2}Q(R_{w},b,q^{\prime})}}\right)$ for some $\delta_{0}>0$, the ridge estimator $\hat{\boldsymbol{a}}$ satisfies
-$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.136)
+There exists a $4q$-th order polynomial $Q(R_{\boldsymbol{w}},b,q^{\prime})$ of $R_{\boldsymbol{w}}=\max_{j}\mathopen{}\mathclose{{}\left\|{\boldsymbol{w}_{j}}}\right\|$, $b=(b_{j})_{j=1}^{N}$ such that, if we set $\lambda=\Theta\mathopen{}\mathclose{{}\left(\sqrt{\frac{2}{T_{2}\delta_{0}}N^{2}Q(R_{w},b,q^{\prime})}}\right)$ for some $\delta_{0}>0$, the ridge estimator $\hat{\boldsymbol{a}}$ satisfies
+$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.119)
 with probability $1-\delta_{0}$.
-Therefore, by taking $T_{2}=\tilde{\Theta}((N^{4}Q(R_{w},b,q)+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}])\varepsilon^{-2})$, and $N=\tilde{\Theta}(\varepsilon^{-1})$, we have
-$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon.$
+Hence taking $T_{2}=\tilde{\Theta}((N^{4}Q^{2}(R_{w},b,q^{\prime})+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}]^{2})\varepsilon^{-4})$ and $N=\tilde{\Theta}(\varepsilon^{-1})$, we have
+$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon^{2}.$
 Proof.
@@
 Part (1).
-Here, we first bound the second term $\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}$.
-Since ^$\hat{\mathcal{L}}(f_{\hat{\boldsymbol{a}}})+\lambda\|\hat{\boldsymbol{a}}\|^{2}\leq\hat{\mathcal{L}}(f_{\boldsymbol{a}^{*}})+\lambda\|\boldsymbol{a}^{*}\|^{2}$, we have that
+We first bound the term $\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}$.
+Since ^$\hat{\mathcal{L}}(f_{\hat{\boldsymbol{a}}})+\lambda\|\hat{\boldsymbol{a}}\|^{2}\leq\hat{\mathcal{L}}(f_{\boldsymbol{a}^{*}})+\lambda\|\boldsymbol{a}^{*}\|^{2}$, we have
 $\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{T_{2}})}^{2}+\lambda\|\hat{\boldsymbol{a}}\|^{2}$ (B.120)
@@
 $=$ $2\sqrt{\frac{1}{T_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]},$ (B.130)
-where $(\sigma_{i})_{i=1}^{T_{2}}$ is the i.i.d. Rademacher sequence which is independent of $(\boldsymbol{x}_{i})_{i=1}^{T_{2}}$. Hence, Markov’s inequality yields that
+where $(\sigma_{i})_{i=1}^{T_{2}}$ is the i.i.d. Rademacher sequence independent of $(\boldsymbol{x}_{i})_{i=1}^{T_{2}}$. Hence, Markov’s inequality yields
 $\mathopen{}\mathclose{{}\left|\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{T_{2}})}^{2}-\|f_{\hat{\boldsymbol{a}}}\|_{L^{2}(P_{x})}^{2}}\right|$ $=2\|\hat{\boldsymbol{a}}\|^{2}\sqrt{\frac{1}{T_{2}\delta_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]},$
-with probabilty $1-\delta_{2}$.
+with probability $1-\delta_{2}$.
 The third term in Eq. (B.125) can be evaluated as
@@
 $\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}+\mathopen{}\mathclose{{}\left(\frac{\lambda}{4}-\sqrt{\frac{2}{T_{2}\delta_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}}\right)\|\hat{\boldsymbol{a}}\|^{2}$
-$\leq\tilde{O}(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda}\mathopen{}\mathclose{{}\left(\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{1}}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{3}}+\frac{\mathbb{E}_{\boldsymbol{x}}[(f^{*})^{4}]}{\delta_{3}}}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$
+$\leq\tilde{O}(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda}\mathopen{}\mathclose{{}\left(\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{1}}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{3}}+\frac{\mathbb{E}_{\boldsymbol{x}}[(f^{*}(\boldsymbol{x}))^{4}]}{\delta_{3}}}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$
 with probability $1-\sum_{j=1}^{4}\delta_{j}$. Hence, by setting $\lambda\geq 8\sqrt{\frac{2}{T_{2}\delta_{2}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}$, we have that
 $\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$
-$\leq\tilde{O}(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda}\mathopen{}\mathclose{{}\left(\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{1}}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}{\delta_{3}}+\frac{\mathbb{E}_{\boldsymbol{x}}[(f^{*})^{4}]}{\delta_{3}}}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2}.$
+$\leq\tilde{O}(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda}\mathopen{}\mathclose{{}\left(\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{2}]}{\delta_{1}}+\frac{\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}{\delta_{3}}+\frac{\mathbb{E}_{\boldsymbol{x}}[(f^{*}(\boldsymbol{x}))^{4}]}{\delta_{3}}}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2}.$
 When the activation function $\sigma$ is a polynomial, then each $\psi_{j}(\boldsymbol{x})=\sigma(\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle+b_{j})$ is an order $q$-polynomial of a Gaussian random variable $\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle$ which has mean 0 and variance $\mathbb{E}[\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle^{2}]=\|\boldsymbol{w}_{j}\|^{2}=\tilde{O}(1)$.
@@
 By combining evaluations of (1) and (2) together, if we let $\lambda=8\sqrt{\frac{2}{T_{2}\delta_{0}}\mathbb{E}_{\boldsymbol{x}}[\|\psi(\boldsymbol{x})\|^{4}]}$ for some $\delta_{0}>0$, (by ignoring polylogarithmic factors) we obtain that
-$!\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.148)
+$\!\!\!\!\!\!\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*}(\boldsymbol{x}))^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.131)
 with probability $1-4\delta_{0}$.
-Thus, since $\|\boldsymbol{a}^{*}\|^{2}=\tilde{O}(N)$, by setting $T_{2}=\tilde{\Theta}((N^{4}Q(R_{w},b,q^{\prime})+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}])\varepsilon^{-2})$, and $N=\tilde{\Theta}(\varepsilon^{-1})$, we obtain that (B.136)$\lesssim\varepsilon$.
+Thus, since $\|\boldsymbol{a}^{*}\|^{2}=\tilde{O}(N)$, by setting $T_{2}=\tilde{\Theta}((N^{4}Q^{2}(R_{w},b,q^{\prime})+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}]^{2})\varepsilon^{-4})$, and $N=\tilde{\Theta}(\varepsilon^{-1})$, we obtain that (B.119)$\lesssim\varepsilon^{2}$.
 ∎
-Finally, we provide the approximation guarantee: If $\sigma$ is a degree-$q$ polynomial, we have the following result, which follows Lemmas 29 and 30 of [OSSW24].
+Higher Generative Exponent Functions.
+For general link functions, under Assumption 4 and the bounded fourth moment of the link function, we have the following counterpart of Lemma 20, which provides the formal statement of Proposition 10.
+Lemma 21.
+Suppose that $\mathbb{E}[\sigma_{*}(\boldsymbol{\theta}^{\top}x)^{4}]<\infty$ and Assumption 4 hold.
+Then, by setting $\lambda=\tilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\frac{N^{2}}{T_{2}\delta_{0}}}}\right)$ for some $\delta_{0}>0$, the ridge estimator $\hat{\boldsymbol{a}}$ satisfies
+$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim\varepsilon^{2}+\frac{1}{\sqrt{T_{2}\delta_{0}}}\mathopen{}\mathclose{{}\left(N^{2}C_{4}+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{1}{\sqrt{T_{2}\delta_{0}}}\|\boldsymbol{a}^{*}\|^{2},$ (B.132)
+with probability $1-\delta_{0}$. By taking $T_{2}=\tilde{\Theta}((N^{4}+N^{2})\varepsilon^{-4})$, we have
+$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon^{2}.$
+Furthermore, applying Lemma 12 and 13 yields that, when $\sigma_{*}=\sum_{j=0}^{\infty}\alpha_{j}\mathsf{He}_{j}$ satisfies $\sum_{j=0}^{\infty}j^{2}j!\alpha_{j}^{2}$ and $\mathbb{E}[\sigma_{*}(\boldsymbol{\theta}^{\top}x)^{4}]$ are bounded, with a properly designed randomized activation in Lemma 12, by taking $N=\tilde{\Theta}(\varepsilon^{-7})$ and $T_{2}=\tilde{\Theta}(\varepsilon^{-32})$, Algorithm 1 yields
+$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon^{2},$
+with probability $1-o_{d}(1)$.
+Proof.
+The proof is identical to that of Lemma 20, with the difference being that we replace the bounded moment assumptions with $\mathbb{E}[\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{4}]<\infty$ or Assumption 4.
+∎
+Approximation Guarantee.
+Note that for non-polynomial link function with generative exponent $p_{*}\geq 3$, the approximation error is already controlled in Assumption 4 based on [BBSS22, Lemma 4.4, 4.5] (using activation function with a ReLU component).
+If $\sigma_{*}$ is a degree-$q$ polynomial, we have the following approximation result using polynomial activation, which follows Lemmas 29 and 30 of [OSSW24a].
 Lemma 22.
@@
 Then, there exists $a_{1},\dots,a_{N}$ such that
-$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1}+\varepsilon)$ (B.149)
+$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1}+\varepsilon)$ (B.133)
 with high probability, where $(\boldsymbol{x},y)$ is a random sample, and we omit dependence on the degree $q$ in the big-$O$ notation. Moreover, we have $\sum_{j=1}^{N}a_{j}^{2}=\tilde{O}(N)$.
-We rely on the following result.
+Lemma 22 can be established from the following result in [OSSW24a].
 Lemma 23.
@@
 $\mathbb{E}[\bar{v}(b;h)\sigma(\delta s+b)]=h(s).$ (B.134)
-Proof.  When $g_{q}(s)=\sigma(s)$ is a degree-$q$ polynomial,
-$g_{q}(s)=\int_{b=-q}^{0}\sigma(s+b)\mathrm{d}b$ (B.151)
-is also a degree-$q$ polynomial. Let us repeatedly define
-$g_{q-i}(s):=g_{q-(i-1)}(s+1)-g_{q-(i-1)}(s)\quad(i=1,2,\cdots,q),$ (B.152)
-and let $(c_{i,j})$ be coefficients so that $(s-1)^{i}=\sum_{j=0}^{i}c_{i,j}s^{j}$ holds for all $z$. Then, by induction, $g_{i}(s)$ is a degree-$i$ polynomial. Moreover, we have
-$g_{q-i}(s)$ $=\sum_{j=0}^{i}c_{i,j}\int_{b=-q}^{0}\sigma(s+b+j)\mathrm{d}b$ (B.153)
-$=2C_{b}\mathbb{E}_{b\sim\mathrm{Unif}([-C_{b},C_{b}])}\bigg{[}\bigg{(}\sum_{j=0}^{i}c_{i,j}\mathbbm{1}[j-q\leq b\leq j]\bigg{)}\sigma(s+b)\bigg{]},$ (B.154)
-when $C_{b}\geq q$. Therefore, for any polynomial $h(s)$ with its degree at most $q$, there exists $\bar{v}(b;h)$ with $|\bar{v}(b;h)|\lesssim C_{b}$ such that for all $s$,
-$\mathbb{E}[\bar{v}(b;h)\sigma(\delta s+b)]=h(s).$ (B.155)
-Proof of Lemma 15. We now discretize Lemma 16. We focus on $N^{\prime}$ neurons that satisfy $\|\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta}\|\leq\varepsilon$ (by letting $a_{j}=0$ otherwise).
-For $A=\tilde{\Theta}(N^{\prime})=\tilde{\Theta}(N)$ (with a small hidden constant), we consider $2A$ intervals $[-C_{b},C_{b}(-1+\frac{1}{A})),[C_{b}(-1+\frac{1}{A}),C_{b}(-1+\frac{2}{A})),\cdots,[C_{b}(1-\frac{1}{A}),C_{b}]$.
-By taking the hidden constant sufficiently small, for each interval there exists at least one $b_{j}$.
-Then, for $b_{j}$ corresponding to $[C_{b}(-1+\frac{i}{A}),C_{b}(-1+\frac{i+1}{A}))$, we set $a_{j}=\frac{N}{2}\int_{C_{b}(-1+\frac{i}{A})}^{C_{b}(-1+\frac{i+1}{A}))}\bar{v}(b;h)\mathrm{d}b$.
-Here we note that $|a_{j}|=\tilde{O}(1)$ holds for all $j$.
-If each interval contains more than one $b_{j}$, we ignore all but one by letting $a_{j}=0$ except for the one.
-By doing so, because of Lipschitzness of $\sigma$, we have
-$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(s+b_{j})-h(s)}\right|=\tilde{O}(N)$ (B.156)
-for all $s=\tilde{O}(1)$. Because $|\boldsymbol{\theta}^{\top}\boldsymbol{x}^{t}|=\tilde{O}(1)$ with high probability, we have
-$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1})$ (B.157)
-with high probability. Finally, because $\|\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta}\|\leq\varepsilon$, we have
-$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})}\right|=\tilde{O}\big{(}(\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta})^{\top}\boldsymbol{x}\big{)}=\tilde{O}(\varepsilon).$ (B.158)
-Combining (B.157) and (B.158), we obtain the assertion. ∎